### v2

# Transfer learning for high-dimensional Factor-augmented sparse linear model

Abstract
In this paper, we study transfer learning for high-dimensional factor-augmented sparse linear models, motivated by applications in economics and finance where strongly correlated predictors and latent factor structures pose major challenges for reliable estimation. Our framework simultaneously mitigates the impact of high correlation and removes the additional contributions of latent factors, thereby reducing potential model misspecification in conventional linear modeling. In such settings, the target dataset is often limited, but multiple heterogeneous auxiliary sources may provide additional information. We develop transfer learning procedures that effectively leverage these auxiliary datasets to improve estimation accuracy, and establish non-asymptotic $\ell_{1}$- and $\ell_{2}$-error bounds for the proposed estimators. To prevent negative transfer, we introduce a data-driven source detection algorithm capable of identifying informative auxiliary datasets and prove its consistency. In addition, we provide a hypothesis testing framework for assessing the adequacy of the factor model, together with a procedure for constructing simultaneous confidence intervals for the regression coefficients of interest. Numerical studies demonstrate that our methods achieve substantial gains in estimation accuracy and remain robust under heterogeneity across datasets. Overall, our framework offers a theoretical foundation and a practically scalable solution for incorporating heterogeneous auxiliary information in settings with highly correlated features and latent factor structures.
Keywords: Transfer learning, high-dimensional statistics, Lasso, Factor-augmented sparse linear model, negative transfer.
1 Introduction
In many modern statistical and machine learning applications, data are often collected from multiple related sources that share similar yet non-identical underlying structures. Directly combining these heterogeneous datasets may lead to biased estimation due to distributional discrepancies, whereas analyzing the target dataset alone can suffer from a limited sample size. Transfer learning offers a principled framework to leverage auxiliary information from source datasets to improve estimation performance on the target task.
Despite transfer learning has achieved remarkable empirical success, its statistical characterization and the theoretical justifications of related methodologies remain relatively underexplored. Several fundamental challenges persist: (i) what to transfer, (ii) how to transfer, and (iii) when to transfer (pan2009survey). In practice, it is often unclear which source datasets are informative, and incorporating uninformative or mismatched sources may lead to negative transfer. A growing body of work seeks to develop statistical foundations for transfer learning. Examples include nonparametric classification (cai2021transfer), transfer reinforcement learning (chai2025deep), and high-dimensional regression models such as linear model (li2022transfer), Gaussian graphical model (li2023transfer), and generalized linear models (bastani2021predicting; tian2023transfer). Extensions have been made to smoothed quantile regression (zhang2025transfer), multinomial regression (yang2025debiased), and adaptive Huber regression models (yang2025communication). The similarity between the source and target datasets is typically quantified via the distance between their regression coefficients (li2022transfer; tian2023transfer; zhang2025transfer; yang2025debiased; yang2025communication), whereas gu2024robust investigated an angle-based similarity measure under a random-coefficients framework, where both the sample size and dimensionality grow proportionally to infinity. There are also some other related work in recent years. For example, gross2016data; ollier2017regression explored the “data shared Lasso” in the multi-task learning framework.
However, for the linear model, as pointed out by fan2020factor, when the covariates are highly correlated-particularly in financial, economic, and medical scenarios, the traditional Lasso-type estimators often fail to achieve consistent model selection. This issue was further addressed by fan2024latent, who utilized a factor model to extract the highly correlated part. Specifically, they proposed the Factor Augmented sparse linear regression model (FARM), defined as follows:
$y$ $=\boldsymbol{f}^{\top}\boldsymbol{\gamma}^{*}+\boldsymbol{u}^{\top}\boldsymbol{\beta}^{*}+\mathcal{E},$ (1.1)
$\boldsymbol{x}$ $=\boldsymbol{B}\boldsymbol{f}+\boldsymbol{u},$
where $y\in\mathbb{R}$ is a scalar response variable, $\boldsymbol{x}\in\mathbb{R}^{p}$ is the covariate, $\boldsymbol{f}\in\mathbb{R}^{K}$ is the latent factors, $\boldsymbol{B}\in\mathbb{R}^{p\times K}$ is the factor loading matrix, and $\boldsymbol{u}\in\mathbb{R}^{p}$ is a idiosyncratic component. The parameter $\boldsymbol{\beta}^{*}\in\mathbb{R}^{p}$ is the true regression parameter vector that is of interest, and $\boldsymbol{\gamma}^{*}\in\mathbb{R}^{K}$ quantifies the contribution of the latent factor $\boldsymbol{f}$. The FARM model can be viewed as a generalization of the sparse linear model, as it can be rewritten as:
$y$ $=\boldsymbol{f}^{\top}\boldsymbol{\varphi}^{*}+\boldsymbol{x}^{\top}\boldsymbol{\beta}^{*}+\mathcal{E},$
$\boldsymbol{x}$ $=\boldsymbol{B}\boldsymbol{f}+\boldsymbol{u},$
where $\boldsymbol{\varphi}^{*}=\boldsymbol{\gamma}^{*}-\boldsymbol{B}^{\top}\boldsymbol{\beta}^{*}$ measures the extra contribution of $\boldsymbol{f}$ to $y$ beyond $\boldsymbol{x}$. It enhances sparse linear regression by incorporating informative directions spanned by the common factors $\boldsymbol{f}$. From another point of view, the FARM model can be viewed as a generalization of the factor regression model (stock2002forecasting; bing2021prediction) as it utilizes additional information in $\boldsymbol{u}$. Therefore, the strong performance of FARM arises from its ability to jointly capture both the common factors $\boldsymbol{f}$ and idiosyncratic components $\boldsymbol{u}$ in explaining the response.
Although the FARM model effectively addresses the challenge of highly correlated covariates by separating common factors and idiosyncratic components, it is primarily designed for a single dataset. In practice, the target dataset may have a limited sample size, and related auxiliary datasets often contain useful information that can enhance estimation accuracy. To this end, combining the FARM framework with transfer learning is particularly appealing.
Motivated by these insights, we propose a transfer learning procedure for FARM (Trans-FARM) in this paper. The Trans-FARM allows information to be shared across multiple related datasets while accounting for the heterogeneity in both the factor structure and the idiosyncratic components. By removing the effects of strong correlations among covariates and the additional variations induced by latent factors, Trans-FARM yields more accurate estimation of regression coefficients. Compared with standard FARM, it improves estimation by borrowing similarity from informative sources; compared with transfer learning methods for linear models (li2022transfer; tian2023transfer), it provides more reliable estimation due to their misspecification of model. Furthermore, the proposed framework enables statistical testing on regression parameters, which can be applied to assess the validity of transfer learning for factor regression in lorenzi2016transfer.
Specifically, our model is formulated as follows. Given a target dataset $\{\boldsymbol{x}_{i}^{(0)},y_{i}^{(0)}\}_{i=1}^{n_{0}}$ and $K$ independent source datasets $\left\{\{\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)}\}_{i=1}^{n_{k}}\right\}_{k=1}^{K}$, the model can be written as the following matrix form:
[…]
$\boldsymbol{X}^{(k)}$ $=\boldsymbol{F}_{k}{\boldsymbol{B}_{k}}^{\top}+\boldsymbol{U}_{k},\quad k=0,\cdots,K.$
[…]
Compared with the Trans-Lasso and Trans-GLM methods in li2022transfer; tian2023transfer, our proposed Trans-FARM possesses several merits. First, the latent factors $\boldsymbol{f}^{(k)}$ are treated as unobserved confounders. By explicitly accounting for these confounding effects, our model removes the extra contribution of latent factors to the response, thereby yielding a cleaner target of estimation. Second, even when the latent factors do not provide additional explanatory power for the response, our estimator achieves the same convergence rate as the Trans-Lasso estimator in li2022transfer; tian2023transfer. Moreover, since the sub-Gaussian assumption on $\boldsymbol{x}$ fails under strong feature correlations, existing transfer-learning theories do not cover such settings. Our framework addresses this gap by allowing the design matrix to exhibit high correlation. Our main contributions are summarized as follows:
  • A Robust Transfer Learning Framework for Confounded Data. We develop a unified transfer learning framework that explicitly accounts for latent factors and highly correlated features. This approach effectively removes confounding effects and prevents model misspecification, which is crucial for applications in economics, finance, and biomedicine. The framework systematically leverages auxiliary datasets to enhance the estimation of regression coefficients, providing a solid foundation for both accurate estimation and valid inference in complex data environments.
  • Theoretical Guarantees and a Safeguard Against Negative Transfer. We establish the non-asymptotic $\ell_{1}$- and $\ell_{2}$-error bounds for our proposed Oracle-Trans-FARM estimator, demonstrating its efficiency. Theoretically, it achieves a convergence rate that matches state-of-the-art methods when factors are absent, while strictly outperforming them when latent factors are present. Furthermore, to ensure this superior performance in practice, we propose a novel source detection algorithm with proven consistency, which acts as a safeguard by identifying and utilizing only beneficial auxiliary datasets to avoid negative transfer.
  • Valid Inference for Heterogeneous Effects. Moving beyond point estimation, we construct a comprehensive inferential toolkit for the high-dimensional transfer learning setting. This includes a hypothesis testing framework for assessing factor model adequacy and a method for constructing simultaneous confidence intervals for regression coefficients. A key innovation is a studentized approach that yields coefficient-specific interval lengths, thereby accurately capturing the heterogeneous influence of different features on the response.
The rest of this paper is structured as follows. Section 2 introduces our estimation procedure for the unobserved factors and idiosyncratic components, and presents an oracle transfer learning algorithm (Algorithm 1 under the FARM framework (where we know which sources are informative). Section 3 develops the theoretical results for the estimator attained via Algorithm 1, including the $\ell_{1}$ and $\ell_{2}$- estimation error bounds, and introduces the source detection algorithm together with its detection consistency. The procedure for simultaneous inference is presented at the end of Section 3. Section 4 presents simulation results, and Section 5 presents our conclusions.
Notation: For a vector $\mathbf{v}=(v_{1},\cdots,v_{p})^{\top}\in\mathbb{R}^{p}$, denote $\|\mathbf{v}\|=(\sum_{i=1}^{p}|v_{i}|^{2})^{1/2}$, $\|\mathbf{v}\|_{1}=\sum_{i=1}^{p}|v_{i}|$, $\|\mathbf{v}\|_{\infty}=\max_{1\leq i\leq p}|v_{i}|$. For $\mathcal{S}\subseteq[p]$, let $\mathcal{S}^{c}$ be the complement of $\mathcal{S}$, i.e., $\mathcal{S}^{c}=[p]\backslash\mathcal{S}$, where we use the standard notation $[p]$ to denote $\{1,\ldots,p\}$ here. For a matrix $\mathbf{A}=(a_{ij})_{1\leq i\leq k_{1},1\leq j\leq k_{2}}\in\mathbb{R}^{k_{1}\times k_{2}}$, denote $\|\mathbf{A}\|_{\max}=\max_{1\leq i\leq k_{1},1\leq j\leq k_{2}}|a_{ij}|$, $\|\mathbf{A}\|_{\ell_{1}}=\sum_{i=1}^{k_{1}}\sum_{j=1}^{k_{2}}|a_{ij}|$, $\|\mathbf{A}\|_{\infty}=\max_{1\leq i\leq k_{1}}\sum_{j=1}^{k_{2}}|a_{ij}|$, $\|\mathbf{A}\|_{1}=\max_{1\leq j\leq k_{2}}\sum_{i=1}^{k_{1}}|a_{ij}|$, and let $\|\mathbf{A}\|_{2}$ be the spectral norm of $\mathbf{A}$. We use the standard notation $\boldsymbol{A}_{i,j}$ to represent the $(i,j)$-th element of $\boldsymbol{A}$, and use $\boldsymbol{A}_{i,-j}$ to denote the $i$-th row without the $(j,j)$-th element of $\boldsymbol{A}$ for a given matrix $\boldsymbol{A}$. We define the sub-Gaussian norm of a sub-Gaussian random variable $X$ as $\|X\|_{\psi_{2}}=\inf\{t>0:\mathbb{E}\exp(X^{2}/t^{2})\leq 2\}$, and the sub-exponential norm of a sub-exponential random variable $Y$ as $\|Y\|_{\psi_{1}}=\inf\{t>0:\mathbb{E}\exp(|Y|/t)\leq 2\}$. For any two sequences $\{a_{n}\}$ and $\{b_{n}\}$, we use $a_{n}=O(b_{n}),a_{n}=o(b_{n}),$ and $a_{n}\asymp b_{n}$ to represent that there exists $c,C>0$ such that $|a_{n}|\leq C|b_{n}|$ for all $n$, $\lim_{n\to\infty}|a_{n}|/|b_{n}|=0$, and $0<c<|a_{n}/b_{n}|<C<\infty$, respectively. $a_{n}\lesssim b_{n}$ means $a_{n}=O(b_{n})$. For a sequence of random variables $\{X_{n}\}$, we use the notation $Z_{n}=O_{P}(1)$ if $\lim_{M\to\infty}\lim_{n\to\infty}\mathbb{P}(|Z_{n}|>M)=0$, and $Z_{n}=o_{P}(1)$ if $\lim_{n\to\infty}\mathbb{P}(|Z_{n}|>\varepsilon)=0$ for any $\varepsilon>0$. For any two sequences of random variables $\{X_{n}\}$ and $\{Y_{n}\}$, we say $X_{n}=O_{P}(Y_{n})$ if $X_{n}/Y_{n}=O_{P}(1)$, and $X_{n}=o_{P}(Y_{n})$ if $X_{n}/Y_{n}=o_{P}(1)$.
2 Preliminaries and Problem Statement
2.1 Factor estimation
The FARM is a basic tool to analyze the high-dimensional data with correlated features, as it provides a unified framework that integrates dimension reduction and sparse regression by simultaneously accounting for the common factors $\boldsymbol{f}$ and idiosyncratic components $\boldsymbol{u}$ in explaining the response. In this section, we consider how to estimate the FARM.
Recall that we suppose the observations $\{(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)},\boldsymbol{f}_{i}^{(k)},\boldsymbol{u}_{i}^{(k)},\mathcal{E}_{i}^{(k)})\}_{i=1}^{n}$ are generated from model (1.2):
$y^{(k)}$ $=\boldsymbol{f}^{(k)\top}\boldsymbol{\gamma}^{(k)}+\boldsymbol{u}^{(k)\top}\boldsymbol{w}^{(k)}+\mathcal{E}^{(k)},$
$\boldsymbol{x}^{(k)}$ $=\boldsymbol{B}_{k}\boldsymbol{f}^{(k)}+\boldsymbol{u}^{(k)}.$
According to (1.2), the model can be expressed in matrix form as
[…]
$\boldsymbol{X}^{(k)}$ $=\boldsymbol{F}_{k}\boldsymbol{B}_{k}^{\top}+\boldsymbol{U}_{k},\quad k=0,\cdots,K.$
It can also be written as
[…]
$\boldsymbol{X}^{(k)}$ $=\boldsymbol{F}_{k}\boldsymbol{B}_{k}^{\top}+\boldsymbol{U}_{k},\quad k=0,\cdots,K,$
where $\boldsymbol{\varphi}^{(k)}=\boldsymbol{\gamma}^{(k)}-\boldsymbol{B}_{k}^{\top}\boldsymbol{w}^{(k)}$ quantifies the additional effect of factors $\boldsymbol{F}_{k}$ to $\boldsymbol{Y}^{(k)}$.
Since the latent factor $\boldsymbol{f}_{i}^{(k)}$, the idiosyncratic component $\boldsymbol{u}_{i}^{(k)}$, and the loading matrix $\boldsymbol{B}_{k}$ are all needed to be estimated, we first provide their estimators. Note that $\boldsymbol{f}_{i}$ is not identifiable because $\boldsymbol{x}_{i}^{(k)}=\boldsymbol{B}_{k}\boldsymbol{f}_{i}^{(k)}+\boldsymbol{u}_{i}^{(k)}=\boldsymbol{B}_{k}\boldsymbol{T}(\boldsymbol{T}^{-1}\boldsymbol{f}_{i}^{(k)})+\boldsymbol{u}_{i}^{(k)}$ for any non-singular matrix $\boldsymbol{T}\in\mathbb{R}^{r_{k}\times r_{k}}$. To ensure identifiability, we impose the following normalization:
$\mbox{cov}(\boldsymbol{f}^{(k)})=\boldsymbol{I}_{r_{k}}\text{ and }\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k}\text{ is diagonal.}$
We also assume that $\mathcal{E}^{(k)}$, $\boldsymbol{f}^{(k)}$ and $\boldsymbol{u}^{(k)}$ are independent with each other, and $\mathbb{E}[\mathcal{E}^{(k)}]=0$.
Then the estimates simply follows by solving the constrained least squares problem for $k=0,\ldots,K$:
$(\hat{\boldsymbol{F}}_{k},\hat{\boldsymbol{B}}_{k})=\mathop{\arg\min}_{\boldsymbol{F}\in\mathbb{R}^{n\times r_{k}},\boldsymbol{B}\in\mathbb{R}^{p\times r_{k}}}\|\boldsymbol{X}^{(k)}-\boldsymbol{F}\boldsymbol{B}^{\top}\|_{\mathbb{F}}^{2}$
$\text{subject to }\frac{1}{n}\boldsymbol{F}^{\top}\boldsymbol{F}=\boldsymbol{I}_{r_{k}}\text{, and }\boldsymbol{B}^{\top}\boldsymbol{B}\text{ is diagonal.}$
Simple calculation yields that $\hat{\boldsymbol{F}}_{k}/\sqrt{n_{k}}$ is the eigenvectors corresponding to the top $r_{k}$ eigenvectors of $\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top}$, $\hat{\boldsymbol{B}}_{k}=(\hat{\boldsymbol{F}}_{k}^{\top}\hat{\boldsymbol{F}}_{k})^{-1}\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{X}^{(k)}=\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{X}^{(k)}/n_{k}$, and the estimator for the remainder term is $\hat{\boldsymbol{U}}_{k}=\boldsymbol{X}^{(k)}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{B}}_{k}^{\top}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}/n_{k})\boldsymbol{X}^{(k)}$. Then we have $\boldsymbol{X}^{(k)}=\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{B}}_{k}^{\top}+\hat{\boldsymbol{U}}_{k}$, and $\hat{\boldsymbol{F}}_{k}$, $\hat{\boldsymbol{B}}_{k}$, and $\hat{\boldsymbol{U}}_{k}$ are estimations for $\boldsymbol{F}_{k}$, $\boldsymbol{B}_{k}$, and $\boldsymbol{U}_{k}$ in model (2.1).
Remark 2.1.
If the intercept is considered, i.e., the first column of $\boldsymbol{X}^{(k)}$ is all $1$, let $\boldsymbol{X}_{-1}^{(k)}=(\boldsymbol{x}_{1,-1}^{(k)},\cdots,\boldsymbol{x}_{n,-1}^{(k)})^{\top}\in\mathbb{R}^{n_{k}\times(p-1)}$ denote the submatrix of $\boldsymbol{X}^{(k)}$ comprising its second through $p$th columns. By removing the intercept term, similar calculation yields that $\hat{\boldsymbol{F}}_{k}/\sqrt{n_{k}}$ is the eigenvectors corresponding to the top $r_{k}$ eigenvectors of $\boldsymbol{X}_{-1}^{(k)}{\boldsymbol{X}_{-1}^{(k)}}^{\top}$, $\hat{\boldsymbol{B}}_{k,-1}=\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{X}_{-1}^{(k)}/n_{k}$, and $\hat{\boldsymbol{U}}_{k,-1}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}/n_{k})\boldsymbol{X}_{-1}^{(k)}$. Then we can write $\hat{\boldsymbol{B}}_{k}=(\boldsymbol{0}_{r_{k}},\hat{\boldsymbol{B}}_{k,-1}^{\top})^{\top}\in\mathbb{R}^{p\times r_{k}}$, $\hat{\boldsymbol{U}}_{k}=(\boldsymbol{1}_{n_{k}},\hat{\boldsymbol{U}}_{k,-1})\in\mathbb{R}^{n_{k}\times p}$ such that $\boldsymbol{X}^{(k)}=\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{B}}_{k}^{\top}+\hat{\boldsymbol{U}}_{k}$, and $\hat{\boldsymbol{F}}_{k}$, $\hat{\boldsymbol{B}}_{k}$, and $\hat{\boldsymbol{U}}_{k}$ are estimations for $\boldsymbol{F}_{k}$, $\boldsymbol{B}_{k}$, and $\boldsymbol{U}_{k}$, respectively.
To attain the consistency of these estimations, we first introduce some regularity assumptions.
Assumption 1 (Assumptions on $\boldsymbol{f}^{(k)}$, $\boldsymbol{u}^{(k)}$, and $\boldsymbol{B}_{k}$ (for the $k$th dataset)).
  1. (a)
    Assume that $\boldsymbol{\nu}^{(k)}=(\boldsymbol{u}^{(k)\top},\boldsymbol{f}^{(k)\top})^{\top}$ is sub-Gaussian with $\|\boldsymbol{\nu}^{(k)}\|_{\psi_{2}}\leq v_{0}$.
  2. (b)
    There exists a constant $C>1$ such that $p/C\leq\lambda_{\min}(\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k})\leq\lambda_{\max}(\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k})\leq pC$. Moreover, we assume that $n_{k}=O(p)$.
  3. (c)
    There exists a constant $\Upsilon>0$ such that $\|\boldsymbol{B}_{k}\|_{\max}\leq\Upsilon$ and $\mathbb{E}|\boldsymbol{u}^{(k)\top}\boldsymbol{u}^{(k)}-\mbox{tr}(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)})|^{4}\leq\Upsilon p^{2}$.
  4. (d)
    There exists a positive constant $\iota<1$ such that $\iota\leq\lambda_{\min}(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)})$, $\|\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\|_{1}\leq 1/\iota$ and $\min_{1\leq i,\ell\leq p}\mbox{Var}(u_{i}^{(k)}u_{\ell}^{(k)})\geq\iota$.
Remark 2.2.
Assumption 1 is a standard assumption in the high-dimensional factor model, see bai2003inferential; fan2013large; fan2024latent.
The consistency of factor estimations is summarized in the following lemma, which follows from Lemmas D.2 and D.3 in wang2017asymptotics.
Lemma 2.1.
Define $\boldsymbol{H}_{k}=n_{k}^{-1}\boldsymbol{V}_{k}^{-1}\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{F}_{k}\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k}$, where $\boldsymbol{V}_{k}\in\mathbb{R}^{r_{k}\times r_{k}}$ is a diagonal matrix with its diagonal elements being the first $r_{k}$ largest eigenvalues of $n_{k}^{-1}\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)\top}}$. If Assumption 1 holds for any $k\in\{0,\ldots,K\}$, and $\log n_{k}=o(p)$, for $k\in\{0,\ldots,K\}$, we then have for the $k$th dataset,
  • $\|\hat{\boldsymbol{F}}_{k}-\boldsymbol{F}_{k}\boldsymbol{H}_{k}^{\top}\|_{\mathbb{F}}^{2}=O_{P}(n_{k}/p+1/n_{k})$, $\|\hat{\boldsymbol{F}}_{k}-\boldsymbol{F}_{k}\boldsymbol{H}_{k}^{\top}\|_{\max}^{2}=O_{P}((1/p+1/n_{k}^{2})\log^{2}n_{k})$.
  • For any $\mathcal{J}\subset\{1,\cdots,p\}$, $\max_{j\in\mathcal{J}}\sum_{i=1}^{n_{k}}|\hat{u}_{ij}^{(k)}-u_{ij}^{(k)}|^{2}=O_{P}(\log|\mathcal{J}|+n_{k}/p).$
  • $\|\boldsymbol{H}_{k}^{\top}\boldsymbol{H}_{k}-\boldsymbol{I}_{r_{k}}\|=O_{P}(1/n_{k}+1/p)$.
  • $\max_{j\in[p]}\|\hat{\boldsymbol{b}}_{j}^{(k)}-\boldsymbol{H}_{k}\boldsymbol{b}_{j}^{(k)}\|_{2}^{2}=O_{P}(\log p/n_{k})$, $\|\hat{\boldsymbol{B}}_{k}-\boldsymbol{B}_{k}\boldsymbol{H}_{k}^{-1}\|_{\max}^{2}=O_{P}(\log p/n_{k})$.
Remark 2.3.
The true number of latent factors $r_{k}$ in each dataset is unknown in practice, and determining $r_{k}$ in a data-driven manner is a crucial task. A rich literature has proposed various approaches for estimating $r_{k}$ (bai2002determining; lam2012factor; ahn2013eigenvalue; fan2022estimating). Our theoretical results remain valid as long as $r_{k}$ is replaced by any consistent estimator $\hat{r}_{k}$, i.e., we only require
$\mathbb{P}(\hat{r}_{k}=r_{k})\to 1,\quad\text{as }\quad n_{k}\to\infty.$
Therefore, we assume throughout this paper that the number of factors $r_{k}$ is known. For empirical implementation, we adopt the eigenvalue ratio method (lam2012factor; ahn2013eigenvalue) to determine $r_{k}$. Specifically, let $\lambda_{i}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)\top}})$ denote the eigenvalues of the Gram matrix $\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)\top}}$. Then the estimator of the number of factors is given by
$\hat{r}_{k}=\mathop{\arg\max}_{i\leq\mathcal{K}}\frac{\lambda_{i}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top})}{\lambda_{i+1}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top})},$
where $1\leq\mathcal{K}\leq n_{k}$ is a prescribed upper bound for $r_{k}$.
2.2 Factor Augmented Transfer Learning
Firstly, we consider the way to attain an estimator for each dataset. It is commonly presumed that only a small subset of predictors exerts influence on the response variable under the high-dimensional setting. This corresponds to assuming that the true coefficient vector $\boldsymbol{w}^{(k)}$ is sparse. Accordingly, the regularized estimators of the unknown parameter vectors $\boldsymbol{w}^{(k)}$ and $\boldsymbol{\gamma}^{(k)}$ for $k=0,\cdots,K$ in our factor-augmented linear model are
$(\hat{\boldsymbol{w}}^{(k)},\hat{\boldsymbol{\gamma}}^{(k)})=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p},\boldsymbol{\gamma}\in\mathbb{R}^{r_{k}}}\left\{\frac{1}{2n_{k}}\|\boldsymbol{Y}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}-\hat{\boldsymbol{F}}_{k}\boldsymbol{\gamma}\|_{2}^{2}+\lambda\|\boldsymbol{w}\|_{1}\right\},$ (2.3)
where $\lambda>0$ is a tuning parameter. Let $\hat{\boldsymbol{P}}_{k}=n_{k}^{-1}\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}$ be the projection matrix. Denote $\tilde{\boldsymbol{Y}}^{(k)}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{P}}_{k})\boldsymbol{Y}^{(k)}$. Recall that $\hat{\boldsymbol{U}}_{k}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}/n_{k})\boldsymbol{X}^{(k)}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{P}}_{k})\boldsymbol{X}^{(k)}$, which implies that $\hat{\boldsymbol{F}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}=\boldsymbol{0}_{r_{k}\times p}$ and (2.3) can be equivalently written as
$\hat{\boldsymbol{w}}^{(k)}$ $=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\left\{\frac{1}{2n_{k}}\|\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}\|_{2}^{2}+\lambda\|\boldsymbol{w}\|_{1}\right\}$ (2.4)
$=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\left\{\frac{1}{2n_{k}}\|(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{P}}_{k})(\boldsymbol{Y}^{(k)}-\boldsymbol{X}^{(k)}\boldsymbol{w})\|_{2}^{2}+\lambda\|\boldsymbol{w}\|_{1}\right\}$
$\hat{\boldsymbol{\gamma}}^{(k)}$ $=(\hat{\boldsymbol{F}}_{k}^{\top}\hat{\boldsymbol{F}}_{k})^{-1}\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{Y}^{(k)}=\frac{1}{n_{k}}\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{Y}^{(k)}$
Inspired by the insights of transfer learning, we denote $\boldsymbol{\beta}=\boldsymbol{w}^{(0)}$ as the true parameter of the target data and let $\boldsymbol{\delta}^{(k)}=\boldsymbol{\beta}-\boldsymbol{w}^{(k)}$ as the $k$th constract vector. We use the $\ell_{1}$-distance $\|\boldsymbol{\delta}^{(k)}\|_{1}$ to measure the similarity of the target dataset and the $k$th source dataset. We begin with considering a general transfer learning algorithm on FARM when the transferable set $\mathcal{A}_{\eta}$ is known, where $\mathcal{A}_{\eta}=\{1\leq k\leq K:\|\boldsymbol{\delta}^{(k)}\|_{1}\leq\eta\}$ for some $\eta$. According to (2.4), we are motivated to attain the initial estimation of $\boldsymbol{\beta}$ through the following regularized loss function, denoted by $\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}$:
$\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\left\{\frac{1}{2(n_{\mathcal{A}_{\eta}}+n_{0})}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}\|_{2}^{2}+\lambda_{\boldsymbol{w}}\|\boldsymbol{w}\|_{1}\right\},$ (2.5)
where $n_{\mathcal{A}_{\eta}}=\sum_{k\in\mathcal{A}_{\eta}}n_{k}$, and $\lambda_{\boldsymbol{w}}$ is a tuning parameter. The equation above can be understood as finding a solution that converges to its population version $\boldsymbol{w}^{\mathcal{A}_{\eta}}$ under certain regularization conditions, where $\boldsymbol{w}^{\mathcal{A}_{\eta}}$ satisfies
$\boldsymbol{w}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\mathbb{E}\left[\frac{1}{2(n_{\mathcal{A}_{\eta}}+n_{0})}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\boldsymbol{Y}^{(k)}-\boldsymbol{U}_{k}\boldsymbol{w}-\boldsymbol{F}_{k}\boldsymbol{\gamma}^{(k)}\|_{2}^{2}\right].$ (2.6)
[…]
$\mathbb{E}\left[\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})/n_{k}\right]=\mathbb{E}\left[\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{u}^{(k)}\boldsymbol{u}^{(k)\top}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})\right]=\boldsymbol{0}_{p},$
where $\alpha_{k}=n_{k}/(n_{\mathcal{A}_{\eta}}+n_{0})$. Hence, $\boldsymbol{w}^{\mathcal{A}_{\eta}}$ has a explicit form: $\boldsymbol{w}^{\mathcal{A}_{\eta}}=(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}})^{-1}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\boldsymbol{w}^{(k)}$, where $\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}}=\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}$, and $\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}=\mathbb{E}[\boldsymbol{u}^{(k)}\boldsymbol{u}^{(k)\top}]$.
Inspired by bastani2021predicting; tian2023transfer; li2022transfer; zhang2025transfer, define $\boldsymbol{\delta}^{\mathcal{A}_{\eta}}=\boldsymbol{\beta}$, we attain its estimator $\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}$ via solving:
$\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{\delta}\in\mathbb{R}^{p}}\left\{\frac{1}{2n_{0}}\|\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}(\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\boldsymbol{\delta})\|_{2}^{2}+\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\delta}\|_{1}\right\},$ (2.7)
where $\lambda_{\boldsymbol{\delta}}$ is a tuning parameter. We regard $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}$ as the estimator of $\boldsymbol{\beta}$. The algorithm is summarized in Algorithm 1.
Input: Target data $(\boldsymbol{X}^{(0)},\boldsymbol{Y}^{(0)})$, source data $\{(\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)})\}_{k=1}^{K}$, penalty parameters $\lambda_{\boldsymbol{w}}$ and $\lambda_{\boldsymbol{\delta}}$, transferring set $\mathcal{A}_{\eta}$, factor bound $\mathcal{K}$.
Output: Estimated coefficient vector $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}$.
1 Preprocessing step: Estimate $\hat{r}_{k}=\arg\max_{i\leq\mathcal{K}}\dfrac{\lambda_{i}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top})}{\lambda_{i+1}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top})}$. Let $\hat{\boldsymbol{F}}_{k}/\sqrt{n_{k}}$ be the eigenvectors corresponding to the top $\hat{r}_{k}$ eigenvalues of $\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top}$. Compute $\hat{\boldsymbol{B}}_{k}=\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{X}^{(k)}/n_{k}$ and $\hat{\boldsymbol{U}}_{k}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}/n_{k})\boldsymbol{X}^{(k)}$.
2 Transferring step: Compute (2.5), i.e.,
$\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\left\{\frac{1}{2(n_{\mathcal{A}_{\eta}}+n_{0})}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}\|_{2}^{2}+\lambda_{\boldsymbol{w}}\|\boldsymbol{w}\|_{1}\right\}$
3 Debiasing step: Compute (2.7), i.e.,
$\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{\delta}\in\mathbb{R}^{p}}\left\{\frac{1}{2n_{0}}\|\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}(\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\boldsymbol{\delta})\|_{2}^{2}+\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\delta}\|_{1}\right\}$
4 Let $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}$.
Algorithm 1 Oracle-Trans-FARM Algorithm
3 Statistical analysis
3.1 Estimation error bounds for Oracle-Trans-FARM algorithm
Several technical assumptions are required to establish the theoretical results.
Assumption 2 (Assumption on $\mathcal{E}^{(k)}$).
There exists a positive constant $c<\infty$ such that $\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\mathcal{E}^{(k)}\|_{\psi_{2}}\leq c$.
Assumption 3.
There exists a positive constant $C_{1}<\infty$ such that $\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}})^{-1}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\|_{1}\leq C_{1}$, where $\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}}=\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}$, and $\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}=\mathbb{E}[\boldsymbol{u}^{(k)}\boldsymbol{u}^{(k)\top}]$.
Assumption 2 is standard and mild in high-dimensional regression models (fan2024latent; zhang2017simultaneous; tian2023transfer). Assumption 3 is also commonly adopted in transfer learning settings (tian2023transfer; zhang2025transfer). In particular, Assumption 3 is reasonable as it does not require exact sparsity of $(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}})^{-1}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}$ and is satisfied, for example, by Toeplitz-type covariance structures. This assumption can be further interpreted as a constraint on the heterogeneity between the target idiosyncratic component and the source idiosyncratic component; see Condition 4 in li2023transfer for a detailed discussion. However, since the predicators are not sub-Gaussian if factors exist, our assumptions can be regarded as weaker conditions. Moreover, this assumption can be further simplified as a restriction on heterogeneity between target predictors and source predictors. More discussions can be found in Condition 4 of li2023transfer. This assumption can also be relaxed at the cost of slightly weaker estimation bounds, following arguments similar to those in Theorem 3 of tian2023transfer.
Now, we are ready to present our main results for the estimation error bounds of the Oracle Trans-FARM algorithm proposed in Algorithm 1. Given the transferable set $\mathcal{A}_{\eta}\subset\{1,\ldots,K\}$, we consider the parameter space
[…]
Theorem 3.1.
Suppose that Assumptions 1-2 hold for $k\in\{0\}\cup\mathcal{A}_{\eta}$ and Assumption 3 holds. Assume that $\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}n_{k}=O(p)$, $\eta=O(1)$, the following (3.1) holds for a constant $c$ small enough:
$\max\left\{\sqrt{\frac{s\log p}{n_{0}}},\frac{s(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}},\frac{s}{p}\right\}<c,$ (3.1)
and
$\max\left\{\frac{(|\mathcal{A}_{\eta}|+1)(1\vee\eta\log^{\frac{1}{2}}p\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2}\log^{-\frac{1}{2}}p)}{\sqrt{n_{\mathcal{A}_{\eta}}+n_{0}}},\frac{\mathcal{V}_{n_{0},p}\|\varphi^{(0)}\|_{2}}{\sqrt{n_{0}\log p}}\right\}=O(1),$ (3.2)
where
$\mathcal{V}_{n_{k},p}=\frac{n_{k}}{p}+\sqrt{\frac{\log p}{n_{k}}}+\sqrt{\frac{n_{k}\log p}{p}}.$
[…]
$\|\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}\|_{2}$ $\lesssim\sqrt{\frac{\log p}{n_{0}}}\eta+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}$
$\|\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}\|_{1}$ $\lesssim s\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\eta+\sqrt{s\eta}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}$
Remark 3.1.
We provide some illustrations of the conditions in Theorem 3.1. The dimension-type condition $n_{k}=O(p)$ is widely applied in factor analysis to ensure consistency. Conditions (3.1)-(3.2) are also mild. Consider a regime that $\{n_{k}\}_{k\in\{0\}\cup\mathcal{A}_{\eta}}$ are the same order, the additional contribution of factors is bounded by $\|\varphi^{(k)}\|_{2}=O_{P}(\sqrt{\log p})$, and $\eta=O(1)$, (3.2) and the second term in (3.1) changes into $\sqrt{(|\mathcal{A}_{\eta}|+1)\log p/n_{0}}=O(1)$ and $s\log p/n_{0}<c$ for some small constant $c$.
From Theorem 3.1, the $\ell_{1}$- and $\ell_{2}$-error bounds of our estimator improve upon those in fan2024latent provided that $\eta\ll s\sqrt{\log p/n_{0}}$ and $n_{\mathcal{A}_{\eta}}+n_{0}\gg n_{0}$. This demonstrates the advantage gained through transfer learning.
3.2 Consistency of Source detection
In this subsection, we consider how to detect the transferable set $\mathcal{A}_{\eta}$ through a data-driven manner, which is not known in practice. Since the performance of transfer learning deeply depends on the transferable set, the negative transfer (pan2009survey; ge2014handling) would happen if irrelevant sources are included in the transferable set. Therefore, a reliable procedure for identifying which source datasets can be effectively leveraged is crucial.
Motivated by tian2023transfer; yang2025communication; yang2025debiased, we develop a source detection procedure as follows. We begin by randomly partitioning the target dataset into three folds, denoted by $\{(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]})\}_{r=1}^{3}$. Note that we choose three folds and suppose $n_{0}$ is divisible by $3$ only for convenience; alternative numbers of folds are also examined in our simulation studies. Next, for each source dataset, we apply the transferring step using every two folds of target data and evaluate the resulting model on the remaining fold. We compute the corresponding value of the given loss function on the remaining target fold and average across the three splits to obtain the cross-validated loss $\hat{L}_{0}^{(k)}$ for each source $k$. As a benchmark, we also fit the Lasso estimator using every pair of target folds and evaluate its performance on the remaining fold, yielding the average cross-validated loss $\hat{L}_{0}^{(0)}$, which serves as the loss of the target-only model. Finally, we compare the difference $\hat{L}_{0}^{(k)}-\hat{L}_{0}^{(0)}$ with a pre-specified threshold. We then include source $k$ in $\widehat{\mathcal{A}}$ if $\hat{L}_{0}^{(k)}-\hat{L}_{0}^{(0)}$ is below this threshold. Write the $r$-th fold of target data is $(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]})$ with the data included in it denoting by $\mathcal{R}^{[r]}\subset[n_{0}]$. Let $\boldsymbol{e}^{(0)[r]}$ be the $n_{0}$-dimensional vector with $\boldsymbol{e}_{\mathcal{R}^{[r]}}^{(0)[r]}=1$ and $\boldsymbol{e}_{[n_{0}]\backslash\mathcal{R}^{[r]}}^{(0)[r]}=0$. For any regression coefficient estimate $\boldsymbol{w}$, the loss function under our FARM setting on the $r$-th fold of target data is
$\hat{L}_{0}^{[r]}(\boldsymbol{w})=\frac{1}{n_{0}/3}\|{\boldsymbol{e}^{(0)[r]}}^{\top}(\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}\boldsymbol{w})\|_{2}^{2},$ (3.3)
The details are shown in Algorithm 2.
Input: target data $(\boldsymbol{X}^{(0)},\boldsymbol{Y}^{(0)})$, source data $\{(\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)})\}_{k=1}^{K}$, a constant $\epsilon_{0}>0$, penalty parameters $\{\{\lambda^{(k)[r]}\}_{k=0}^{K}\}_{r=1}^{3}$, constants $\epsilon_{0}$ and $\hat{\sigma}$.
Output: the estimated coefficient vector $\hat{\boldsymbol{\beta}}$, and the determined transferring set $\widehat{\mathcal{A}}$
1 Preprocessing step: Run step 1 (Preprocessing step) of Algorithm 1.
2 Transferable source detection: Randomly divide $(\boldsymbol{X}^{(0)},\boldsymbol{Y}^{(0)})$ into three sets of equal size as $\{(\boldsymbol{X}^{(0)[i]},\boldsymbol{Y}^{(0)[i]})\}_{i=1}^{3}$
3 for $r=1$ to 33 do
4 $\hat{\boldsymbol{\beta}}^{[r]}\leftarrow$ fit the Lasso on $\{(\boldsymbol{X}^{(0)[i]},\boldsymbol{Y}^{(0)[i]})\}_{i=1}^{3}\backslash(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]})$ with penalty parameter $\lambda^{(0)[r]}$
5 $\hat{\boldsymbol{w}}^{(k)[r]}\leftarrow$ run step 2 (transferring step) in Algorithm 1 with $(\{(\boldsymbol{X}^{(0)[i]},\boldsymbol{Y}^{(0)[i]})\}_{i=1}^{3}\backslash(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]}))\cup(\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)})$ and penalty parameter $\lambda^{(k)[r]}$ for all $k\neq 0$
6   Compute the loss $\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})$ for $k\neq 0$ and $\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})$ on $(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]})$.
7
8 end for
[…]
10 $\widehat{\mathcal{A}}\leftarrow\{k\neq 0:\hat{L}_{0}^{(k)}\leq\hat{L}_{0}^{(0)}+\epsilon_{0}\hat{\sigma}^{2}\}$
11 $\widehat{\mathcal{A}}-$Trans-GLM: $\hat{\boldsymbol{\beta}}\leftarrow$ run Algorithm 1 using $\{(\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)})\}_{k\in\{0\}\cup\widehat{\mathcal{A}}}$
12 Output $\hat{\boldsymbol{\beta}}$ and $\widehat{\mathcal{A}}$
Algorithm 2 Trans-FARM Algorithm
To ensure the consistency of the source detection algorithm, we impose some assumptions below.
Assumption 4.
  • (a) There exists a positive constant $C_{2}<\infty$ such that for every $k$,
    $\max_{i\in\{0,k\}}\left\|\left(\frac{n_{0}}{n_{0}+n_{k}}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}+\frac{n_{k}}{n_{0}+n_{k}}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\right)^{-1}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\right\|_{1}\leq C_{2}.$
    Moreover, assume that
    $\frac{\mathcal{V}_{n_{k},p}\|\boldsymbol{\varphi}^{(k)}\|_{2}}{\sqrt{(n_{k}+n_{0})\log p}}=O(1).$
  • (b) There exists a positive constant $\tilde{\eta}$ such that $\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{1}\leq\tilde{\eta}$ for $k\in\mathcal{A}_{\eta}^{c}$.
  • (c) Assume that for some $\eta$,
    $\epsilon_{0}$ $\gtrsim\frac{s\log p}{n_{0}}+\eta\sqrt{\frac{\log p}{n_{0}}}+\frac{\eta^{2}}{p}+s^{2}\left(\frac{\log p}{n_{0}}\right)^{\frac{3}{2}},\quad\text{and}$
    $\inf_{k\in\mathcal{A}_{\eta}^{c}}\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{2}^{2}$ $\gtrsim\epsilon_{0}\vee\left(\frac{s\log p}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}(\tilde{\eta}\vee\tilde{\eta}^{2})+\tilde{\eta}^{\frac{3}{2}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}+\tilde{\eta}\sqrt{\frac{s\log p}{n_{0}+n_{k}}}\right).$
  • (d) Suppose that $\sigma_{0}^{2}=\mbox{Var}(\mathcal{E}^{(0)})$ with $0<\sigma_{0}^{2}<\infty$. There exists a $0<\Delta_{\sigma}\leq 1$ such that $|\hat{\sigma}/\sigma_{0}-1|=O_{P}(\Delta_{\sigma})$, and $\Delta_{\sigma}=O(1)$.
Remark 3.2.
Assumption 4 (a)(b)(c) ensures that the transferring step can be applied to attain some estimator for the consistency of source detection. Similar conditions can be found in tian2023transfer; yang2025communication; yang2025debiased; zhang2025transfer. Assumption 4 (d) is assumed only for convenience. On one hand, as long as $\hat{\sigma}^{2}$ used in Algorithm 2 is the order of a constant, our source detection algorithm can be proved to be consistent. This type of choice of $\hat{\sigma}$ is similar to tian2023transfer; yang2025communication; yang2025debiased. On the other hand, the source detection is not sensitive to the choice of $\epsilon_{0}\hat{\sigma}^{2}$ in practice (yang2025debiased).
Under Assumption 4, the following result shows that our source detection algorithm is consistent.
Theorem 3.2.
Suppose that Assumptions 1-2 hold for $k\in[K]$ and Assumptions 34 also hold. In addition, assume that $\max_{k\in[K]}n_{k}=O(p)$, $\eta=O(1)$, and conditions (3.1)-(3.2) are satisfied. If we choose $\lambda^{(0)[r]}\asymp\sqrt{\log p/n_{0}}$ and $\lambda^{(k)[r]}\asymp\sqrt{\log p/(n_{0}+n_{k})}$, then $\mathbb{P}(\widehat{\mathcal{A}}=\mathcal{A}_{\eta})\to 1$.
3.3 Inference
3.3.1 Adequacy test of factor model
Factor regression is widely used across various fields as an effective dimensionality reduction tool. This naturally raises the question of whether the assumed factor structure is adequate for a given dataset, in which case the FAQR model provides a meaningful alternative. More specifically, we consider testing the following hypotheses:
$H_{0}:\boldsymbol{\beta}=0\quad\text{versus}\quad\boldsymbol{\beta}\neq 0$ (3.4)
for the target FARM in (2.1) with $k=0$ and $\boldsymbol{\beta}=\boldsymbol{w}^{(0)}$. We only consider high-dimensional simultaneous hypothesis tests for the $p$ features by leveraging the high-dimensional Gaussian approximation and bootstrap approaches (zhang2017simultaneous; fan2024latent; cai2025statistical) as it can also be easily extended to fixed-dimensional inference van2014asymptotically and any subset of $[p]$. The key is to attain a debiased-type estimator via constructing the inverse matrix of the Hessian matrix of the population loss function on the target dataset, i.e. $\boldsymbol{\Theta}=(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)})^{-1}$.
Note that the transfer learning is also extended to attain a better estimator of $\boldsymbol{\Theta}$ by using nodewise regression on the target dataset and the informative dataset $\widehat{\mathcal{A}}_{\eta}$ (tian2023transfer; yang2025communication; yang2025debiased). However, there is no guarantee that the informative sets for regression coefficients and design matrix are the same. Another problem is that the regressors $\hat{\boldsymbol{U}}^{(k)}$ in FARM are estimated such that estimating $\boldsymbol{\Theta}$ via transfer learning does not improve estimation accuracy. Hence, we construct the debiased estimator for $\boldsymbol{\beta}$ as
[…]
where $\hat{\boldsymbol{\beta}}$ is the estimator attained via Algorithm 1, and $\hat{\boldsymbol{\Theta}}$ is an estimator of $\boldsymbol{\Theta}$ by estimating the inverse matrix of $\hat{\boldsymbol{U}}^{(k)\top}\hat{\boldsymbol{U}}^{(k)}/n_{k}$. We do not constrain $\boldsymbol{\Theta}$ to be attained from any specific type, but only need to satisfy the following assumption.
Assumption 5.
There exist positive $\Delta_{1}$, $\Delta_{\max}$, $\Delta_{\infty}$, and $0<\Delta_{\sigma}\leq 1$ such that $|\hat{\sigma}/\sigma_{0}-1|=O_{P}(\Delta_{\sigma})$, and
$\|\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\|_{\max}=O_{P}(\Delta_{1}),\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\max}=O_{P}(\Delta_{\max}),\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\infty}=O_{P}(\Delta_{\infty}),$
where $\sigma_{0}^{2}=\mbox{Var}(\mathcal{E}^{(0)})$, and $\Delta_{1}$, $\Delta_{\max}$, $\Delta_{\infty}$, and $\Delta_{\sigma}$ satisfies
$\sqrt{n_{0}\log p}\Delta_{1}\mathcal{R}_{1}+\Delta_{\infty}\log p\to 0\quad\text{and}\quad\left(\mathcal{V}_{n_{0},p}\|\boldsymbol{\varphi}^{(0)}\|_{2}+\sqrt{\log p}\right)\|\boldsymbol{\Theta}\|_{\infty}\sqrt{\frac{\log p}{n_{0}}}\to 0,$
where $\mathcal{R}_{1}$ is the $\ell_{1}$-error bound for $\hat{\boldsymbol{\beta}}$ established in Theorem 3.1. We further assume that $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\infty}=O_{P}(\|\boldsymbol{\Theta}\|_{\infty})$ without loss of generality.
This assumption on the estimation of $\hat{\boldsymbol{\Theta}}$ can be easily satisfied for the CLIME-type estimator (cai2011constrained; javanmard2014confidence; yan2023confidence; cai2025statistical) and the nodewise regression (zhang2014confidence; van2014asymptotically; zhang2017simultaneous). For the assumption on the estimation of $\hat{\sigma}$, one can adopt the refitted cross-validation approach in fan2012variance or the scaled-Lasso (sun2012scaled) to attain a consistent estimation of $\sigma_{0}$ by following fan2024latent. More details can be seen in Section 3.1 of fan2024latent and references therein.
Theorem 3.3.
Suppose that $\hat{\boldsymbol{\beta}}$ satisfies the error bounds shown in 3.1, and Assumptions 1-2 hold for $k=0$ and Assumption 5 also holds, then there exists a $p$-dimensional Gaussian random vector $\boldsymbol{Q}$ with mean $\boldsymbol{0}_{p}$ and variance $\sigma^{2}\boldsymbol{\Theta}$ such that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\infty}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}\|_{\infty}\leq t)\right|=o(1).$
From Theorem 3.3, it is natural to attain the $(1-\alpha)$ quantile of $\boldsymbol{Q}$, denoted by $c_{1-\alpha}$, and then construct the test statistic as $\mathbb{I}(\sqrt{n_{0}}\|\tilde{\boldsymbol{\beta}}\|_{\infty}>c_{1-\alpha})$. However, when $p$ is large, it is still unclear whether the critical value can be reliably estimated by sampling from the $p$-dimensional Gaussian distribution with an estimated covariance matrix. Hence, we adopt the multiplier bootstrap procedure. Let $e_{i}\sim\mathbb{N}(0,1),i=1,\cdots,n_{0}$ be independent of the data $\mathcal{D}=\{(\boldsymbol{x}_{i}^{(0)},y_{i}^{(0)},\boldsymbol{f}_{i}^{(0)},\boldsymbol{u}_{i}^{(0)},\mathcal{E}_{i}^{(0)})\}_{i=1}^{n_{0}}$, and consider
$\hat{\boldsymbol{Q}}^{e}=\hat{\sigma}\hat{\boldsymbol{\Theta}}\frac{1}{\sqrt{n_{0}}}\sum_{i=1}^{n_{0}}e_{i}\hat{\boldsymbol{u}}_{i}^{(0)}.$
Denote the $\alpha$-quantile of $\|\hat{\boldsymbol{Q}}^{e}\|_{\infty}$ by
$\hat{c}(\alpha)=\inf_{t\in\mathbb{R}}\{t\in\mathbb{R}:\mathbb{P}_{e}(\|\hat{\boldsymbol{Q}}^{e}\|_{\infty}\leq t)\geq\alpha\},$
where $\mathbb{P}_{e}(\mathcal{H})$ represents the probability of the event $\mathcal{H}$ with respect to $e_{1},\cdots,e_{n}$. The validity of the bootstrap approach is established by the following theorem.
Theorem 3.4.
Under the conditions of Theorem 3.3, if
$\log p(\Delta_{1}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}+\Delta_{\sigma})=o(1),$
we have
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}^{e}\|_{\infty}\leq t|\mathcal{D}\right)-\mathbb{P}\left(\|\boldsymbol{Q}\|_{\infty}\leq t\right)\right|=o(1).$
Following Theorem 3.3 and Theorem 3.4, our test statistic for (3.4) is given by
$\mathbb{I}(\sqrt{n_{0}}\|\tilde{\boldsymbol{\beta}}\|_{\infty}>\hat{c}(1-\alpha)).$
Compared with the test statistic in fan2024latent, our procedure is more reliable because the estimator $\hat{\boldsymbol{\beta}}$ obtained from Algorithm 1 provides a more accurate estimate of $\boldsymbol{\beta}$. In particular, Assumption 5 implies that part of the inferential error captured by $\mathcal{R}_{1}$ can be substantially improved as long as $\mathcal{R}_{1}$ is smaller than the traditional $\ell_{1}$-error $s\sqrt{\log p/n_{0}}$ in fan2024latent.
3.3.2 Confidence interval construction
In this section, we provide methods to construct confidence intervals. By employing the same de-biasing estimator in (3.5) and invoking the results in Theorems 3.33.4, we can further construct both entrywise confidence intervals (van2014asymptotically; javanmard2014confidence; tian2023transfer) and groupwise simultaneous confidence intervals (zhang2017simultaneous; fan2024latent; yan2023confidence; cai2025statistical). We focus on the construction of groupwise simultaneous confidence intervals, as the entrywise intervals are simpler and arise as a direct byproduct of the groupwise procedure. Let $\mathcal{G}\subset[p]$ be the indices of parameters of interest, and denote $\|\boldsymbol{Q}\|_{\mathcal{G}}=\|\boldsymbol{Q}_{\mathcal{G}}\|_{\infty}$ as the maximum norm on the sub-vector $\boldsymbol{Q}_{\mathcal{G}}$. The following corollary presents the validity of constructing simultaneous confidence intervals for $\boldsymbol{\beta}_{\mathcal{G}}$ via bootstrap procedure.
Corollary 3.1.
Under the conditions of Theorem 3.4, there exists a $p$-dimensional Gaussian random vector $\boldsymbol{Q}$ with mean $\boldsymbol{0}_{p}$ and variance $\sigma^{2}\boldsymbol{\Theta}$ such that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\mathcal{G}}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}\|_{\mathcal{G}}\leq t)\right|=o(1),$
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}^{e}\|_{\mathcal{G}}\leq t|\mathcal{D}\right)-\mathbb{P}\left(\|\boldsymbol{Q}\|_{\mathcal{G}}\leq t\right)\right|=o(1).$
From Corollary 3.1, let the $\alpha$-quantile of $\|\hat{\boldsymbol{Q}}^{e}\|_{\mathcal{G}}$ by
$\hat{c}_{\mathcal{G}}(\alpha)=\inf_{t\in\mathbb{R}}\{t\in\mathbb{R}:\mathbb{P}_{e}(\|\hat{\boldsymbol{Q}}^{e}\|_{\mathcal{G}}\leq t)\geq\alpha\},$
the simultaneous $\alpha$-level confidence interval for $\boldsymbol{\beta}_{\mathcal{G}}$ can be constructed as
$\left[\tilde{\beta}_{i}-n_{0}^{-1/2}\hat{c}_{\mathcal{G}}(1-\alpha),\tilde{\beta}_{i}+n_{0}^{-1/2}\hat{c}_{\mathcal{G}}(1-\alpha)\right],i\in\mathcal{G}.$
[…]
$\hat{\boldsymbol{Q}}_{stu}^{e}=\hat{\sigma}\hat{\boldsymbol{S}}^{-1/2}\hat{\boldsymbol{\Theta}}\frac{1}{\sqrt{n_{0}}}\sum_{i=1}^{n_{0}}e_{i}\hat{\boldsymbol{u}}_{i}^{(0)},$
and denote the $\alpha$-quantile of $\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\mathcal{G}}$ by
$\hat{c}_{\mathcal{G},stu}(\alpha)=\inf_{t\in\mathbb{R}}\{t\in\mathbb{R}:\mathbb{P}_{e}(\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\mathcal{G}}\leq t)\geq\alpha\}.$
The validity of the studentized bootstrap procedure is established in the following theorem.
Theorem 3.5.
Under the conditions of Theorem 3.4, further assume that $\Delta_{\infty}\|\boldsymbol{\Theta}\|_{\infty}\sqrt{\log p/n_{0}}=o(1)$, then there exists a $p$-dimensional Gaussian random vector $\boldsymbol{Q}_{stu}$ with mean $\boldsymbol{0}_{p}$ and variance $\sigma^{2}\boldsymbol{S}^{-1/2}\boldsymbol{\Theta}\boldsymbol{S}^{-1/2}$ such that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n}\hat{\boldsymbol{S}}^{-1/2}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\mathcal{G}}\leq t\right)-\mathbb{P}(\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\mathcal{G}}\leq t)\right|=o(1),$
and
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\infty}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}_{stu}\|_{\infty}\leq t)\right|=o(1).$
Following Theorem 3.5, the simultaneous $\alpha$-level confidence interval with varying length for $\boldsymbol{\beta}_{\mathcal{G}}$ can be constructed as follows:
$\left[\tilde{\beta}_{i}-n_{0}^{-1/2}(\hat{\boldsymbol{\Theta}}_{i,i})^{1/2}\hat{c}_{\mathcal{G},stu}(1-\alpha),\tilde{\beta}_{i}+n_{0}^{-1/2}(\hat{\boldsymbol{\Theta}}_{i,i})^{1/2}\hat{c}_{\mathcal{G},stu}(1-\alpha)\right],i\in\mathcal{G}.$
The complete steps for obtaining $\tilde{\boldsymbol{\beta}}$ and constructing simultaneous confidence intervals with varying lengths are summarized in Algorithm 3. The corresponding procedure for the non-studentized version follows an analogous structure and is therefore omitted for brevity.
Input: target data $(\boldsymbol{X}^{(0)},\boldsymbol{Y}^{(0)})$, $\mathcal{G}$, estimations $\hat{\boldsymbol{\Theta}}$ and $\hat{\sigma}$.
Output: the simultaneous confidence interval for $\boldsymbol{\beta}_{\mathcal{G}}$
1 Run Algorihm 2 to attain $\hat{\boldsymbol{\beta}}$, and calculate $\tilde{\boldsymbol{\beta}}$ via (3.5).
2 for $\ell=1$ to $B$ do
3   Draw standard normal random variables $e_{\ell,i},i=1,\cdots,n_{0}$.
4   Calculate $\hat{\boldsymbol{Q}}_{stu,\ell}^{e}=\hat{\sigma}\hat{\boldsymbol{S}}^{-1/2}\hat{\boldsymbol{\Theta}}\sum_{i=1}^{n_{0}}e_{\ell,i}\hat{\boldsymbol{u}}_{i}^{(0)}$
5 end for
6Simultaneous confidence intervals construction: Find the $1-\alpha$-quantile of $\{\|\hat{\boldsymbol{Q}}_{stu,\ell}^{e}\|_{\mathcal{G}}\}_{\ell=1}^{B}$, represented as $\hat{c}_{\mathcal{G},stu}(1-\alpha)$. The simultaneous $\alpha$-level confidence interval for $\beta_{i}$ is $\mathcal{I}_{i}=\left[\tilde{\beta}_{i}-n_{0}^{-1/2}(\hat{\boldsymbol{\Theta}}_{i,i})^{1/2}\hat{c}_{\mathcal{G},stu}(1-\alpha),\tilde{\beta}_{i}+n_{0}^{-1/2}(\hat{\boldsymbol{\Theta}}_{i,i})^{1/2}\hat{c}_{\mathcal{G},stu}(1-\alpha)\right],i\in\mathcal{G}.$ Output the simultaneous confidence intervals $\{\mathcal{I}_{i}\}_{i\in\mathcal{G}}$.
Algorithm 3 Simultaneous confidence interval construction
4 Simulation
In this section, we provide numerical experiments to illustrate the finite sample properties of our proposed Trans-FARM estimator under the FARM. In the simulation part, we study the performance of different methods under various settings of $h$. The methods include only-FARM (Algorithm 1 with $\mathcal{A}$ being $\varnothing$), Trans-FARM (Algorithm 2), Oracle-Trans-FARM (Algorithm 1 with $\mathcal{A}=\mathcal{A}_{\eta}$), Pooled-Trans-FARM (Algorithm 1 with $\mathcal{A}$ being all sources), only-Lasso (Lasso on target data), Trans-Lasso (tian2023transfer, Algorithm 2), Oracle-Trans-Lasso (Algorithm 1 in (tian2023transfer) with $\mathcal{A}=\mathcal{A}_{\eta}$), Pooled-Trans-Lasso (Algorithm 1 in (tian2023transfer) with $\mathcal{A}$ being all sources). The $\epsilon_{0}\hat{\sigma}^{2}$ in Trans-FARM is set as $2\hat{L}_{0}^{(0)}$ (see Algorithm 2), similar to the setting in tian2023transfer. All experiments are conducted in R.
We generate the data from the FARM model as follows. We let $r_{0}=\ldots=r_{K}=2$, and the design matrix is generated by
$\boldsymbol{X}^{(k)}=\boldsymbol{F}_{k}\boldsymbol{B}_{k}^{\top}+\boldsymbol{U}_{k}\in\mathbb{R}^{n_{k}\times p}$
[…]
We set $n_{0}=300$, $n_{1}=\ldots=n_{K}=300$, $p=500$, $s=20$ and $\eta=5$ or $\eta=10$. We let $K=10$ and the informative set $\mathcal{A}_{\eta}$ is random chosen from $\{1,\cdots,K\}$ with $|\mathcal{A}_{\eta}|$ varies from 0 to $10$. We exhibit the $\ell_{2}$ errors of the $8$ estimators in Figure 1.
Figure 1: Comparison of $\ell_{2}$ errors under varying number of informative sets ($|\mathcal{A}_{\eta}|$) and $\eta$ with $K=10$, $n_{0}=n_{1}=\ldots=n_{K}=300$, $p=500$, and $s=20$, and $r_{\mathrm{p}}=2\times 10^{3}$.
From Figure 1, we draw several conclusions. First, across all experimental regimes, FARM-based estimators dominate Lasso-based estimators in tian2023transfer, consistent with the theoretical expectation that Lasso suffers from model misspecification in the presence of latent factors. Second, the Oracle-Trans-FARM procedure exhibits strictly smaller estimation error than the Target-only estimators, with the error decreasing monotonically in the number of informative sources, which aligns with the theoretical benefits of borrowing strength across similar datasets. Moreover, the benefits of Oracle-Trans-FARM compared with Oracle-Trans-Lasso increase as the number of informative sources increases, which states that Oracle-Trans-FARM can capture more information from the informative sources in this setting. Third, the errors of Trans-FARM align with the Oracle-Trans-FARM, indicating that our source detection algorithm accurately recovers the underlying transferable set. Finally, the poor performance of Pooled-Trans-FARM relative to Trans-FARM highlights the necessity of the source detection algorithm. These results collectively show that Trans-FARM achieves near-oracle accuracy while adaptively identifying informative sources.
5 Conclusion
In this paper, we study the factor-augmented sparse linear model within a transfer learning framework, aiming to address the challenge of strong correlations among predictors and to account for potential latent factor structures, along with leveraging these auxiliary datasets to improve estimation accuracy. This setting is particularly relevant to economic, financial, and medical applications, where the target dataset is often limited in size but accompanied by multiple heterogeneous auxiliary sources that can help enhance the estimation accuracy of regression coefficients. We propose transfer learning procedures that effectively exploit these auxiliary datasets and establish non-asymptotic error bounds for the resulting estimators under both $\ell_{1}$- and $\ell_{2}$- estimation errors. To mitigate the risk of negative transfer, we further develop a transferable source detection method in a data-driven manner and provide a thorough theoretical analysis of its consistency. We also adopt an inferential framework to test the adequacy of the factor model and construct valid simultaneous confidence intervals for regression coefficients of interest. The proposed simultaneous confidence intervals, constructed via a studentized bootstrap procedure, are especially noteworthy in that their lengths vary across coefficients, thereby reflecting the heterogeneous influence of different features on the response. Simulation experiments and a real-data study demonstrate the practical advantages of our methodology, showing that the proposed procedures yield substantial improvements in estimation accuracy, deliver reliable uncertainty quantification, and remain robust across heterogeneous environments.
Acknowledgments
Appendix A Proof
Denote
$\hat{L}(\boldsymbol{w})$ $=\frac{1}{2(n_{\mathcal{A}_{\eta}}+n_{0})}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}\|_{2}^{2},$
$\nabla\hat{L}(\boldsymbol{w})$ $=-\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\hat{\boldsymbol{U}}_{k}^{\top}(\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}),$
$\delta\hat{L}(\boldsymbol{t})$ $=\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}}+\boldsymbol{t})-\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})-{\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})}^{\top}\boldsymbol{t}.$
$D(\boldsymbol{\Delta})$ $=\boldsymbol{\Delta}^{\top}\left(\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}\right)\boldsymbol{\Delta}$
$\hat{D}(\boldsymbol{\Delta})$ $=\boldsymbol{\Delta}^{\top}\left(\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}\right)\boldsymbol{\Delta}$
Proof.
Let $\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}-\boldsymbol{w}^{\mathcal{A}_{\eta}}$ and $\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}$ be the bias of the the transferring step and eventual estimator respectively. We firstly consider the error bound for $\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$ from which we are committed to derive the error bound for $\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}$.
Step 1: Bounds for $\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$.
By the definition of $\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}$, Lemma A.1, and the convexity of $\hat{L}$, under the event $\mathcal{E}_{\boldsymbol{w}}=\{\lambda_{\boldsymbol{w}}\geq 2\|\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})\|_{\infty}\}$, it follows that
$0\leq\delta\hat{L}(\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$ $\leq\lambda_{\boldsymbol{w}}(\|\boldsymbol{w}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})-\lambda_{\boldsymbol{w}}(\|\hat{\boldsymbol{w}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\|\hat{\boldsymbol{w}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})-{\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})}^{\top}\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$ (A.1)
$\leq\lambda_{\boldsymbol{w}}(\|\boldsymbol{w}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})-\lambda_{\boldsymbol{w}}(\|\hat{\boldsymbol{w}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\|\hat{\boldsymbol{w}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})+\frac{1}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}$
$\leq\frac{3}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\lambda_{\boldsymbol{w}}(\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\|\hat{\boldsymbol{w}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})+\frac{1}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}$
$\leq\frac{3}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{1}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{w}}\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}$
$\leq\frac{3}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{1}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{w}}C_{1}\eta$
It can be implied from (A.1) that $\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}\leq 3\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta$ such that $\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\in\mathbb{C}_{\eta}$ where $\mathbb{C}_{\eta}=\{\boldsymbol{t}:\|\boldsymbol{t}_{\mathcal{S}^{c}}\|_{1}\leq 3\|\boldsymbol{t}_{\mathcal{S}}\|_{1}+4C_{1}\eta\}$. Meanwhile, it is straightforward that $t\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\in\mathbb{C}_{\eta}$ for any $t\in(0,1)$.
We term to consider the event
$\mathcal{E}_{\boldsymbol{w}}^{\prime}=\left\{\frac{D(\boldsymbol{\Delta})}{\|\boldsymbol{\Delta}\|_{2}^{2}}\geq\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\frac{\|\boldsymbol{\Delta}\|_{1}}{\|\boldsymbol{\Delta}\|_{2}}\right),\quad\text{for all }\|\boldsymbol{\Delta}\|_{2}\leq 1\right\}.$
We aim to show that conditional on $\mathcal{E}_{\boldsymbol{w}}\cap\mathcal{E}_{\boldsymbol{w}}^{\prime}$,
$\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\leq 8\kappa_{2}C_{1}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\eta+3\kappa_{1}^{-1}\sqrt{s}\lambda_{\boldsymbol{w}}+2\sqrt{\frac{\lambda_{\boldsymbol{w}}C_{1}\eta}{\kappa_{1}}}+\sqrt{\frac{2e_{\eta}}{\kappa_{1}}}:=c_{\mathcal{T}}$
through proof by contradiction, where $e_{\eta}$ is defined as $e_{\eta}=c\eta^{2}((|\mathcal{A}_{\eta}|+1)\log p/(n_{\mathcal{A}_{\eta}}+n_{0})+1/p)$ for some positive constant $c<\infty$. If this claim does not hold, we could find some $t_{0}\in(0,1)$ such that $\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}=t\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$ satisfying $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\geq c_{\mathcal{T}}$ and $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\leq 1$ and as long as $c_{\mathcal{T}}\leq 1$. Let $\tilde{\boldsymbol{w}}^{\mathcal{A}_{\eta}}=\boldsymbol{w}^{\mathcal{A}_{\eta}}+\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$. The optimality of $\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}$ implies that $F(\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\leq 0$ with $F(\boldsymbol{t})=\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}}+\boldsymbol{t})-\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})+\lambda_{\boldsymbol{w}}\|\boldsymbol{w}^{\mathcal{A}_{\eta}}+\boldsymbol{t}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}$. The convexity of function $F(\cdot)$ and $F(\boldsymbol{0}_{p})=0$ yield that $F(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\leq tF(\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\leq 0$. Following a same trick of (A.1), together with the convexity of function $F(\cdot)$ yield that
$F(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\geq\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})^{\top}\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})+\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}$
$\geq$ $-\frac{1}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$
$\geq$ $-\frac{1}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{3}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\lambda_{\boldsymbol{w}}(\|\tilde{\boldsymbol{w}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$
$\geq$ $\frac{1}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{3}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}-2\lambda_{\boldsymbol{w}}\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}+D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})-D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$
$\geq$ $\frac{1}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{3}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}-2\lambda_{\boldsymbol{w}}C_{1}\eta+\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\frac{\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}}{\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}}\right)\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}^{2}+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})-D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$
$\geq$ $\kappa_{1}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}^{2}-\kappa_{1}\kappa_{2}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}-\frac{3}{2}\sqrt{s}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{2}-2\lambda_{\boldsymbol{w}}C_{1}\eta-|\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})-D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})|$
Note that $\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\in\mathbb{C}_{\eta}$, which deduces that
$\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}\leq 4\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta\leq 4\sqrt{s}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{2}+4C_{1}\eta,$
and by Lemma A.6, it is straightforward that
$|\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})-D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})|$ $\leq\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}^{2}\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\left\|\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}(\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}-\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k})\right\|_{\max}$
$=O_{P}\left(\left(s\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}^{2}+\eta^{2}\right)\left(\frac{(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}+\frac{1}{p}\right)\right)$
Hence, we have
$\frac{1}{2}\kappa_{1}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}^{2}-\left(4\kappa_{1}\kappa_{2}C_{1}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\eta+\frac{3}{2}\sqrt{s}\lambda_{\boldsymbol{w}}\right)\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}-2\lambda_{\boldsymbol{w}}C_{1}\eta-e_{\eta}\leq 0$
as long as $16\kappa_{2}\sqrt{s\log p/(n_{\mathcal{A}_{\eta}}+n_{0})}\leq 1$ and $s(|\mathcal{A}_{\eta}|+1)\log p/(n_{\mathcal{A}_{\eta}}+n_{0})+s/p<c$ for a constant $c$ small enough. However, the left-hand side of the inequality would be positive if the claim $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\geq c_{\mathcal{T}}$ holds. This leads to a contradiction, implying that $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\leq c_{\mathcal{T}}$, i.e.,
$\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}$ $\lesssim\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\eta+\sqrt{s}\lambda_{\boldsymbol{w}}+\sqrt{\lambda_{\boldsymbol{w}}\eta}+\eta\sqrt{\frac{(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\frac{\eta}{\sqrt{p}}$
$\lesssim\eta\sqrt{\frac{(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\sqrt{s}\lambda_{\boldsymbol{w}}+\sqrt{\lambda_{\boldsymbol{w}}\eta}+\frac{\eta}{\sqrt{p}}$
By choosing $\lambda_{\boldsymbol{w}}\asymp\sqrt{\log p/(n_{\mathcal{A}_{\eta}}+n_{0})}$, then we obatin
$\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\lesssim\eta\sqrt{\frac{(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\sqrt{\frac{s\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\left(\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}$
and
$\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}$ $\leq 4\sqrt{s}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{2}+4C_{1}\eta$
$\lesssim\eta\sqrt{\frac{s(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+s\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\left(\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}\right)^{\frac{1}{4}}\sqrt{s\eta}+\frac{\sqrt{s}\eta}{\sqrt{p}}+\eta$
$\lesssim\eta\sqrt{\frac{s(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+s\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\eta$
It remains to show that the probability of $\mathcal{E}_{\boldsymbol{w}}\cap\mathcal{E}_{\boldsymbol{w}}^{\prime}$ occurs with high probability. Lemma A.2 ensures that $\mathcal{E}_{\boldsymbol{w}}^{\prime}$ occurs with high probability. Lemma A.4 ensures that $\mathcal{E}_{\boldsymbol{w}}$ occurs with high probability if $\eta=O(1)$ and
$\frac{(|\mathcal{A}_{\eta}|+1)(1\vee\eta\log^{\frac{1}{2}}p\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2}\log^{-\frac{1}{2}}p)}{\sqrt{n_{\mathcal{A}_{\eta}}+n_{0}}}=O(1).$
Step 2: Bounds for $\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}$. Define
$\hat{L}^{(0)}(\boldsymbol{w})$ $=\frac{1}{2n_{0}}\|\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}\boldsymbol{w}\|_{2}^{2},\quad\nabla\hat{L}^{(0)}(\boldsymbol{w})=\frac{1}{n_{0}}\hat{\boldsymbol{U}}_{0}^{\top}(\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}\boldsymbol{w}),$
$\boldsymbol{\delta}^{\mathcal{A}_{\eta}}$ $=\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}},\quad\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}},\quad\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta},$
$\delta\hat{L}^{(0)}(\boldsymbol{\Delta})$ $=\hat{L}^{(0)}(\boldsymbol{\beta}+\boldsymbol{\Delta})-\hat{L}^{(0)}(\boldsymbol{\beta})-{\nabla\hat{L}^{(0)}(\boldsymbol{\beta})}^{\top}\boldsymbol{\Delta}.$
Similar to (A.1), under the event $\mathcal{E}_{\boldsymbol{\delta}}=\{\lambda_{\boldsymbol{\delta}}\geq 2\|\nabla\hat{L}^{(0)}(\boldsymbol{\beta})\|_{\infty}\}$, we have
$0\leq\delta\hat{L}^{(0)}(\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})$ $\leq\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}\|_{1}+\frac{1}{2}\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}$ (A.2)
$\leq\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}\|_{1}+\frac{1}{2}\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}$
$\leq\frac{3}{2}\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\Delta}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{\delta}}\|(\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})_{\mathcal{S}^{c}}\|_{1}-\frac{1}{2}\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\Delta}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}$
Since $\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}=\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}=(\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}})-\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$ with $\|\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}\leq C_{1}\eta$ by Lemma A.1, we conclude that
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}\leq 4\|\hat{\boldsymbol{\Delta}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+4(\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+C_{1}\eta)\leq 4\sqrt{s}\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}+4\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta.$
We term to consider the event
$\mathcal{E}_{\boldsymbol{\delta}}^{\prime}=\left\{\frac{D^{(0)}(\boldsymbol{\Delta})}{\|\boldsymbol{\Delta}\|_{2}^{2}}\geq\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\frac{\|\boldsymbol{\Delta}\|_{1}}{\|\boldsymbol{\Delta}\|_{2}}\right),\quad\text{for all }\|\boldsymbol{\Delta}\|_{2}\leq 1\right\},$
where $D^{(0)}(\boldsymbol{\Delta})=\boldsymbol{\Delta}^{\top}\left(\boldsymbol{U}_{0}^{\top}\boldsymbol{U}_{0}/n_{k}\right)\boldsymbol{\Delta}$.
We aim to show that conditional on $\mathcal{E}_{\boldsymbol{\delta}}\cap\mathcal{E}_{\boldsymbol{\delta}}^{\prime}$,
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\leq 8\kappa_{2}C_{1}\sqrt{\frac{\log p}{n_{0}}}\eta+8\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+\sqrt{\frac{2e_{\eta}^{\prime}}{\kappa_{1}}}:=c_{\boldsymbol{\Delta}}$
through proof by contradiction, where $e_{\eta}^{\prime}$ is defined as $e_{\eta}^{\prime}=c(\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}^{2}+\eta^{2})\left(\log p/n_{0}+1/p\right)+2\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{\delta}}C_{1}\eta$ for some positive constant $c<\infty$. If this claim does not hold, we could find some $t_{0}\in(0,1)$ such that $\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}=t\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}$ satisfying $\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\geq c_{\boldsymbol{\Delta}}$ and $\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\leq 1$ and as long as $c_{\boldsymbol{\Delta}}\leq 1$. Also, $\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}\leq 4\sqrt{s}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}+4\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta.$
Similar to Step 1, define $\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})=\boldsymbol{\Delta}^{\top}\left(\hat{\boldsymbol{U}}_{0}^{\top}\hat{\boldsymbol{U}}_{0}/n_{k}\right)\boldsymbol{\Delta}$, then we have
$|\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})-D^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})|$ $\leq\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}^{2}\frac{1}{n_{0}}\left\|\boldsymbol{U}_{0}^{\top}\boldsymbol{U}_{0}-\hat{\boldsymbol{U}}_{0}^{\top}\hat{\boldsymbol{U}}_{0}\right\|_{\max}$
$=O_{P}\left(\left(s\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}^{2}+\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}^{2}+\eta^{2}\right)\left(\frac{\log p}{n_{0}}+\frac{1}{p}\right)\right).$
Write $\tilde{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}=\boldsymbol{\beta}+\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}$. The optimality of $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}$ implies that $G(\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})\leq 0$ with $G(\boldsymbol{t})=\hat{L}^{(0)}(\boldsymbol{\beta}+\boldsymbol{t})-\hat{L}^{(0)}(\boldsymbol{\beta})+\lambda_{\boldsymbol{w}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\boldsymbol{t}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}$. The convexity of function $G(\cdot)$ and $G(\boldsymbol{0}_{p})=0$ yield that $G(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})\leq tG(\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})\leq 0$. Following a same trick of Step 1, conditional on $\mathcal{E}_{\boldsymbol{\delta}}\cap\mathcal{E}_{\boldsymbol{\delta}}^{\prime}$, one attains
$G(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})\geq\nabla\hat{L}^{(0)}(\boldsymbol{\beta})^{\top}\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}+\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})+\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}$
$\geq$ $-\frac{1}{2}\lambda_{\boldsymbol{\delta}}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}+\lambda_{\boldsymbol{\delta}}\|\tilde{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}+\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})$
$\geq$ $-2\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}+D^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})-|\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})-D^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})|$
$\geq$ $-2\lambda_{\boldsymbol{\delta}}(\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+C_{1}\eta)+\kappa_{1}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}^{2}-\kappa_{1}\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}-|\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})-D^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})|$
It follows that
$\frac{1}{2}\kappa_{1}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}^{2}-\left(4\kappa_{1}\kappa_{2}C_{1}\sqrt{\frac{\log p}{n_{0}}}\eta+4\kappa_{1}\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}\right)\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}-e_{\eta}^{\prime}\leq 0$
as long as $s\log p/n_{0}+s/p\lesssim c$ for a constant $c$ small enough, where $e_{\eta}^{\prime}=c(\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}^{2}+\eta^{2})\left(\log p/n_{0}+1/p\right)+2\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{\delta}}C_{1}\eta$ for some positive constant $c<\infty$. However, the left-hand side of the inequality would be positive if the claim $\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\geq c_{\boldsymbol{\Delta}}$ holds. This leads to a contradiction, implying that $\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\leq c_{\boldsymbol{\Delta}}$, i.e.,
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\lesssim$ $\sqrt{\frac{\log p}{n_{0}}}\eta+\sqrt{\frac{\log p}{n_{0}}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+\sqrt{e_{\eta}^{\prime}}$
$\lesssim$ $\sqrt{\frac{\log p}{n_{0}}}\eta+\frac{s\log p}{\sqrt{n_{0}(n_{\mathcal{A}_{\eta}}+n_{0})}}+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\frac{s}{\sqrt{p}}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}$
as long as $s(|\mathcal{A}_{\eta}|+1)\log p/(n_{\mathcal{A}_{\eta}}+n_{0})=O(1)$. Furthermore, assume that $(s\log p/n_{0})^{\frac{1}{2}}=O(1)$, we have
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\lesssim\sqrt{\frac{\log p}{n_{0}}}\eta+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}.$
Meanwhile,
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}$ $\leq 4\sqrt{s}\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}+4\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta$
$\lesssim s\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\eta+\sqrt{s\eta}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}$
It remains to show that the probability of $\mathcal{E}_{\boldsymbol{\delta}}\cap\mathcal{E}_{\boldsymbol{\delta}}^{\prime}$ occurs with high probability. Lemma A.3 ensures that $\mathcal{E}_{\boldsymbol{w}}^{\prime}$ occurs with high probability. Lemma A.5 ensures that $\mathcal{E}_{\boldsymbol{w}}$ occurs with high probability under conditions (3.1) and (3.2). ∎
Proof of Theorem 2.
We write $\hat{\boldsymbol{U}}_{0}^{[r]}=\boldsymbol{e}^{(0)[r]\top}\hat{\boldsymbol{U}}_{0}$, and
$L_{0}^{[r]}(\boldsymbol{w})=\mathbb{E}\left[\frac{1}{n_{0}/3}\|\boldsymbol{e}^{(0)[r]\top}(\boldsymbol{Y}^{(0)}-\boldsymbol{U}_{0}\boldsymbol{w}-\boldsymbol{F}_{0}\boldsymbol{\gamma}^{(k)})\|_{2}^{2}\right]$
for convenience. First, we consider $k\in\mathcal{A}_{\eta}$. For $r\in\{1,2,3\}$, $\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})$ can be decomposed as
$\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})=$ $\{\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})+L_{0}^{[r]}(\boldsymbol{\beta})\}+\{L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-L_{0}^{[r]}(\boldsymbol{\beta})\}$
$+\{\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})-\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-L_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+L_{0}^{[r]}(\boldsymbol{\beta})\}+\{L_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})-L_{0}^{[r]}(\boldsymbol{\beta})\}$
$=$ $I_{1}+I_{2}+I_{3}+I_{4}.$
For $I_{2}$, we have
$|I_{2}|=|L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-L_{0}^{[r]}(\boldsymbol{\beta})|=|\frac{1}{2}(\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta})^{\top}\mathbb{E}\left[\frac{\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}}{n_{0}/3}\right](\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta})|\lesssim\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{2}^{2}$
by the definition of $\boldsymbol{\beta}$. Similarly, we have $|I_{4}|\lesssim\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}^{2}$. By Lemma A.7, we have
$\max\{|I_{1}|,|I_{3}|\}\lesssim$ $\sqrt{\frac{\log p}{n_{0}}}(\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}\vee\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{1}\vee\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}^{2}\vee\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{2}^{2})$
$+\frac{\log p}{n_{0}}(\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}^{2}\vee\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{1}^{2}).$
By a similar proof of Theorem 3.1, one attains
$\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}\lesssim\sqrt{\frac{s\log p}{n_{0}}},\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}\lesssim s\sqrt{\frac{\log p}{n_{0}}},$
and
$\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{2}$ $\lesssim\sqrt{\frac{\log p}{n_{0}}}\eta+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{0}+n_{k}}},$
$\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{1}$ $\lesssim s\sqrt{\frac{\log p}{n_{0}+n_{k}}}+\eta+\sqrt{s\eta}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}$
under Assumptions 1-4. Hence, as long as $\eta=O(1)$, we conclude that
$|I_{1}|+|I_{2}|+|I_{3}|+|I_{4}|\lesssim$ $\frac{s\log p}{n_{0}}+\eta\sqrt{\frac{\log p}{n_{0}}}+\frac{\eta^{2}}{p}+\left(s\sqrt{\frac{\log p}{n_{0}}}+\eta\right)\sqrt{\frac{\log p}{n_{0}}}+\frac{\log p}{n_{0}}\left(s\sqrt{\frac{\log p}{n_{0}}}+\eta\right)^{2}$
$\lesssim$ $\frac{s\log p}{n_{0}}+\eta\sqrt{\frac{\log p}{n_{0}}}+\frac{\eta^{2}}{p}+s^{2}\left(\frac{\log p}{n_{0}}\right)^{\frac{3}{2}}$
Therefore, under Assumption 4 (d), we have
$\mathbb{P}$ $\left(\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})\geq\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+\epsilon_{0}\hat{\sigma}^{2}\right)\leq\mathbb{P}\left(\epsilon_{0}\hat{\sigma}^{2}\leq|I_{1}|+|I_{2}|+|I_{3}|+|I_{4}|\right)$
$\leq\mathbb{P}\left(\epsilon_{0}\hat{\sigma}^{2}\lesssim\frac{s\log p}{n_{0}}+\eta\sqrt{\frac{\log p}{n_{0}}}+\frac{\eta^{2}}{p}+s^{2}\left(\frac{\log p}{n_{0}}\right)^{\frac{3}{2}}\right)\to 0.$
Now we term to $k\in\mathcal{A}_{\eta}^{c}$. Let
$\boldsymbol{w}^{(0,k)}=\mathop{\arg\min}_{\boldsymbol{w}\in\mathbb{R}^{p}}\mathbb{E}\left[\frac{1}{2n_{0}/3+n_{k}}\left(\|(\boldsymbol{1}_{n_{0}}-\boldsymbol{e}^{(0)[r]})^{\top}(\boldsymbol{Y}^{(0)}-\boldsymbol{U}_{0}\boldsymbol{w}-\boldsymbol{F}_{0}\boldsymbol{\gamma}^{(k)})\|_{2}^{2}+\|\boldsymbol{Y}^{(k)}-\boldsymbol{U}_{k}\boldsymbol{w}-\boldsymbol{F}_{k}\boldsymbol{\gamma}^{(k)}\|_{2}^{2}\right)\right].$
For $r\in\{1,2,3\}$, $\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})$ can be decomposed as
$\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})$
$=$ $\{\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\boldsymbol{w}^{(0,k)})-L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})+L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})\}+\{L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})\}$
$+\{\hat{L}_{0}^{[r]}(\boldsymbol{w}^{(0,k)})-\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})+L_{0}^{[r]}(\boldsymbol{\beta})\}+\{L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})-L_{0}^{[r]}(\boldsymbol{\beta})\}$
$+\{\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})-\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-L_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+L_{0}^{[r]}(\boldsymbol{\beta})\}+\{L_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})-L_{0}^{[r]}(\boldsymbol{\beta})\}$
$=$ $E_{1}+E_{2}+E_{3}+E_{4}+E_{5}+E_{6}.$
Similar to the proof for $k\in\mathcal{A}_{\eta}$ case, we have
$\max\{|E_{1}|,|E_{3}|,|E_{5}|\}\lesssim$ $\left(\sqrt{\frac{\log p}{n_{0}}}(\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{1}\vee\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{1}\vee 1\right)\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{1}\sqrt{\frac{\log p}{n_{0}}}$
$+\left(\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{1}\vee 1\right)\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{1}\sqrt{\frac{\log p}{n_{0}}}$
$+\left(\sqrt{\frac{\log p}{n_{0}}}\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}\vee 1\right)\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}\sqrt{\frac{\log p}{n_{0}}},$
$+\sqrt{\frac{\log p}{n_{0}}}(\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}^{2}\vee\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}^{2}\vee\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}^{2})$
and $|E_{6}|\lesssim\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}^{2}$. For $E_{2}$, we have
$|E_{2}|$ $\leq|\mathbb{E}[\nabla L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})]^{\top}(\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)})+\frac{1}{2}v_{0}\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}^{2}$
$\leq v_{0}\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}+\frac{1}{2}v_{0}\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}^{2}.$
For $E_{4}$, we obtain $E_{4}\geq\lambda_{\min}(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)})\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}^{2}:=\underline{\kappa}\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}^{2}$. Assumption 4 ensures that we have
$\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}$ $\lesssim\sqrt{\frac{\log p}{n_{0}}}\tilde{\eta}+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\tilde{\eta}}+\frac{\tilde{\eta}}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{0}+n_{k}}},$
$\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{1}$ $\lesssim s\sqrt{\frac{\log p}{n_{0}+n_{k}}}+\tilde{\eta}+\sqrt{s\tilde{\eta}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}$
with probability at least $1-C\exp(-Cn_{0})$. Moreover, by a similar proof of Lemma A.1, we have
$\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}\leq\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{1}\lesssim\tilde{\eta}$.
Therefore, we have
$|E_{1}|+|E_{2}|+|E_{3}|+|E_{5}|+|E_{6}|\lesssim\frac{s\log p}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}(\tilde{\eta}\vee\tilde{\eta}^{2})+\tilde{\eta}^{\frac{3}{2}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}+\tilde{\eta}\sqrt{\frac{s\log p}{n_{0}+n_{k}}}.$
Hence, we conclude that for $k\in\mathcal{A}_{\eta}^{c}$,
$\mathbb{P}$ $\left(\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})\leq\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+\epsilon_{0}\hat{\sigma}^{2}\right)\leq\mathbb{P}\left(E_{4}\leq\epsilon_{0}\hat{\sigma}^{2}+|E_{1}|+|E_{2}|+|E_{3}|+|E_{5}|+|E_{6}|\right).$
As long as
$\inf_{k\in\mathcal{A}_{\eta}^{c}}\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{2}^{2}\gtrsim\epsilon_{0}\vee\left(\frac{s\log p}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}(\tilde{\eta}\vee\tilde{\eta}^{2})+\tilde{\eta}^{\frac{3}{2}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}+\tilde{\eta}\sqrt{\frac{s\log p}{n_{0}+n_{k}}}\right),$
the probability approaches $1$. Note that by union bounds, we have
$\mathbb{P}(\widehat{\mathcal{A}}\neq\mathcal{A}_{\eta})\leq\sum_{r=1}^{3}\sum_{k\in\mathcal{A}_{\eta}}\mathbb{P}\left(\hat{L}_{0}(\hat{\boldsymbol{w}}^{(k)[r]})\geq\hat{L}_{0}(\hat{\boldsymbol{\beta}}^{[r]})+\epsilon_{0}\hat{\sigma}^{2}\right)+\sum_{r=1}^{3}\sum_{k\in\mathcal{A}_{\eta}^{c}}\mathbb{P}\left(\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})\leq\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+\epsilon_{0}\hat{\sigma}^{2}\right).$
Here, the probability above tends to 0 since we can change all the error bounds previously analyzed that contained $\log p$ into $\log(Kp)\lesssim(\log K)\vee(\log p)$ and we assume that $K\lesssim p$ without loss of generality. ∎
Proof of Theorem 3.3.
Firstly, define $\boldsymbol{\gamma}_{j}^{(0)}={\arg\min}_{\boldsymbol{\gamma}\in\mathbb{R}^{p-1}}\mathbb{E}[(\boldsymbol{u}_{j}^{(0)}-\boldsymbol{u}_{-j}^{(0)}\boldsymbol{\gamma})^{2}]=(\boldsymbol{\Sigma}_{\boldsymbol{u},-j,-j}^{(0)})^{-1}\boldsymbol{\Sigma}^{(0)}_{\boldsymbol{u},-j,j}$ and $\boldsymbol{\omega}_{j}^{(0)}$ be the $p$-dimensional vector with elements $(\boldsymbol{\omega}_{j}^{(0)})_{j}=1$ and $(\boldsymbol{\omega}_{j}^{(0)})_{\ell}=-\gamma_{j\ell}^{(0)}$ for $j\neq\ell$. The following example states the specific form of $\Delta_{1}$, $\Delta_{\max}$, and $\Delta_{\infty}$ on Assumption 5.
Example 1 ((fan2024latent, Proposition E.1)).
We have
$\|\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{\Sigma}}_{\boldsymbol{u}}^{(0)}\|_{\max}$ $=O_{P}\left(\sqrt{\frac{\log p}{n_{0}}}+\frac{1}{\sqrt{p}}+\max_{j\in[p]}\frac{\mathcal{V}_{n_{0},p}}{n_{0}}\|\boldsymbol{B}_{0}^{\top}\boldsymbol{\omega}_{j}^{(0)}\|_{2}\right),$
$\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\max}$ $=O_{P}\left(\max_{j\in[p]}\sqrt{|\mathcal{S}_{j}|\left(\frac{\log p}{n_{0}}+\frac{1}{p}\right)}+\max_{j\in[p]}\frac{\mathcal{V}_{n_{0},p}}{n_{0}}\sqrt{|\mathcal{S}_{j}|}\|\boldsymbol{B}_{0}^{\top}\boldsymbol{\omega}_{j}^{(0)}\|_{2}\right),$
$\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\infty}$ $=O_{P}\left(\max_{j\in[p]}|\mathcal{S}_{j}|\sqrt{\frac{\log p}{n_{0}}+\frac{1}{p}}+\max_{j\in[p]}\frac{\mathcal{V}_{n_{0},p}}{n_{0}}|\mathcal{S}_{j}|\|\boldsymbol{B}_{0}^{\top}\boldsymbol{\omega}_{j}^{(0)}\|_{2}\right).$
First, we have the following decomposition:
[…]
The last term in (A.3) can be bounded as
$\|(\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{\Sigma}}_{\boldsymbol{u}})(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\max}\leq\|\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{\Sigma}}_{\boldsymbol{u}}\|_{\max}\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{1}=\Delta_{1}\mathcal{R}_{1},$
where
$\mathcal{R}_{1}=\sqrt{\frac{\log p}{n_{0}}}\eta+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}$
is the $\ell_{1}$ error bound established in Theorem 3.1. Additionally, we have
$\|\frac{1}{n_{0}}\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{U}}_{0}^{\top}\boldsymbol{F}_{0}\boldsymbol{\mathcal{\varphi}}^{(0)}\|_{\infty}=O_{P}(\mathcal{V}_{n_{0},p}\|\boldsymbol{\Theta}\|_{\infty}\|\boldsymbol{\mathcal{\varphi}}^{(0)}\|_{2}/n_{0}),$
by Lemma A.6 and
[…]
$=O_{P}\left(\frac{\sqrt{\log p}}{n_{0}}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}\sqrt{\frac{\log p}{n_{0}}}\right).$
[…]
$\Delta_{n_{0},p}$ $=\sqrt{n_{0}}\Delta_{1}\mathcal{R}_{1}+\frac{\mathcal{V}_{n_{0},p}\|\boldsymbol{\Theta}\|_{\infty}\|\boldsymbol{\mathcal{\varphi}}^{(0)}\|_{2}}{\sqrt{n_{0}}}+\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}\sqrt{\log p}=o\left(1\right).$
Additionally, since $\lambda_{\max}(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)})\leq\|\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\|_{1}\leq 1/\iota$, we have $\min_{j\in[p]}\boldsymbol{\Theta}_{j,j}\geq\lambda_{\min}(\boldsymbol{\Theta})\geq\iota$. Then, $\{\boldsymbol{\Theta}_{j}\boldsymbol{u}_{i}^{(0)}\mathcal{E}_{i}^{(0)}\}_{i=1}^{n_{0}}$ are i.i.d. zero-mean sub-exponential random variables with $\sigma^{2}\boldsymbol{\Theta}_{j,j}\geq\sigma^{2}\iota$ and
$\max_{j\in[p]}\|\boldsymbol{\Theta}_{j}\boldsymbol{u}_{i}^{(0)}\mathcal{E}_{i}^{(0)}\|_{\psi_{1}}\leq\frac{\|\boldsymbol{u}_{i}^{(0)}\|_{\psi_{1}}\|\mathcal{E}_{i}^{(0)}\|_{\psi_{1}}}{\iota}.$
By Lemma A.8, we have
[…]
By Nazarov’s inequality (chernozhukov2023high, Proposition 1), since the variance of each element in $\boldsymbol{Q}$ is bounded away from 0 and $\infty$ by assumptions, we can infer that
$\sup_{t\in\mathbb{R}}\mathbb{P}\left(|\|\boldsymbol{Q}\|_{\infty}-t|<\delta\log^{-1/2}(p)\right)\leq C\delta\log^{-1/2}(p)\{\sqrt{2\log p}+2\}.$
By choosing a $\delta$ small enough, the right hand side of the inequality can be arbitrarily small. From Theorem 3.4, the remainder is $o_{P}(1)$, and thus applying Lemma A.9,
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n_{0}}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\infty}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}\|_{\infty}\leq t)\right|=o(1).$
Proof of Theorem 3.4.
[…]
$\|\boldsymbol{\Upsilon}-\hat{\boldsymbol{\Upsilon}}\|_{\max}$ $\leq\hat{\sigma}^{2}\|\hat{\boldsymbol{\Theta}}\|_{\infty}\|\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{\Sigma}}_{\boldsymbol{u}}^{(0)}\|_{\max}+\hat{\sigma}^{2}\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\infty}\|\hat{\boldsymbol{\Theta}}\|_{\max}+|\sigma^{2}-\hat{\sigma}^{2}|\|\boldsymbol{\Theta}\|_{\max}$
$=O_{P}(\Delta_{1}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}+\Delta_{\sigma})$
Thus, $\log p\|\sigma^{2}\boldsymbol{\Theta}-\hat{\boldsymbol{\Upsilon}}\|_{\max}=o_{P}(1)$ by assumptions. Since the smallest eigenvalue of $\sigma^{2}\boldsymbol{\Theta}$ is bounded away from 0, by the Gaussian comparision result (chernozhukov2023nearly, Lemma 2.1), it can be derived that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}^{e}\|_{\infty}\leq t|\mathcal{D}\right)-\mathbb{P}\left(\|\boldsymbol{Q}\|_{\infty}\leq t\right)\right|=o(1).$
Proof of Corollary 3.1.
The proof simply follows from the proof of Theorems 3.3-3.4 by considering the subset $\mathcal{G}$. ∎
Proof of Theorem 3.5.
[…]
$\|\hat{\boldsymbol{S}}-\boldsymbol{S}\|_{\infty}\leq\|\boldsymbol{\Theta}-\hat{\boldsymbol{\Theta}}\|_{\infty}=O_{P}(\Delta_{\infty}).$
Since $\iota\leq\lambda_{\min}(\boldsymbol{\Theta})\leq\min_{j\in[p]}\boldsymbol{\Theta}_{j,j}\leq\max_{j\in[p]}\boldsymbol{\Theta}_{j,j}\leq\lambda_{\max}(\boldsymbol{\Theta})\leq\iota^{-1}$, which implies that the diagonal elements of $\boldsymbol{\Theta}$ are bounded away from 0, it can be inferred that $\|\hat{\boldsymbol{S}}^{-1/2}-\boldsymbol{S}^{-1/2}\|_{\infty}=O_{P}(\Delta_{\infty}).$ Then it can be concluded that
[…]
$=$ $O_{P}\left(\Delta_{\infty}\|\boldsymbol{\Theta}\|_{\infty}\sqrt{\frac{\log p}{n_{0}}}+\Delta_{n_{0},p}\right),$
where $\Delta_{n_{0},p}$ is defined in the proof of Theorem 3.3. The error bound above is $o_{P}(1)$ by assumptions. Let $\boldsymbol{Q}^{stu}$ be a $p$-dimensional Gaussian random vector with mean $\mathbf{0}$ and variance $\sigma^{2}\boldsymbol{S}^{-1/2}\boldsymbol{\Theta}\boldsymbol{S}^{-1/2}$. Similar to the proof of Theorem 3.3, it can be verified that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n}\hat{\boldsymbol{S}}^{-1/2}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\mathcal{G}}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}^{stu}\|_{\mathcal{G}}\leq t)\right|=o(1).$
Following the proof and notation of Theorem 3.4, we have $\hat{\boldsymbol{Q}}^{e}_{stu}\sim\mathbb{N}(\boldsymbol{0}_{p},\hat{\boldsymbol{S}}^{-1/2}\hat{\boldsymbol{\Upsilon}}\hat{\boldsymbol{S}}^{-1/2})$. Since $\hat{\boldsymbol{S}}^{-1/2}$ and $\boldsymbol{S}^{-1/2}$ are bounded diagonal matrices, it can be deduced that
$\|\hat{\boldsymbol{S}}^{-1/2}\hat{\boldsymbol{\Upsilon}}\hat{\boldsymbol{S}}^{-1/2}-\boldsymbol{S}^{-1/2}\boldsymbol{\Upsilon}\boldsymbol{S}^{-1/2}\|_{\max}\leq$ $\|\hat{\boldsymbol{S}}^{-1/2}-\boldsymbol{S}^{-1/2}\|_{\max}\|\hat{\boldsymbol{\Upsilon}}\|_{\max}\|\hat{\boldsymbol{S}}^{-1/2}\|_{\max}$
$+\|\boldsymbol{S}^{-1/2}\|_{\max}\|\hat{\boldsymbol{\Upsilon}}-\hat{\boldsymbol{\Upsilon}}\|_{\max}\|\hat{\boldsymbol{S}}^{-1/2}\|_{\max}$
$+\|\boldsymbol{S}^{-1/2}\|_{\max}\|\boldsymbol{\Upsilon}\|_{\max}\|\hat{\boldsymbol{S}}^{-1/2}-\boldsymbol{S}^{-1/2}\|_{\max}$
$=$ $O_{P}(\Delta_{1}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}+\Delta_{\sigma})=o_{P}\left((\log p)^{-1}\right)$
by assumptions. Similar to the proof of Theorem 3.4, we obtain that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\mathcal{G}}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}^{stu}\|_{\mathcal{G}}\leq t)\right|=o(1).$
Lemma A.1.
Under Assumption 3, we have $\|\boldsymbol{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}\|_{1}=\|\boldsymbol{w}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}\|_{1}\leq C_{1}\eta$.
Proof of Lemma A.1.
By definition, we have
$\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})=\boldsymbol{0}_{p},$
which indicates that
$\sum_{k\in\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}}(\boldsymbol{w}^{(k)}-\boldsymbol{\beta})=\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}}(\boldsymbol{w}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}).$
It follows that $\|\boldsymbol{w}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}\|_{1}\leq\sum_{k\in\mathcal{A}_{\eta}}\alpha_{k}\|(\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}})^{-1}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}}\|_{1}\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{1}\leq C_{1}\eta$. ∎
Lemma A.2 (Lemma 2 of tian2023transfer).
Suppose that Assumption 1 holds for $k\in\{0\}\cup\mathcal{A}_{\eta}$, then there exists some positive constant $\kappa_{1},\kappa_{2},C_{3}$ and $C_{4}$ such that
$\frac{D(\boldsymbol{\Delta})}{\|\boldsymbol{\Delta}\|_{2}^{2}}\geq\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\frac{\|\boldsymbol{\Delta}\|_{1}}{\|\boldsymbol{\Delta}\|_{2}}\right),\quad\text{for all }\|\boldsymbol{\Delta}\|_{2}\leq 1$
with probability at least $1-C_{3}\exp(-C_{4}(n_{\mathcal{A}_{\eta}}+n_{0}))$.
Proof of Lemma A.2.
The proof simply follows from Lemma 2 of tian2023transfer. ∎
Lemma A.3 (Proposition 2 of negahban2009unified).
Suppose that Assumption 1 holds for $k\in\{0\}\cup\mathcal{A}_{\eta}$, then there exists some positive constant $\kappa_{1},\kappa_{2},C_{3}$ and $C_{4}$ such that
$\frac{D^{(0)}(\boldsymbol{\Delta})}{\|\boldsymbol{\Delta}\|_{2}^{2}}\geq\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\frac{\|\boldsymbol{\Delta}\|_{1}}{\|\boldsymbol{\Delta}\|_{2}}\right),\quad\text{for all }\|\boldsymbol{\Delta}\|_{2}\leq 1$
with probability at least $1-C_{3}\exp(-C_{4}n_{0})$, where $D^{(0)}(\boldsymbol{\Delta})=\boldsymbol{\Delta}^{\top}\left(\boldsymbol{U}_{0}^{\top}\boldsymbol{U}_{0}/n_{k}\right)\boldsymbol{\Delta}$.
Proof of Lemma A.2.
The proof simply follows from Proposition 2 of negahban2009unified. ∎
Lemma A.4.
Suppose that Assumptions 1-2 hold for $k\in\{0\}\cup\mathcal{A}_{\eta}$, if $\eta=O(1)$, we have
$\|\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})\|_{\infty}=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)(\sqrt{\log p}\vee\eta\log p\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2})}{n_{\mathcal{A}_{\eta}}+n_{0}}+\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right).$
Proof of Lemma A.4.
Note that
$-\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})$ $=\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\hat{\boldsymbol{U}}_{k}^{\top}(\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}^{(k)})+\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})$
$:$ $=E_{1}+E_{2}.$
For $E_{1}$, since
[…]
$=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)(\sqrt{\log p}\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2})}{n_{\mathcal{A}_{\eta}}+n_{0}}+\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right).$
For $E_{2}$, according to Lemma A.1 and the definition of $\mathcal{A}_{\eta}$, we have $\|\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}\leq\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{1}+\|\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}\leq(C_{1}+1)\eta$. This equation, together with Lemma A.6 implies that
$\left\|\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}(\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}-\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k})(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})\right\|_{\infty}$
$\leq$ $\frac{|\mathcal{A}_{\eta}|+1}{(n_{\mathcal{A}_{\eta}}+n_{0})}\|\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}-\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}\|_{\max}\|\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)\eta\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}\right)$
Moreover, by the definition of $\boldsymbol{w}^{\mathcal{A}_{\eta}}$, we have $\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})/(n_{\mathcal{A}_{\eta}}+n_{0})$ is a mean of the sample $\boldsymbol{u}_{i}^{(k)}\boldsymbol{u}_{i}^{(k)\top}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})$ with expectation $\boldsymbol{0}_{p}$ and
$\|\boldsymbol{u}_{i}^{(k)}\boldsymbol{u}_{i}^{(k)\top}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})\|_{\psi_{1}}\leq\|\boldsymbol{u}_{i}^{(k)}\|_{\psi_{2}}^{2}\|\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}\lesssim\eta.$
It follows from Bernstein’s inequality that
$\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})=O_{P}\left(\eta\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right).$
Hence, we have
$\|E_{2}\|_{\infty}=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)\eta\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}+\eta\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right),$
which indicates that
$\|\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})\|_{\infty}=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)(\sqrt{\log p}\vee\eta\log p\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2})}{n_{\mathcal{A}_{\eta}}+n_{0}}+\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right)$
if $\eta=O(1)$. ∎
Lemma A.5.
Suppose that Assumptions 1-2 hold for $k=0$, we have
$\|\nabla\hat{L}^{(0)}(\boldsymbol{\beta})\|_{\infty}=O_{P}\left(\frac{\sqrt{\log p}\vee\mathcal{V}_{n_{0},p}\|\varphi^{(0)}\|_{2}}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}\right).$
Proof of Lemma A.5.
Similar to the proof Lemma A.4, we have
[…]
$=$ $O_{P}\left(\frac{\sqrt{\log p}\vee\mathcal{V}_{n_{0},p}\|\varphi^{(0)}\|_{2}}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}\right).$
Lemma A.6 (Lemma C.3, Lemma C.4, and Lemma C.6 of fan2024latent).
Suppose that Assumptions 1-2 hold for $k$, and assume that $n_{k}=O(p)$, define
$\mathcal{V}_{n_{k},p}=\frac{n_{k}}{p}+\sqrt{\frac{\log p}{n_{k}}}+\sqrt{\frac{n_{k}\log p}{p}},$
then for any $\boldsymbol{\phi}\in\mathbb{R}^{r_{k}}$ with $\|\boldsymbol{\phi}\|_{2}=1$, we have
[…]
and
$\|\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}-\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}\|_{\max}=O_{P}\left(\frac{n_{k}}{p}+\log p\right).$
Lemma A.7.
Suppose that Assumptions 1-2 hold for $k\in\{1,\ldots,K\}$, for any $1\leq r\leq 3$ and any $\boldsymbol{w}_{1},\boldsymbol{w}_{2}\in\mathbb{R}^{p}$, we have
$|\hat{L}_{0}^{[r]}(\boldsymbol{w}_{1})-\hat{L}_{0}^{[r]}(\boldsymbol{w}_{2})-(L_{0}^{[r]}(\boldsymbol{w}_{1})-L_{0}^{[r]}(\boldsymbol{w}_{2}))|$
$\lesssim$ $\left(\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}\vee\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{1}\vee 1\right)\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}\sqrt{\frac{\log p}{n_{0}}}$
$+\sqrt{\frac{\log p}{n_{0}}}(\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}^{2}+\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{2}),$
where
$L_{0}^{[r]}(\boldsymbol{w})=\mathbb{E}\left[\frac{1}{n_{0}/3}\|\boldsymbol{e}^{(0)[r]\top}(\boldsymbol{Y}^{(0)}-\boldsymbol{U}_{0}\boldsymbol{w}-\boldsymbol{F}_{0}\boldsymbol{\gamma}^{(k)})\|_{2}^{2}\right]$
Proof of Lemma A.7.
Firstly, denote $\hat{\boldsymbol{U}}_{0}^{[r]}=\boldsymbol{e}^{(0)[r]\top}\hat{\boldsymbol{U}}_{0}$ and $\boldsymbol{U}_{0}^{[r]}=\boldsymbol{e}^{(0)[r]\top}\boldsymbol{U}_{0}$, we have
$\hat{L}_{0}^{[r]}(\boldsymbol{w}_{1})-\hat{L}_{0}^{[r]}(\boldsymbol{w}_{2})=\nabla\hat{L}_{0}^{[r]}(\boldsymbol{w}_{2})^{\top}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})+\frac{1}{2}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})^{\top}\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})$
$=$ $\nabla\hat{L}_{0}^{[r]}(\boldsymbol{\beta})^{\top}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})+(\boldsymbol{w}_{2}-\boldsymbol{\beta})^{\top}\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})+\frac{1}{2}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})^{\top}\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})$
Similarly, apply this equation to $L_{0}^{[r]}(\boldsymbol{w}_{1})-L_{0}^{[r]}(\boldsymbol{w}_{2})$, together with the equation above yields that
$\hat{L}_{0}^{[r]}(\boldsymbol{w}_{1})-\hat{L}_{0}^{[r]}(\boldsymbol{w}_{2})-(L_{0}^{[r]}(\boldsymbol{w}_{1})-L_{0}^{[r]}(\boldsymbol{w}_{2}))$
$=$ $(\nabla\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-\nabla L_{0}^{[r]}(\boldsymbol{\beta}))^{\top}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})+(\boldsymbol{w}_{2}-\boldsymbol{\beta})^{\top}\left(\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\right)(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})$
$+\frac{1}{2}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})^{\top}\left(\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\right)(\boldsymbol{w}_{1}-\boldsymbol{w}_{2}):=E_{1}+E_{2}+E_{3}$
By Lemma A.5 and the assumptions, we have $|E_{1}|\leq\|\nabla\hat{L}_{0}^{[r]}(\boldsymbol{\beta})\|_{\infty}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}\lesssim\sqrt{\log p/n_{0}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}$. Now we consider $(n_{0}/3)^{-1}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}$. By following a similar proof of the line of Lemma C.3 and Lemma C.6 in fan2024latent, we have
$\frac{1}{n_{0}/3}\|\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{U}_{0}^{[r]\top}\boldsymbol{U}_{0}^{[r]}\|_{\max}=O_{P}\left(\frac{\log p}{n_{0}}\right).$
Furthermore, since $\|\boldsymbol{a}^{\top}\boldsymbol{u}^{(0)}\boldsymbol{u}^{(0)\top}\boldsymbol{b}\|_{\psi_{1}}\leq\|\boldsymbol{u}^{(0)}\|_{\psi_{2}}^{2}\|a\|_{2}\|b\|_{2}$ for any vectors $\boldsymbol{a}$ and $\boldsymbol{b}$, by the Bernstein’s inequality, one attains that
$\left\|\boldsymbol{a}^{\top}\left(\frac{1}{n_{0}/3}\boldsymbol{U}_{0}^{[r]\top}\boldsymbol{U}_{0}^{[r]}-\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\right)\boldsymbol{b}\right\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{a}\|_{2}\|\boldsymbol{b}\|_{2}\right).$
Hence, we conclude that
$|E_{2}|+|E_{3}|\lesssim$ $\left\|\frac{1}{n_{0}/3}(\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{U}_{0}^{[r]\top}\boldsymbol{U}_{0}^{[r]})\right\|_{\max}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}(\frac{1}{2}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}+\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{1})$
$+\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}(\frac{1}{2}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}+\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{2})$
$\lesssim$ $\frac{\log p}{n_{0}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}(\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}\vee\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{1})$
$+\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}(\frac{1}{2}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}+\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{2})$
The proof is complete by leveraging the bounds for $|E_{1}|$ and $|E_{2}|+|E_{3}|$. ∎
Lemma A.8 ((chernozhukov2013gaussian, Corollary 2.1)).
Let $\boldsymbol{X}_{1},\cdots,\boldsymbol{X}_{n}=(X_{i1},\cdots,X_{id})^{\top}$ be n i.i.d. $d$-dimensional random vectors with mean $\mathbf{0}_{d}$. Suppose that there are some constants $0<C_{1}<C_{2}$ such that $\mathbb{E}[X_{ij}|^{2}\geq C_{1}$ and $X_{ij}$ is sub-exponential with $\|X_{ij}\|_{\psi_{1}}\leq C_{2}$ for all $1\leq j\leq d$. If $\log^{7}(dn)/n=o(1)$, then
$\lim_{n\to\infty}\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\boldsymbol{X}_{i}\right\|_{\infty}\leq t\right)-\mathbb{P}(\|\boldsymbol{N}\|_{\infty}\leq t)\right|=0,$
where $\boldsymbol{N}\sim\mathbb{N}(\mathbf{0}_{d},\mathbb{E}[\boldsymbol{X}_{i}\boldsymbol{X}_{i}^{\top}])$.
Lemma A.9 ((cai2025statistical, Lemma 7)).
For any $\delta>0$ and sequences of random vectors $\{X_{n}\}_{n\in\mathbb{N}},\{Y_{n}\}_{n\in\mathbb{N}},\{Z_{n}\}_{n\in\mathbb{N}}$ satisfying
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}(\|X_{n}\|_{\infty}\leq t)-\mathbb{P}(\|Z_{n}\|_{\infty}\leq t)\right|=\Delta_{n},$
we have that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}(\|X_{n}+Y_{n}\|_{\infty}\leq t)-\mathbb{P}(\|Z_{n}\|_{\infty}\leq t)\right|\leq\mathbb{P}(\|Y_{n}\|_{\infty}>\delta)+\sup_{t\in\mathbb{R}}\mathbb{P}(|\|Z_{n}\|_{\infty}-t|<\delta)+\Delta_{n}.$