### v3

# Transfer learning for high-dimensional Factor-augmented sparse linear model

Abstract
This paper introduces a transfer learning methodology for high-dimensional sparse linear models augmented with latent factors, addressing settings where covariates exhibit strong correlations and latent factor structures, which is motivated by the demand for stable and reliable estimation applications in economics and finance. Our framework jointly corrects for latent factor-induced effects and mitigates multicollinearity, reducing potential model misspecification relative to conventional linear sparse regression. When the target dataset is limited by sample size but multiple heterogeneous auxiliary datasets are accessible, we propose transfer-augmented estimators that leverage auxiliary information to improve estimation accuracy. we establish non-asymptotic $\ell_{1}$- and $\ell_{2}$-error bounds for the resulting estimators. To guard against negative transfer, we develop a data-driven source detection algorithm with provable consistency in selecting informative auxiliary datasets. In addition, we provide a procedure for constructing simultaneous confidence intervals for the regression coefficients of interest. Numerical studies confirm that the proposed approach achieves substantial gains in estimation accuracy and remain robust under heterogeneity across datasets. The work provides a theoretically rigorous and computationally tractable framework for incorporating heterogeneous auxiliary data in high-dimensional factor-structured regression problems.
Keywords: transfer learning, high-dimensional statistics, factor-augmented regression model, latent factor effect, negative transfer
1 Introduction
In many modern statistical and machine learning applications, data are often collected from multiple related sources that share similar yet non-identical underlying structures. Directly combining these heterogeneous datasets may lead to biased estimation due to distributional discrepancies, whereas analyzing the target dataset alone can suffer from a limited sample size. Transfer learning offers a principled framework to leverage auxiliary information from source datasets to improve estimation performance on the target task.
A growing body of work seeks to develop statistical foundations for transfer learning. Examples include nonparametric classification (Cai and Wei, 2021), transfer reinforcement learning (Chai et al., 2025), Gaussian graphical model (Li et al., 2023), and high-dimensional parametric regression models such as linear model (Li et al., 2022), generalized linear models (Bastani, 2021; Tian and Feng, 2023), smoothed quantile regression (Zhang and Zhu, 2025), multinomial regression (Yang et al., 2025b), and adaptive Huber regression models (Yang et al., 2025a). In these parametric regression models, the similarity between the source and target datasets is typically quantified via the distance between their regression coefficients.
However, those existing transfer learning methods for parametric regression mentioned above rely critically on weak dependence on the variables that ensure estimation consistency. In many applications–particularly in finance, economics, and biomedicine–the covariates exhibit strong cross-sectional dependence and are often driven by latent common factors. Under such structures, in the high-dimensional linear regression model, the conditions (e.g., the irrepresentable condition) required for variable selection and estimation consistency of Lasso-type procedures are violated (Fan et al., 2020). Moreover, when the latent factors contribute directly to the response, the linear regression model based only on the observed covariates becomes misspecified. As a result, transfer learning procedures built upon standard high-dimensional parametric regression may perform unreliably in the presence of latent factor effects or highly correlated covariates. This structural issue was considered in Fan et al. (2024), where they proposed the Factor Augmented (sparse linear) Regression Model (FARM), defined as follows:
$y$ $=\boldsymbol{f}^{\top}\boldsymbol{\gamma}^{*}+\boldsymbol{u}^{\top}\boldsymbol{\beta}^{*}+\mathcal{E},$ (1.1)
$\boldsymbol{x}$ $=\boldsymbol{B}\boldsymbol{f}+\boldsymbol{u},$
where $y\in\mathbb{R}$ is a scalar response variable, $\boldsymbol{x}\in\mathbb{R}^{p}$ is the covariate, $\boldsymbol{f}\in\mathbb{R}^{r}$ is the latent factors, $\boldsymbol{B}\in\mathbb{R}^{p\times r}$ is the factor loading matrix, and $\boldsymbol{u}\in\mathbb{R}^{p}$ is a idiosyncratic component. The parameter $\boldsymbol{\beta}^{*}\in\mathbb{R}^{p}$ is the true regression parameter vector that is of interest, and $\boldsymbol{\gamma}^{*}\in\mathbb{R}^{r}$ quantifies the contribution of the latent factor $\boldsymbol{f}$. The FARM model can be viewed as a generalization of the sparse linear model, as it can be rewritten as:
$y$ $=\boldsymbol{f}^{\top}\boldsymbol{\varphi}^{*}+\boldsymbol{x}^{\top}\boldsymbol{\beta}^{*}+\mathcal{E},$
$\boldsymbol{x}$ $=\boldsymbol{B}\boldsymbol{f}+\boldsymbol{u},$
where $\boldsymbol{\varphi}^{*}=\boldsymbol{\gamma}^{*}-\boldsymbol{B}^{\top}\boldsymbol{\beta}^{*}$ measures the extra contribution of $\boldsymbol{f}$ to $y$ beyond $\boldsymbol{x}$. It enhances sparse linear regression by incorporating informative directions spanned by the common factors $\boldsymbol{f}$. From another point of view, the FARM model can be viewed as a generalization of the factor regression model (Stock and Watson, 2002; Bing et al., 2021) as it utilizes additional information in $\boldsymbol{u}$. Therefore, the strong performance of FARM arises from its ability to jointly capture both the common factors $\boldsymbol{f}$ and idiosyncratic components $\boldsymbol{u}$ in explaining the response.
Although the FARM model effectively addresses the challenge of highly correlated covariates and latent factors by separating common factors and idiosyncratic components, in practice, the target dataset may have a limited sample size, and related auxiliary datasets often contain useful information that can enhance estimation accuracy. Since the regression coefficient $\boldsymbol{\beta}$ quantifies the contribution of the covariates $\boldsymbol{X}^{(0)}$ to the outcome $\boldsymbol{Y}^{(0)}$, this makes its estimation and inference a quantity of significant importance (Guo et al., 2022). To this end, we consider procedures for the FARM to improve estimation accuracy by extracting useful information from auxiliary datasets.
Motivated by these insights, we propose a transfer learning procedure for FARM (Trans-FARM) in this paper. The Trans-FARM allows information to be shared across multiple related datasets while accounting for the heterogeneity in both the factor structure and the idiosyncratic components. By removing the effects of strong correlations among covariates and the additional variations induced by latent factors, Trans-FARM yields more accurate estimation of regression coefficients. Compared with standard FARM, it improves estimation by borrowing similarity from informative sources; compared with transfer learning methods for linear models (Li et al., 2022; Tian and Feng, 2023), it provides more reliable estimation due to their misspecification of model. In addition to enhancing estimation, the more reliable conduction of simultaneous inference of $\boldsymbol{\beta}$ is also of key interest by utilizing auxiliary datasets. Moreover, it is often unclear which source datasets are informative in practice, and incorporating uninformative or mismatched sources may lead to negative transfer. So we also present a source detection method for our transfer learning framework. Our proposed framework enables statistical inference on regression parameters, which can be applied to assess the validity of transfer learning for factor regression in Lorenzi et al. (2016). Our main contributions are summarized as follows:
  • A Robust Transfer Learning Framework for Data with High-Correlated Covariates and Latent Factor Structures. We develop a transfer learning framework that explicitly accounts for the data with latent factors and highly correlated covariates, which is particularly relevant in applications from economics, finance, and biomedicine. The framework systematically leverages auxiliary datasets to enhance the estimation of regression coefficients, providing a solid foundation for both accurate estimation and valid inference in complex data environments.
  • Theoretical Guarantees and a Safeguard Against Negative Transfer. We establish the non-asymptotic $\ell_{1}$- and $\ell_{2}$-error bounds for our proposed Trans-FARM estimator, demonstrating its efficiency. Theoretically, it achieves a convergence rate that matches the common transfer learning methods (Li et al., 2022; Tian and Feng, 2023) when factors are absent, while strictly outperforming them when latent factors are present due to their misspecification of models. Furthermore, to ensure this superior performance in practice, we propose a source detection algorithm with proven consistency, which acts as a safeguard by identifying and utilizing only beneficial auxiliary datasets to avoid negative transfer.
  • Valid Inference for Heterogeneous Effects. Moving beyond point estimation, we construct a comprehensive inferential tool for the high-dimensional transfer learning setting by constructing simultaneous confidence intervals for regression coefficients. A key innovation is a studentized approach that yields coefficient-specific interval lengths, thereby accurately capturing the heterogeneous influence of different features on the response.
The rest of this paper is structured as follows. Section 2 introduces some preliminaries and the model formulation. Section 3 develops the theoretical results for our proposed methods, including the $\ell_{1}$ and $\ell_{2}$- estimation error bounds, and introduces the source detection algorithm together with its detection consistency. The procedure for simultaneous inference is presented in Section 4. Section 5 presents simulation results and a real data analysis, and Section 6 presents our conclusion. All proofs are provided in the Appendix.
Notation: For a vector $\mathbf{v}=(v_{1},\cdots,v_{p})^{\top}\in\mathbb{R}^{p}$, denote $\|\mathbf{v}\|=(\sum_{i=1}^{p}|v_{i}|^{2})^{1/2}$, $\|\mathbf{v}\|_{1}=\sum_{i=1}^{p}|v_{i}|$, $\|\mathbf{v}\|_{\infty}=\max_{1\leq i\leq p}|v_{i}|$. For $\mathcal{S}\subseteq[p]$, let $\mathcal{S}^{c}$ be the complement of $\mathcal{S}$, i.e., $\mathcal{S}^{c}=[p]\backslash\mathcal{S}$, where we use the standard notation $[p]$ to denote $\{1,\ldots,p\}$ here. For a matrix $\mathbf{A}=(a_{ij})_{1\leq i\leq k_{1},1\leq j\leq k_{2}}\in\mathbb{R}^{k_{1}\times k_{2}}$, denote $\|\mathbf{A}\|_{\max}=\max_{1\leq i\leq k_{1},1\leq j\leq k_{2}}|a_{ij}|$, $\|\mathbf{A}\|_{\infty}=\max_{1\leq i\leq k_{1}}\sum_{j=1}^{k_{2}}|a_{ij}|$, $\|\mathbf{A}\|_{1}=\max_{1\leq j\leq k_{2}}\sum_{i=1}^{k_{1}}|a_{ij}|$, and let $\|\mathbf{A}\|_{2}$ be the spectral norm of $\mathbf{A}$. We use the standard notation $\boldsymbol{A}_{i,j}$ to represent the $(i,j)$-th element of $\boldsymbol{A}$, and use $\boldsymbol{A}_{i,-j}$ to denote the $i$-th row without the $(j,j)$-th element of $\boldsymbol{A}$ for a given matrix $\boldsymbol{A}$. We define the sub-Gaussian norm of a sub-Gaussian random variable $X$ as $\|X\|_{\psi_{2}}=\inf\{t>0:\mathbb{E}\exp(X^{2}/t^{2})\leq 2\}$, and the sub-exponential norm of a sub-exponential random variable $Y$ as $\|Y\|_{\psi_{1}}=\inf\{t>0:\mathbb{E}\exp(|Y|/t)\leq 2\}$. For any two sequences $\{a_{n}\}$ and $\{b_{n}\}$, we use $a_{n}=O(b_{n}),a_{n}=o(b_{n}),$ and $a_{n}\asymp b_{n}$ to represent that there exists $c,C>0$ such that $|a_{n}|\leq C|b_{n}|$ for all $n$, $\lim_{n\to\infty}|a_{n}|/|b_{n}|=0$, and $0<c<|a_{n}/b_{n}|<C<\infty$, respectively. $a_{n}\lesssim b_{n}$ means $a_{n}=O(b_{n})$. For a sequence of random variables $\{X_{n}\}$, we use the notation $Z_{n}=O_{P}(1)$ if $\lim_{M\to\infty}\lim_{n\to\infty}\mathbb{P}(|Z_{n}|>M)=0$, and $Z_{n}=o_{P}(1)$ if $\lim_{n\to\infty}\mathbb{P}(|Z_{n}|>\varepsilon)=0$ for any $\varepsilon>0$. For any two sequences of random variables $\{X_{n}\}$ and $\{Y_{n}\}$, we say $X_{n}=O_{P}(Y_{n})$ if $X_{n}/Y_{n}=O_{P}(1)$, and $X_{n}=o_{P}(Y_{n})$ if $X_{n}/Y_{n}=o_{P}(1)$.
2 Preliminaries and Model Formulation
Suppose that we have a target dataset $\{\boldsymbol{x}_{i}^{(0)},y_{i}^{(0)}\}_{i=1}^{n_{0}}$ and $K$ independent source datasets $\left\{\{\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)}\}_{i=1}^{n_{k}}\right\}_{k=1}^{K}$. For $k\in\{0,1,\cdots,K\}$, we consider the following model (in matrix form):
[…]
$\boldsymbol{X}^{(k)}$ $=\boldsymbol{F}_{k}{\boldsymbol{B}_{k}}^{\top}+\boldsymbol{U}_{k},\quad k=0,\cdots,K,$
[…]
The matrix form in (2.1) can also be written as
[…]
$\boldsymbol{X}^{(k)}$ $=\boldsymbol{F}_{k}\boldsymbol{B}_{k}^{\top}+\boldsymbol{U}_{k},\quad k=0,\cdots,K,$
where $\boldsymbol{\varphi}^{(k)}=\boldsymbol{\gamma}^{(k)}-\boldsymbol{B}_{k}^{\top}\boldsymbol{w}^{(k)}$ quantifies the additional effect of factors $\boldsymbol{F}_{k}$ to $\boldsymbol{Y}^{(k)}$.
Since the latent factor $\boldsymbol{f}_{i}^{(k)}$, the idiosyncratic component $\boldsymbol{u}_{i}^{(k)}$, and the loading matrix $\boldsymbol{B}_{k}$ are all needed to be estimated, we first provide their estimators. Note that $\boldsymbol{f}_{i}$ is not identifiable because $\boldsymbol{x}_{i}^{(k)}=\boldsymbol{B}_{k}\boldsymbol{f}_{i}^{(k)}+\boldsymbol{u}_{i}^{(k)}=\boldsymbol{B}_{k}\boldsymbol{T}(\boldsymbol{T}^{-1}\boldsymbol{f}_{i}^{(k)})+\boldsymbol{u}_{i}^{(k)}$ for any non-singular matrix $\boldsymbol{T}\in\mathbb{R}^{r_{k}\times r_{k}}$. To ensure identifiability, we impose the following normalization:
$\mbox{cov}(\boldsymbol{f}^{(k)})=\boldsymbol{I}_{r_{k}}\text{ and }\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k}\text{ is diagonal.}$
We also assume that $\mathcal{E}^{(k)}$, $\boldsymbol{f}^{(k)}$ and $\boldsymbol{u}^{(k)}$ are independent with each other, and $\mathbb{E}[\mathcal{E}^{(k)}]=0$.
Then the estimates simply follows by solving the constrained least squares problem for $k=0,\ldots,K$:
$(\hat{\boldsymbol{F}}_{k},\hat{\boldsymbol{B}}_{k})=\mathop{\arg\min}_{\boldsymbol{F}\in\mathbb{R}^{n\times r_{k}},\boldsymbol{B}\in\mathbb{R}^{p\times r_{k}}}\|\boldsymbol{X}^{(k)}-\boldsymbol{F}\boldsymbol{B}^{\top}\|_{\mathbb{F}}^{2}$ (2.3)
$\text{subject to }\frac{1}{n_{k}}\boldsymbol{F}^{\top}\boldsymbol{F}=\boldsymbol{I}_{r_{k}}\text{, and }\boldsymbol{B}^{\top}\boldsymbol{B}\text{ is diagonal.}$
Simple calculation yields that $\hat{\boldsymbol{F}}_{k}/\sqrt{n_{k}}$ is the eigenvectors corresponding to the top $r_{k}$ eigenvectors of $\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top}$, $\hat{\boldsymbol{B}}_{k}=(\hat{\boldsymbol{F}}_{k}^{\top}\hat{\boldsymbol{F}}_{k})^{-1}\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{X}^{(k)}=\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{X}^{(k)}/n_{k}$, and the estimator for the remainder term is $\hat{\boldsymbol{U}}_{k}=\boldsymbol{X}^{(k)}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{B}}_{k}^{\top}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}/n_{k})\boldsymbol{X}^{(k)}$. Then we have $\boldsymbol{X}^{(k)}=\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{B}}_{k}^{\top}+\hat{\boldsymbol{U}}_{k}$, and $\hat{\boldsymbol{F}}_{k}$, $\hat{\boldsymbol{B}}_{k}$, and $\hat{\boldsymbol{U}}_{k}$ are estimations for $\boldsymbol{F}_{k}$, $\boldsymbol{B}_{k}$, and $\boldsymbol{U}_{k}$ in model (2.2).
Remark 2.1.
If the intercept is considered, i.e., the first column of $\boldsymbol{X}^{(k)}$ is all $1$, let $\boldsymbol{X}_{-1}^{(k)}=(\boldsymbol{x}_{1,-1}^{(k)},\cdots,\boldsymbol{x}_{n,-1}^{(k)})^{\top}\in\mathbb{R}^{n_{k}\times(p-1)}$ denote the submatrix of $\boldsymbol{X}^{(k)}$ comprising its second through $p$-th columns. By removing the intercept term, similar calculation yields that $\hat{\boldsymbol{F}}_{k}/\sqrt{n_{k}}$ is the eigenvectors corresponding to the top $r_{k}$ eigenvectors of $\boldsymbol{X}_{-1}^{(k)}{\boldsymbol{X}_{-1}^{(k)}}^{\top}$, $\hat{\boldsymbol{B}}_{k,-1}=\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{X}_{-1}^{(k)}/n_{k}$, and $\hat{\boldsymbol{U}}_{k,-1}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}/n_{k})\boldsymbol{X}_{-1}^{(k)}$. Then we can write $\hat{\boldsymbol{B}}_{k}=(\boldsymbol{0}_{r_{k}},\hat{\boldsymbol{B}}_{k,-1}^{\top})^{\top}\in\mathbb{R}^{p\times r_{k}}$, $\hat{\boldsymbol{U}}_{k}=(\boldsymbol{1}_{n_{k}},\hat{\boldsymbol{U}}_{k,-1})\in\mathbb{R}^{n_{k}\times p}$ such that $\boldsymbol{X}^{(k)}=\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{B}}_{k}^{\top}+\hat{\boldsymbol{U}}_{k}$, and $\hat{\boldsymbol{F}}_{k}$, $\hat{\boldsymbol{B}}_{k}$, and $\hat{\boldsymbol{U}}_{k}$ are estimations for $\boldsymbol{F}_{k}$, $\boldsymbol{B}_{k}$, and $\boldsymbol{U}_{k}$, respectively.
To attain the consistency of these estimations, we first impose some regularity assumptions that are standard in the high-dimensional factor model (Bai, 2003; Fan et al., 2013, 2024).
Assumption 1 (Assumptions on $\boldsymbol{f}^{(k)}$, $\boldsymbol{u}^{(k)}$, and $\boldsymbol{B}_{k}$ (for the $k$-th dataset)).
  1. (a)
    Assume that $\boldsymbol{\nu}^{(k)}=(\boldsymbol{u}^{(k)\top},\boldsymbol{f}^{(k)\top})^{\top}$ is sub-Gaussian with $\|\boldsymbol{\nu}^{(k)}\|_{\psi_{2}}\leq v_{0}$.
  2. (b)
    There exists a constant $C>1$ such that $p/C\leq\lambda_{\min}(\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k})\leq\lambda_{\max}(\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k})\leq pC$. Moreover, we assume that $n_{k}=O(p)$.
  3. (c)
    There exists a constant $\Upsilon>0$ such that $\|\boldsymbol{B}_{k}\|_{\max}\leq\Upsilon$ and $\mathbb{E}|\boldsymbol{u}^{(k)\top}\boldsymbol{u}^{(k)}-\mbox{tr}(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)})|^{4}\leq\Upsilon p^{2}$.
  4. (d)
    There exists a positive constant $\iota<1$ such that $\iota\leq\lambda_{\min}(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)})$, $\|\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\|_{1}\leq 1/\iota$ and $\min_{1\leq i,\ell\leq p}\mbox{Var}(u_{i}^{(k)}u_{\ell}^{(k)})\geq\iota$.
The consistency of factor estimations is summarized in the following lemma, which follows from Lemmas D.2 and D.3 in Wang and Fan (2017).
Lemma 2.1.
Define $\boldsymbol{H}_{k}=n_{k}^{-1}\boldsymbol{V}_{k}^{-1}\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{F}_{k}\boldsymbol{B}_{k}^{\top}\boldsymbol{B}_{k}$, where $\boldsymbol{V}_{k}\in\mathbb{R}^{r_{k}\times r_{k}}$ is a diagonal matrix with its diagonal elements being the first $r_{k}$ largest eigenvalues of $n_{k}^{-1}\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)\top}}$. If Assumption 1 holds for any $k\in\{0,\ldots,K\}$, and $\log n_{k}=o(p)$, for $k\in\{0,\ldots,K\}$, we then have for the $k$-th dataset,
  • $\|\hat{\boldsymbol{F}}_{k}-\boldsymbol{F}_{k}\boldsymbol{H}_{k}^{\top}\|_{\mathbb{F}}^{2}=O_{P}(n_{k}/p+1/n_{k})$, $\|\hat{\boldsymbol{F}}_{k}-\boldsymbol{F}_{k}\boldsymbol{H}_{k}^{\top}\|_{\max}^{2}=O_{P}((1/p+1/n_{k}^{2})\log^{2}n_{k})$.
  • For any $\mathcal{J}\subseteq\{1,\cdots,p\}$, $\max_{j\in\mathcal{J}}\sum_{i=1}^{n_{k}}|\hat{u}_{ij}^{(k)}-u_{ij}^{(k)}|^{2}=O_{P}(\log|\mathcal{J}|+n_{k}/p).$
  • $\|\boldsymbol{H}_{k}^{\top}\boldsymbol{H}_{k}-\boldsymbol{I}_{r_{k}}\|=O_{P}(1/n_{k}+1/p)$.
  • $\max_{j\in[p]}\|\hat{\boldsymbol{b}}_{j}^{(k)}-\boldsymbol{H}_{k}\boldsymbol{b}_{j}^{(k)}\|_{2}^{2}=O_{P}(\log p/n_{k})$, $\|\hat{\boldsymbol{B}}_{k}-\boldsymbol{B}_{k}\boldsymbol{H}_{k}^{-1}\|_{\max}^{2}=O_{P}(\log p/n_{k})$.
Remark 2.2.
The true number of latent factors $r_{k}$ in each dataset is unknown in practice, and determining $r_{k}$ in a data-driven manner is a crucial task. A rich literature has proposed various approaches for estimating $r_{k}$ (Bai and Ng, 2002; Lam and Yao, 2012; Ahn and Horenstein, 2013; Fan et al., 2022). Our theoretical results remain valid as long as $r_{k}$ is replaced by any consistent estimator $\hat{r}_{k}$, i.e., we only require
$\mathbb{P}(\hat{r}_{k}=r_{k})\to 1,\quad\text{as }\quad n_{k}\to\infty.$
Therefore, we assume throughout this paper that the number of factors $r_{k}$ is known. For empirical implementation, we adopt the eigenvalue ratio method (Lam and Yao, 2012; Ahn and Horenstein, 2013) to determine $r_{k}$. Specifically, let $\lambda_{i}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)\top}})$ denote the eigenvalues of the Gram matrix $\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)\top}}$. Then the estimator of the number of factors is given by
$\hat{r}_{k}=\mathop{\arg\max}_{i\leq\mathcal{K}}\frac{\lambda_{i}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top})}{\lambda_{i+1}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top})},$
where $1\leq\mathcal{K}\leq n_{k}$ is a prescribed upper bound for $r_{k}$.
Now we consider the way to attain an estimator for the regression coefficient vector $\boldsymbol{w}^{(k)}$ in each dataset. It is commonly presumed that only a small subset of covariates exerts influence on the response variable under the high-dimensional setting. This corresponds to assuming that the true coefficient vector $\boldsymbol{w}^{(k)}$ is sparse. After attaining the estimators for $\hat{\boldsymbol{F}}_{k}$, $\hat{\boldsymbol{B}}_{k}$, and $\hat{\boldsymbol{U}}_{k}$ via (2.3), the regularized estimators of the unknown parameter vectors $\boldsymbol{w}^{(k)}$ and $\boldsymbol{\gamma}^{(k)}$ for $k=0,\cdots,K$ in FARM model are
$(\hat{\boldsymbol{w}}^{(k)},\hat{\boldsymbol{\gamma}}^{(k)})=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p},\boldsymbol{\gamma}\in\mathbb{R}^{r_{k}}}\left\{\frac{1}{2n_{k}}\|\boldsymbol{Y}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}-\hat{\boldsymbol{F}}_{k}\boldsymbol{\gamma}\|_{2}^{2}+\lambda\|\boldsymbol{w}\|_{1}\right\},$ (2.4)
where $\lambda>0$ is a tuning parameter.
3 Factor Augmented Transfer Learning
3.1 Oracle-Trans-FARM Estimation
Recall the estimators in (2.4). Let $\hat{\boldsymbol{P}}_{k}=n_{k}^{-1}\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}$ be the projection matrix. Denote $\tilde{\boldsymbol{Y}}^{(k)}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{P}}_{k})\boldsymbol{Y}^{(k)}$. Recall that $\hat{\boldsymbol{U}}_{k}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}/n_{k})\boldsymbol{X}^{(k)}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{P}}_{k})\boldsymbol{X}^{(k)}$, which implies that $\hat{\boldsymbol{F}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}=\boldsymbol{0}_{r_{k}\times p}$ and (2.4) can be equivalently written as
$\hat{\boldsymbol{w}}^{(k)}$ $=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\left\{\frac{1}{2n_{k}}\|\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}\|_{2}^{2}+\lambda\|\boldsymbol{w}\|_{1}\right\}$ (3.1)
$=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\left\{\frac{1}{2n_{k}}\|(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{P}}_{k})(\boldsymbol{Y}^{(k)}-\boldsymbol{X}^{(k)}\boldsymbol{w})\|_{2}^{2}+\lambda\|\boldsymbol{w}\|_{1}\right\}$
$\hat{\boldsymbol{\gamma}}^{(k)}$ $=(\hat{\boldsymbol{F}}_{k}^{\top}\hat{\boldsymbol{F}}_{k})^{-1}\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{Y}^{(k)}=\frac{1}{n_{k}}\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{Y}^{(k)}$
Inspired by the insights of transfer learning, we denote $\boldsymbol{\beta}=\boldsymbol{w}^{(0)}$ as the target regression coefficient vector and suppose that $\boldsymbol{\beta}$ is $s$-sparse, i.e., $\|\boldsymbol{\beta}\|_{0}\leq s$ with $s\ll p$. Define the $k$-th contrast vector $\boldsymbol{\delta}^{(k)}=\boldsymbol{\beta}-\boldsymbol{w}^{(k)}$ to quantify the discrepancy between the target and the $k$-th source datasets. We use the $\ell_{1}$-distance $\|\boldsymbol{\delta}^{(k)}\|_{1}$ to measure the similarity of the target dataset and the $k$-th source dataset. We begin with considering a general transfer learning algorithm on FARM when the transferable set $\mathcal{A}_{\eta}$ is known, where $\mathcal{A}_{\eta}=\{1\leq k\leq K:\|\boldsymbol{\delta}^{(k)}\|_{1}\leq\eta\}$ for some $\eta$. Transfer learning under this setting is expected to enhance estimation accuracy of $\boldsymbol{\beta}$ when $\eta$ is sufficiently small, as the target and source parameters are closely aligned. According to (3.1), we are motivated to attain the initial estimation of $\boldsymbol{\beta}$ through the following regularized loss function, denoted by $\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}$:
$\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\left\{\frac{1}{2(n_{\mathcal{A}_{\eta}}+n_{0})}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}\|_{2}^{2}+\lambda_{\boldsymbol{w}}\|\boldsymbol{w}\|_{1}\right\},$ (3.2)
where $n_{\mathcal{A}_{\eta}}=\sum_{k\in\mathcal{A}_{\eta}}n_{k}$, and $\lambda_{\boldsymbol{w}}$ is a tuning parameter. The equation above can be understood as finding a solution that converges to its population version $\boldsymbol{w}^{\mathcal{A}_{\eta}}$ under certain regularization conditions, where $\boldsymbol{w}^{\mathcal{A}_{\eta}}$ satisfies
$\boldsymbol{w}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\mathbb{E}\left[\frac{1}{2(n_{\mathcal{A}_{\eta}}+n_{0})}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\boldsymbol{Y}^{(k)}-\boldsymbol{U}_{k}\boldsymbol{w}-\boldsymbol{F}_{k}\boldsymbol{\gamma}^{(k)}\|_{2}^{2}\right].$ (3.3)
[…]
$\mathbb{E}\left[\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})/n_{k}\right]=\mathbb{E}\left[\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{u}^{(k)}\boldsymbol{u}^{(k)\top}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})\right]=\boldsymbol{0}_{p},$
where $\alpha_{k}=n_{k}/(n_{\mathcal{A}_{\eta}}+n_{0})$. Hence, $\boldsymbol{w}^{\mathcal{A}_{\eta}}$ has a explicit form: $\boldsymbol{w}^{\mathcal{A}_{\eta}}=(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}})^{-1}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\\ \boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\boldsymbol{w}^{(k)}$, where $\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}}=\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}$, and $\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}=\mathbb{E}[\boldsymbol{u}^{(k)}\boldsymbol{u}^{(k)\top}]$.
After attaining $\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}$, we define $\boldsymbol{\delta}^{\mathcal{A}_{\eta}}=\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}}$, and obtain its estimator $\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}$ via solving:
$\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{\delta}\in\mathbb{R}^{p}}\left\{\frac{1}{2n_{0}}\|\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}(\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\boldsymbol{\delta})\|_{2}^{2}+\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\delta}\|_{1}\right\},$ (3.4)
where $\lambda_{\boldsymbol{\delta}}$ is a tuning parameter. We regard $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}$ as the estimator of $\boldsymbol{\beta}$. The algorithm is summarized in Algorithm 1.
Input: Target data $(\boldsymbol{X}^{(0)},\boldsymbol{Y}^{(0)})$, source data $\{(\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)})\}_{k=1}^{K}$, penalty parameters $\lambda_{\boldsymbol{w}}$ and $\lambda_{\boldsymbol{\delta}}$, transferring set $\mathcal{A}_{\eta}$, factor bound $\mathcal{K}$.
Output: Estimated coefficient vector $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}$.
1 Preprocessing step: Estimate $\hat{r}_{k}=\arg\max_{i\leq\mathcal{K}}\dfrac{\lambda_{i}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top})}{\lambda_{i+1}(\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top})}$. Let $\hat{\boldsymbol{F}}_{k}/\sqrt{n_{k}}$ be the eigenvectors corresponding to the top $\hat{r}_{k}$ eigenvalues of $\boldsymbol{X}^{(k)}{\boldsymbol{X}^{(k)}}^{\top}$. Compute $\hat{\boldsymbol{B}}_{k}=\hat{\boldsymbol{F}}_{k}^{\top}\boldsymbol{X}^{(k)}/n_{k}$ and $\hat{\boldsymbol{U}}_{k}=(\boldsymbol{I}_{n_{k}}-\hat{\boldsymbol{F}}_{k}\hat{\boldsymbol{F}}_{k}^{\top}/n_{k})\boldsymbol{X}^{(k)}$.
2 Transferring step: Compute (3.2), i.e.,
$\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{w}\in\mathbb{R}^{p}}\left\{\frac{1}{2(n_{\mathcal{A}_{\eta}}+n_{0})}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}\|_{2}^{2}+\lambda_{\boldsymbol{w}}\|\boldsymbol{w}\|_{1}\right\}$
3 Debiasing step: Compute (3.4), i.e.,
$\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}=\mathop{\arg\min}\limits_{\boldsymbol{\delta}\in\mathbb{R}^{p}}\left\{\frac{1}{2n_{0}}\|\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}(\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\boldsymbol{\delta})\|_{2}^{2}+\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\delta}\|_{1}\right\}$
4 Let $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}$.
Algorithm 1 Oracle-Trans-FARM Algorithm
Before presenting the error bound of $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}$, several technical assumptions below are required.
Assumption 2 (Assumption on $\mathcal{E}^{(k)}$).
There exists a positive constant $c<\infty$ such that $\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\mathcal{E}^{(k)}\|_{\psi_{2}}\leq c$.
Assumption 3.
There exists a positive constant $C_{1}\!\!<\!\!\infty$ such that $\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\!\|(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}})^{-1}\!\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\!\|_{1}\\ \leq\!C_{1}$, where $\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}}=\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}$, and $\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}=\mathbb{E}[\boldsymbol{u}^{(k)}\boldsymbol{u}^{(k)\top}]$.
Assumption 2 is standard and mild in high-dimensional regression models (Fan et al., 2024; Zhang and Cheng, 2017; Tian and Feng, 2023). Assumption 3 is also commonly adopted in transfer learning settings (Tian and Feng, 2023; Zhang and Zhu, 2025). In particular, Assumption 3 is reasonable as it does not require exact sparsity of $(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{\mathcal{A}_{\eta}})^{-1}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}$ and is satisfied, for example, by Toeplitz-type covariance structures. This assumption can be further interpreted as a constraint on the heterogeneity between the target idiosyncratic component and the source idiosyncratic component; see Condition 4 in Li et al. (2023) for a detailed discussion. However, since the covariates are not sub-Gaussian if factors exist, our assumptions can be regarded as weaker conditions. This assumption can also be relaxed at the cost of slightly weaker estimation bounds, following arguments similar to those in Theorem 3 of Tian and Feng (2023).
Now, we are ready to present our main results for the estimation error bounds of the Oracle Trans-FARM algorithm proposed in Algorithm 1. Given the transferable set $\mathcal{A}_{\eta}\subseteq\{1,\ldots,K\}$, we consider the parameter space
[…]
Theorem 3.1.
Suppose that Assumptions 1-2 hold for $k\in\{0\}\cup\mathcal{A}_{\eta}$ and Assumption 3 holds. Assume that $\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}n_{k}=O(p)$, $\eta=O(1)$, the following (3.5) holds for a constant $c$ small enough:
$\max\left\{\sqrt{\frac{s\log p}{n_{0}}},\frac{s(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}},\frac{s}{p}\right\}<c,$ (3.5)
and
$\max\left\{\frac{(|\mathcal{A}_{\eta}|+1)(1\vee\eta\log^{\frac{1}{2}}p\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2}\log^{-\frac{1}{2}}p)}{\sqrt{n_{\mathcal{A}_{\eta}}+n_{0}}},\frac{\mathcal{V}_{n_{0},p}\|\varphi^{(0)}\|_{2}}{\sqrt{n_{0}\log p}}\right\}=O(1),$ (3.6)
where
$\mathcal{V}_{n_{k},p}=\frac{n_{k}}{p}+\sqrt{\frac{\log p}{n_{k}}}+\sqrt{\frac{n_{k}\log p}{p}}.$
[…]
$\|\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}\|_{2}$ $\lesssim\sqrt{\frac{\log p}{n_{0}}}\eta+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}$
$\|\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}\|_{1}$ $\lesssim s\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\eta+\sqrt{s\eta}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}$
Remark 3.1.
We provide some illustrations of the conditions in Theorem 3.1. The dimension-type condition $n_{k}=O(p)$ is widely applied in factor analysis to ensure consistency. Conditions (3.5)-(3.6) are also mild. Consider a regime that $\{n_{k}\}_{k\in\{0\}\cup\mathcal{A}_{\eta}}$ are the same order, the additional contribution of factors is bounded by $\|\varphi^{(k)}\|_{2}=O_{P}(\sqrt{\log p})$, and $\eta=O(1)$, (3.6) and the second term in (3.5) changes into $\sqrt{(|\mathcal{A}_{\eta}|+1)\log p/n_{0}}=O(1)$ and $s\log p/n_{0}<c$ for some small constant $c$.
The proof of Theorem 3.1 can be found in Appendix A.1. From Theorem 3.1, the $\ell_{1}$- and $\ell_{2}$-error bounds of our estimator improve upon those in Fan et al. (2024) provided that $\eta\ll s\sqrt{\log p/n_{0}}$ and $n_{\mathcal{A}_{\eta}}+n_{0}\gg n_{0}$. This demonstrates the advantage gained through transfer learning.
3.2 Source Detection and Its Consistency
In this section, we consider how to detect the transferable set $\mathcal{A}_{\eta}$ through a data-driven manner. Since the performance of transfer learning deeply depends on the transferable set, the negative transfer (Pan and Yang, 2009; Ge et al., 2014) would happen if irrelevant sources are included in the transferable set. Therefore, a reliable procedure for identifying which source datasets can be effectively leveraged is crucial.
Motivated by Tian and Feng (2023); Yang et al. (2025a, b), we develop a source detection procedure as follows. We begin by randomly partitioning the target dataset into three folds, denoted by $\{(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]})\}_{r=1}^{3}$. Note that we choose three folds and suppose $n_{0}$ is divisible by $3$ only for convenience. Next, for each source dataset, we apply the transferring step using every two folds of target data and evaluate the resulting model on the remaining fold. We compute the corresponding value of the given loss function on the remaining target fold and average across the three splits to obtain the cross-validated loss $\hat{L}_{0}^{(k)}$ for each source $k$. As a benchmark, we also fit the Lasso estimator using every pair of target folds and evaluate its performance on the remaining fold, yielding the average cross-validated loss $\hat{L}_{0}^{(0)}$, which serves as the loss of the target-only model. Finally, we compare the difference $\hat{L}_{0}^{(k)}-\hat{L}_{0}^{(0)}$ with a pre-specified threshold. We then include source $k$ in $\widehat{\mathcal{A}}$ if $\hat{L}_{0}^{(k)}-\hat{L}_{0}^{(0)}$ is below this threshold. Write the $r$-th fold of target data is $(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]})$ with the data included in it denoting by $\mathcal{R}^{[r]}\subseteq[n_{0}]$. Let $\boldsymbol{e}^{(0)[r]}$ be the $n_{0}$-dimensional vector with $\boldsymbol{e}_{\mathcal{R}^{[r]}}^{(0)[r]}=1$ and $\boldsymbol{e}_{[n_{0}]\backslash\mathcal{R}^{[r]}}^{(0)[r]}=0$. For any regression coefficient estimate $\boldsymbol{w}$, the loss function under our FARM setting on the $r$-th fold of target data is
$\hat{L}_{0}^{[r]}(\boldsymbol{w})=\frac{1}{n_{0}/3}\|{\boldsymbol{e}^{(0)[r]}}^{\top}(\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}\boldsymbol{w})\|_{2}^{2},$ (3.7)
After obtaining the estimate $\widehat{\mathcal{A}}$, we apply Algorithm 1 with $\widehat{\mathcal{A}}$ as input. The full procedure is described in Algorithm 2.
Input: target data $(\boldsymbol{X}^{(0)},\boldsymbol{Y}^{(0)})$, source data $\{(\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)})\}_{k=1}^{K}$, a constant $\epsilon_{0}>0$, penalty parameters $\{\{\lambda^{(k)[r]}\}_{k=0}^{K}\}_{r=1}^{3}$, constants $\epsilon_{0}$ and $\hat{\sigma}$.
Output: the estimated coefficient vector $\hat{\boldsymbol{\beta}}$, and the determined transferring set $\widehat{\mathcal{A}}$
1 Preprocessing step: Run step 1 (Preprocessing step) of Algorithm 1.
2 Transferable source detection: Randomly divide $(\boldsymbol{X}^{(0)},\boldsymbol{Y}^{(0)})$ into three sets of equal size as $\{(\boldsymbol{X}^{(0)[i]},\boldsymbol{Y}^{(0)[i]})\}_{i=1}^{3}$
3 for $r=1$ to 33 do
4 $\hat{\boldsymbol{\beta}}^{[r]}\leftarrow$ fit the Lasso on $\{(\boldsymbol{X}^{(0)[i]},\boldsymbol{Y}^{(0)[i]})\}_{i=1}^{3}\backslash(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]})$ with penalty parameter $\lambda^{(0)[r]}$
5 $\hat{\boldsymbol{w}}^{(k)[r]}\leftarrow$ run step 2 (transferring step) in Algorithm 1 with $(\{(\boldsymbol{X}^{(0)[i]},\boldsymbol{Y}^{(0)[i]})\}_{i=1}^{3}\backslash(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]}))\cup(\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)})$ and penalty parameter $\lambda^{(k)[r]}$ for all $k\neq 0$
6   Compute the loss $\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})$ for $k\neq 0$ and $\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})$ on $(\boldsymbol{X}^{(0)[r]},\boldsymbol{Y}^{(0)[r]})$.
7
8 end for
[…]
10 $\widehat{\mathcal{A}}\leftarrow\{k\neq 0:\hat{L}_{0}^{(k)}\leq\hat{L}_{0}^{(0)}+\epsilon_{0}\hat{\sigma}^{2}\}$
11 $\widehat{\mathcal{A}}-$Trans-GLM: $\hat{\boldsymbol{\beta}}\leftarrow$ run Algorithm 1 using $\{(\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)})\}_{k\in\{0\}\cup\widehat{\mathcal{A}}}$
12 Output $\hat{\boldsymbol{\beta}}$
Algorithm 2 Trans-FARM Algorithm
To ensure the consistency of the source detection algorithm, we impose some assumptions below.
Assumption 4.
  • (a) There exists a positive constant $C_{2}<\infty$ such that for every $k$,
    $\max_{i\in\{0,k\}}\left\|\left(\frac{n_{0}}{n_{0}+n_{k}}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}+\frac{n_{k}}{n_{0}+n_{k}}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(k)}\right)^{-1}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(i)}\right\|_{1}\leq C_{2}.$
    Moreover, assume that
    $\frac{\mathcal{V}_{n_{k},p}\|\boldsymbol{\varphi}^{(k)}\|_{2}}{\sqrt{(n_{k}+n_{0})\log p}}=O(1).$
  • (b) There exists a positive constant $\tilde{\eta}$ such that $\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{1}\leq\tilde{\eta}$ for $k\in\mathcal{A}_{\eta}^{c}$.
  • (c) Assume that for some $\eta$,
    $\epsilon_{0}$ $\gtrsim\frac{s\log p}{n_{0}}+\eta\sqrt{\frac{\log p}{n_{0}}}+\frac{\eta^{2}}{p}+s^{2}\left(\frac{\log p}{n_{0}}\right)^{\frac{3}{2}},\quad\text{and}$
    $\inf_{k\in\mathcal{A}_{\eta}^{c}}\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{2}^{2}$ $\gtrsim\epsilon_{0}\vee\left(\frac{s\log p}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}(\tilde{\eta}\vee\tilde{\eta}^{2})+\tilde{\eta}^{\frac{3}{2}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}+\tilde{\eta}\sqrt{\frac{s\log p}{n_{0}+n_{k}}}\right),$
    where $\epsilon_{0}$ is defined in Algorithm 2.
  • (d) Suppose that $\sigma_{0}^{2}=\mbox{Var}(\mathcal{E}^{(0)})$ with $0<\sigma_{0}^{2}<\infty$. There exists a $\Delta_{\sigma}(0<\Delta_{\sigma}\leq 1)$ such that $|\hat{\sigma}/\sigma_{0}-1|=O_{P}(\Delta_{\sigma})$, and $\Delta_{\sigma}=O(1)$, where $\hat{\sigma}$ is defined in Algorithm 2.
Remark 3.2.
Assumption 4 (a)(b)(c) ensures that the transferring step can be applied to attain some estimator for the consistency of source detection. Similar conditions can be found in Tian and Feng (2023); Yang et al. (2025a, b); Zhang and Zhu (2025). Assumption 4 (d) is to ensure that $\hat{\sigma}^{2}$ used in Algorithm 2 is the order of a constant such that our source detection algorithm can be proved to be consistent. This type of choice of $\hat{\sigma}$ is similar to Tian and Feng (2023); Yang et al. (2025a, b). As stated in Yang et al. (2025b), the source detection is not sensitive to the choice of $\epsilon_{0}\hat{\sigma}^{2}$ in practice.
Under Assumption 4, Theorem 3.2 establishes the consistency of our proposed source detection algorithm. The proof is provided in Appendix A.2.
Theorem 3.2.
Suppose that Assumptions 1-2 hold for $k\in[K]$ and Assumptions 3-4 also hold. In addition, assume that $\max_{k\in[K]}n_{k}=O(p)$, $\eta=O(1)$, and conditions (3.5)-(3.6) are satisfied. If we choose $\lambda^{(0)[r]}\asymp\sqrt{\log p/n_{0}}$ and $\lambda^{(k)[r]}\asymp\sqrt{\log p/(n_{0}+n_{k})}$, then $\mathbb{P}(\widehat{\mathcal{A}}=\mathcal{A}_{\eta})\to 1$.
4 Simultaneous Inference
In this section, we study the construction of confidence intervals for the regression coefficients $\boldsymbol{\beta}$ in a more stable and reliable way by leveraging auxiliary datasets, as it is important to quantify the uncertainty. Both entrywise confidence intervals (van de Geer et al., 2014; Javanmard and Montanari, 2014; Tian and Feng, 2023) and groupwise simultaneous confidence intervals (Zhang and Cheng, 2017; Fan et al., 2024; Yan et al., 2023; Cai et al., 2025) can be obtained. We focus on the construction of groupwise simultaneous confidence intervals, as the entrywise intervals are simpler and arise as a direct byproduct of the groupwise procedure. Let $\mathcal{G}\subseteq[p]$ be the indices of parameters of interest, and denote $\|\boldsymbol{Q}\|_{\mathcal{G}}=\|\boldsymbol{Q}_{\mathcal{G}}\|_{\infty}$ as the maximum norm on the sub-vector $\boldsymbol{Q}_{\mathcal{G}}$. The key is to attain a debiased-type estimator via constructing the inverse matrix of the Hessian matrix of the population loss function on the target dataset, i.e. $\boldsymbol{\Theta}=(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)})^{-1}$.
Note that the transfer learning is also extended to attain a better estimator of $\boldsymbol{\Theta}$ by using nodewise regression on the target dataset and the informative dataset $\widehat{\mathcal{A}}_{\eta}$ (Tian and Feng, 2023; Yang et al., 2025a, b). However, there is no guarantee that the informative sets for regression coefficients and design matrix are the same. Another problem is that the regressors $\hat{\boldsymbol{U}}^{(k)}$ in FARM are estimated such that estimating $\boldsymbol{\Theta}$ via transfer learning does not improve estimation accuracy. Hence, we construct the debiased estimator for $\boldsymbol{\beta}$ as
[…]
where $\hat{\boldsymbol{\beta}}$ is the estimator attained via Algorithm 1, and $\hat{\boldsymbol{\Theta}}$ is an estimator of $\boldsymbol{\Theta}$ by estimating the inverse matrix of $\hat{\boldsymbol{U}}^{(0)\top}\hat{\boldsymbol{U}}^{(0)}/n_{0}$. We do not constrain $\boldsymbol{\Theta}$ to be attained from any specific type, but only need to satisfy the following assumption.
Assumption 5.
There exist positive $\Delta_{1}$, $\Delta_{\max}$, $\Delta_{\infty}$, and $0<\Delta_{\sigma}\leq 1$ such that $|\hat{\sigma}/\sigma_{0}-1|=O_{P}(\Delta_{\sigma})$, and
$\|\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\|_{\max}=O_{P}(\Delta_{1}),\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\max}=O_{P}(\Delta_{\max}),\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\infty}=O_{P}(\Delta_{\infty}),$
where $\sigma_{0}^{2}=\mbox{Var}(\mathcal{E}^{(0)})$, and $\Delta_{1}$, $\Delta_{\max}$, $\Delta_{\infty}$, and $\Delta_{\sigma}$ satisfies
$\sqrt{n_{0}\log p}\Delta_{1}\mathcal{R}_{1}+\Delta_{\infty}\log p\to 0\quad\text{and}\quad\left(\mathcal{V}_{n_{0},p}\|\boldsymbol{\varphi}^{(0)}\|_{2}+\sqrt{\log p}\right)\|\boldsymbol{\Theta}\|_{\infty}\sqrt{\frac{\log p}{n_{0}}}\to 0,$
where $\mathcal{R}_{1}$ is the $\ell_{1}$-error bound for $\hat{\boldsymbol{\beta}}$ established in Theorem 3.1. We further assume that $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\infty}=O_{P}(\|\boldsymbol{\Theta}\|_{\infty})$ without loss of generality.
This assumption on the estimation of $\hat{\boldsymbol{\Theta}}$ can be easily satisfied for the CLIME-type estimator (Cai et al., 2011; Javanmard and Montanari, 2014; Yan et al., 2023; Cai et al., 2025) and the nodewise regression (Zhang and Zhang, 2014; van de Geer et al., 2014; Zhang and Cheng, 2017). For the assumption on the estimation of $\hat{\sigma}$, one can adopt the refitted cross-validation approach in Fan et al. (2012) or the scaled-Lasso (Sun and Zhang, 2012) to attain a consistent estimation of $\sigma_{0}$ by following Fan et al. (2024).
Now we begin by the case that $\mathcal{G}=[p]$ and extend it to any subset $\mathcal{G}\subseteq[p]$. The following theorem establishes a uniform Gaussian approximation for the sup-norm of the debiased estimator $\tilde{\boldsymbol{\beta}}$, which forms the basis for simultaneous inference.
Theorem 4.1.
Suppose that $\hat{\boldsymbol{\beta}}$ satisfies the error bounds shown in Theorem 3.1, and Assumptions 1-2 hold for $k=0$ and Assumption 5 also holds, then there exists a $p$-dimensional Gaussian random vector $\boldsymbol{Q}$ with mean $\boldsymbol{0}_{p}$ and variance $\sigma^{2}\boldsymbol{\Theta}$ such that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\infty}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}\|_{\infty}\leq t)\right|=o(1).$
The proof of Theorem 4.1 is provided in Appendix A.3. From Theorem 4.1, it is natural to attain the $(1-\alpha)$ quantile of $\boldsymbol{Q}$, denoted by $c_{1-\alpha}$, and then construct the simultaneous confidence intervals. However, when $p$ is large, it is still unclear whether the sup-norm of $\boldsymbol{Q}$ can be reliably estimated by sampling from the $p$-dimensional Gaussian distribution with an estimated covariance matrix. Hence, we adopt the multiplier bootstrap procedure. Let $e_{i}\sim\mathbb{N}(0,1),i=1,\cdots,n_{0}$ be independent of the data $\mathcal{D}=\{(\boldsymbol{x}_{i}^{(0)},y_{i}^{(0)},\boldsymbol{f}_{i}^{(0)},\boldsymbol{u}_{i}^{(0)},\mathcal{E}_{i}^{(0)})\}_{i=1}^{n_{0}}$, and consider
$\hat{\boldsymbol{Q}}^{e}=\hat{\sigma}\hat{\boldsymbol{\Theta}}\frac{1}{\sqrt{n_{0}}}\sum_{i=1}^{n_{0}}e_{i}\hat{\boldsymbol{u}}_{i}^{(0)}.$
Denote the $\alpha$-quantile of $\|\hat{\boldsymbol{Q}}^{e}\|_{\infty}$ by
$\hat{c}(\alpha)=\inf_{t\in\mathbb{R}}\{t\in\mathbb{R}:\mathbb{P}_{e}(\|\hat{\boldsymbol{Q}}^{e}\|_{\infty}\leq t)\geq\alpha\},$
where $\mathbb{P}_{e}(\mathcal{H})$ represents the probability of the event $\mathcal{H}$ with respect to $e_{1},\cdots,e_{n}$. The validity of the bootstrap approach is established by the following theorem.
Theorem 4.2.
Under the conditions of Theorem 4.1, if
$\log p(\Delta_{1}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}+\Delta_{\sigma})=o(1),$
we have
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}^{e}\|_{\infty}\leq t|\mathcal{D}\right)-\mathbb{P}\left(\|\boldsymbol{Q}\|_{\infty}\leq t\right)\right|=o(1).$
The proof of Theorem 4.2 can be found in Appendix A.4. Following Theorem 4.1 and Theorem 4.2, our simultaneous confidence interval for $\boldsymbol{\beta}$ is given by
$\left[\tilde{\beta}_{i}-n_{0}^{-1/2}\hat{c}_{\mathcal{G}}(1-\alpha),\tilde{\beta}_{i}+n_{0}^{-1/2}\hat{c}_{\mathcal{G}}(1-\alpha)\right],i\in[p].$
In contrast to the target-only inferential scheme of Fan et al. (2024), our procedure is more reliable because the estimator $\hat{\boldsymbol{\beta}}$ obtained from Algorithm 1 provides a more accurate estimate of $\boldsymbol{\beta}$. In particular, Assumption 5 implies that part of the inferential error captured by $\mathcal{R}_{1}$ can be substantially improved as long as $\mathcal{R}_{1}$ is smaller than the traditional $\ell_{1}$-error $s\sqrt{\log p/n_{0}}$ in Fan et al. (2024).
Now we extend the previous results for any $\mathcal{G}\subseteq[p]$. The following corollary presents the validity of constructing simultaneous confidence intervals for $\boldsymbol{\beta}_{\mathcal{G}}$ via bootstrap procedure.
Corollary 4.1.
Under the conditions of Theorem 4.2, there exists a $p$-dimensional Gaussian random vector $\boldsymbol{Q}$ with mean $\boldsymbol{0}_{p}$ and variance $\sigma^{2}\boldsymbol{\Theta}$ such that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\mathcal{G}}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}\|_{\mathcal{G}}\leq t)\right|=o(1),$
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}^{e}\|_{\mathcal{G}}\leq t|\mathcal{D}\right)-\mathbb{P}\left(\|\boldsymbol{Q}\|_{\mathcal{G}}\leq t\right)\right|=o(1).$
The proof of Corollary 4.1 is presented in Appendix A.5. From Corollary 4.1, let the $\alpha$-quantile of $\|\hat{\boldsymbol{Q}}^{e}\|_{\mathcal{G}}$ by
$\hat{c}_{\mathcal{G}}(\alpha)=\inf_{t\in\mathbb{R}}\{t\in\mathbb{R}:\mathbb{P}_{e}(\|\hat{\boldsymbol{Q}}^{e}\|_{\mathcal{G}}\leq t)\geq\alpha\},$
the simultaneous $\alpha$-level confidence interval for $\boldsymbol{\beta}_{\mathcal{G}}$ can be constructed as
$\left[\tilde{\beta}_{i}-n_{0}^{-1/2}\hat{c}_{\mathcal{G}}(1-\alpha),\tilde{\beta}_{i}+n_{0}^{-1/2}\hat{c}_{\mathcal{G}}(1-\alpha)\right],i\in\mathcal{G}.$
[…]
$\hat{\boldsymbol{Q}}_{stu}^{e}=\hat{\sigma}\hat{\boldsymbol{S}}^{-1/2}\hat{\boldsymbol{\Theta}}\frac{1}{\sqrt{n_{0}}}\sum_{i=1}^{n_{0}}e_{i}\hat{\boldsymbol{u}}_{i}^{(0)},$
and denote the $\alpha$-quantile of $\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\mathcal{G}}$ by
$\hat{c}_{\mathcal{G},stu}(\alpha)=\inf_{t\in\mathbb{R}}\{t\in\mathbb{R}:\mathbb{P}_{e}(\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\mathcal{G}}\leq t)\geq\alpha\}.$
The validity of the studentized bootstrap procedure is established in the following theorem.
Theorem 4.3.
Under the conditions of Theorem 4.2, further assume that $\Delta_{\infty}\|\boldsymbol{\Theta}\|_{\infty}\sqrt{\log p/n_{0}}=o(1)$, then there exists a $p$-dimensional Gaussian random vector $\boldsymbol{Q}_{stu}$ with mean $\boldsymbol{0}_{p}$ and variance $\sigma^{2}\boldsymbol{S}^{-1/2}\boldsymbol{\Theta}\boldsymbol{S}^{-1/2}$ such that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n}\hat{\boldsymbol{S}}^{-1/2}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\mathcal{G}}\leq t\right)-\mathbb{P}(\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\mathcal{G}}\leq t)\right|=o(1),$
and
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\infty}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}_{stu}\|_{\infty}\leq t)\right|=o(1).$
The proof of Theorem 4.3 is given in Appendix A.6. Following Theorem 4.3, the simultaneous $\alpha$-level confidence interval with varying length for $\boldsymbol{\beta}_{\mathcal{G}}$ can be constructed as follows:
$\left[\tilde{\beta}_{i}-n_{0}^{-1/2}(\hat{\boldsymbol{\Theta}}_{i,i})^{1/2}\hat{c}_{\mathcal{G},stu}(1-\alpha),\tilde{\beta}_{i}+n_{0}^{-1/2}(\hat{\boldsymbol{\Theta}}_{i,i})^{1/2}\hat{c}_{\mathcal{G},stu}(1-\alpha)\right],i\in\mathcal{G}.$
The complete steps for obtaining $\tilde{\boldsymbol{\beta}}$ and constructing simultaneous confidence intervals with varying lengths are summarized in Algorithm 3. The corresponding procedure for the non-studentized version follows an analogous structure and is therefore omitted for brevity.
Input: target data $(\boldsymbol{X}^{(0)},\boldsymbol{Y}^{(0)})$, $\mathcal{G}$, estimations $\hat{\boldsymbol{\Theta}}$ and $\hat{\sigma}$.
Output: the simultaneous confidence interval for $\boldsymbol{\beta}_{\mathcal{G}}$
1 Run Algorithm 2 to attain $\hat{\boldsymbol{\beta}}$, estimate $\hat{\boldsymbol{\Theta}}$ using CLIME or nodewise regression, and calculate $\tilde{\boldsymbol{\beta}}$ via (4.1).
2 for $\ell=1$ to $B$ do
3   Draw standard normal random variables $e_{\ell,i},i=1,\cdots,n_{0}$.
4   Calculate $\hat{\boldsymbol{Q}}_{stu,\ell}^{e}=\hat{\sigma}\hat{\boldsymbol{S}}^{-1/2}\hat{\boldsymbol{\Theta}}\sum_{i=1}^{n_{0}}e_{\ell,i}\hat{\boldsymbol{u}}_{i}^{(0)}$
5 end for
6Simultaneous confidence intervals construction: Find the $(1-\alpha)$-quantile of $\{\|\hat{\boldsymbol{Q}}_{stu,\ell}^{e}\|_{\mathcal{G}}\}_{\ell=1}^{B}$, represented as $\hat{c}_{\mathcal{G},stu}(1-\alpha)$. The simultaneous $\alpha$-level confidence interval for $\beta_{i}$ is $\mathcal{I}_{i}=\left[\tilde{\beta}_{i}-n_{0}^{-1/2}(\hat{\boldsymbol{\Theta}}_{i,i})^{1/2}\hat{c}_{\mathcal{G},stu}(1-\alpha),\tilde{\beta}_{i}+n_{0}^{-1/2}(\hat{\boldsymbol{\Theta}}_{i,i})^{1/2}\hat{c}_{\mathcal{G},stu}(1-\alpha)\right],i\in\mathcal{G}.$ Output the simultaneous confidence intervals $\{\mathcal{I}_{i}\}_{i\in\mathcal{G}}$.
Algorithm 3 Simultaneous confidence interval construction
5 Numerical Studies
In this section, we provide numerical experiments to illustrate the finite-sample performance of our proposed Trans-FARM estimator under the FARM framework. The code is made available at https://github.com/BoFuxjtu/Trans-FARM.git.
5.1 Simulations of Estimation
In this section, we study the estimation performance of different methods under various settings of $h$. The methods include Only-FARM (Algorithm 1 with $\mathcal{A}$ being $\varnothing$; Fan et al. (2024)), Trans-FARM (Algorithm 2), Oracle-Trans-FARM (Algorithm 1 with $\mathcal{A}=\mathcal{A}_{\eta}$), Pooled-Trans-FARM (Algorithm 1 with $\mathcal{A}$ being all sources), Only-Lasso (Lasso on target data), Trans-Lasso (Tian and Feng, 2023, Algorithm 2), Oracle-Trans-Lasso (Algorithm 1 in Tian and Feng (2023) with $\mathcal{A}=\mathcal{A}_{\eta}$), Pooled-Trans-Lasso (Algorithm 1 in Tian and Feng (2023) with $\mathcal{A}$ being all sources). The $\epsilon_{0}\hat{\sigma}^{2}$ in Trans-FARM is set as $2\hat{L}_{0}^{(0)}$ (see Algorithm 2), similar to the setting in Tian and Feng (2023). All experiments are conducted in R.
We generate the data from the FARM model as follows. We let $r_{0}=\ldots=r_{K}=2$, and the design matrix is generated by
$\boldsymbol{X}^{(k)}=\boldsymbol{F}_{k}\boldsymbol{B}_{k}^{\top}+\boldsymbol{U}_{k}\in\mathbb{R}^{n_{k}\times p}$
[…]
Figure 1: Comparison of $\ell_{2}$ errors averaged over $200$ replications versus the number of informative sets ($|\mathcal{A}_{\eta}|$) across four cases with $\eta\in\{5,10\}$, $K=10$, $n_{0}=n_{1}=\ldots=n_{K}=300$, $p=500$, and $s=20$.
We set $n_{0}=300$, $n_{1}=\ldots=n_{K}=300$, $p=500$, $s=20$ and $\eta=5$ or $\eta=10$. We let $K=10$ and the informative set $\mathcal{A}_{\eta}$ is random chosen from $\{1,\cdots,K\}$ with $|\mathcal{A}_{\eta}|$ varies from 0 to $10$. We repeat simulation $200$ times and exhibit the $\ell_{2}$ errors of the $8$ estimators under the four scenarios in Figure 1: (a) $\boldsymbol{U}_{k}\sim\mathbb{N}(\boldsymbol{0}_{p},\boldsymbol{\Sigma}_{k})$, $\boldsymbol{E}^{(k)}\sim\mathbb{N}(0,1)$; (b) $\boldsymbol{U}_{k}\sim\mathbb{N}(\boldsymbol{0}_{p},\boldsymbol{\Sigma}_{k})$, $\boldsymbol{E}^{(k)}\sim t_{5}$; (c) $\boldsymbol{U}_{k}\sim t_{10}(\boldsymbol{0},\boldsymbol{\Sigma}_{k})$, $\boldsymbol{E}^{(k)}\sim\mathbb{N}(0,1)$; (d) $\boldsymbol{U}_{k}\sim t_{10}(\boldsymbol{0},\boldsymbol{\Sigma}_{k})$, $\boldsymbol{E}^{(k)}\sim t_{5}$.
From Figure 1, we draw several conclusions. First, across all experimental regimes, FARM-based estimators dominate Lasso-based estimators in Tian and Feng (2023), consistent with the theoretical expectation that Lasso suffers from model misspecification in the presence of latent factors. Second, the Oracle-Trans-FARM procedure exhibits strictly smaller estimation error than the Target-only estimators, with the error decreasing monotonically in the number of informative sources, which aligns with the theoretical benefits of borrowing strength across similar datasets. Moreover, the benefits of Oracle-Trans-FARM compared with Oracle-Trans-Lasso increase as the number of informative sources increases, which states that Oracle-Trans-FARM can capture more information from the informative sources in this setting. Third, the errors of Trans-FARM align with the Oracle-Trans-FARM, indicating that our source detection algorithm accurately recovers the underlying transferable set. Finally, the poor performance of Pooled-Trans-FARM relative to Trans-FARM highlights the necessity of the source detection algorithm. These results collectively show that Trans-FARM achieves near-oracle accuracy while adaptively identifying informative sources.
5.2 Simulations of Simultaneous Inference
In this section, we generate data following the same model as in Section 5.1, with the only differences being that we fix the number of informative sets to $|\mathcal{A}_{\eta}|=5$, the number of non-zero elements in $\boldsymbol{\beta}$ to $s=5$, and $\eta=5$. The data is simulated with the settings $n_{1}=\ldots n_{K}=500$, and we let $n_{0}\in\{200,500\}$ and $p\in\{200,500\}$. The parameters of interest correspond to the first $p/4$ features, and our objective is to construct the simultaneous confidence intervals for these features. We compare our proposed method (Algorithm 3), denoted by Trans-FARM, with the approach extended from Fan et al. (2024) (denoted by Only-FARM). We estimate the $\boldsymbol{\Theta}$ using the nodewise regression (Zhang and Zhang, 2014; van de Geer et al., 2014; Zhang and Cheng, 2017) on target data, denoted by $\hat{\boldsymbol{\Theta}}$, and apply $\hat{\boldsymbol{\Theta}}$ to Trans-FARM and Only-FARM. The results (in $500$ replications) of average coverage probability (ACP) and average length (AL) of $95\%$ confidence intervals under four cases in Section 5.1 are presented in Table 1.
Table 1: Comparison of ACP and AL averaged over $500$ replications for $p,n_{0}\in\{200,500\}$.
Methods Case $p=200,n_{0}=200$ $p=500,n_{0}=200$ $p=200,n_{0}=500$ $p=500,n_{0}=500$
ACP AL ACP AL ACP AL ACP AL
Only-FARM (a) 0.700 0.552 0.764 0.584 0.950 1.791 0.856 0.375
Trans-FARM 0.938 0.546 0.964 0.575 0.956 2.077 0.958 0.373
Only-FARM (b) 0.680 0.702 0.726 0.747 0.928 2.230 0.804 0.481
Trans-FARM 0.948 0.695 0.962 0.737 0.942 2.640 0.962 0.478
Only-FARM (c) 0.696 0.504 0.746 0.536 0.964 1.838 0.786 0.337
Trans-FARM 0.956 0.495 0.936 0.520 0.960 2.098 0.946 0.334
Only-FARM (d) 0.638 0.640 0.706 0.682 0.946 2.335 0.688 0.433
Trans-FARM 0.948 0.627 0.954 0.662 0.956 2.660 0.956 0.430
It can be seen from Table 1 that Only-FARM attains coverage close to $95\%$ ACP only when $p=200,~n_{0}=500$. In contrast, Trans-FARM consistently achieves coverage rates close to $95\%$ under all settings. Meanwhile, the ALs of Only-FARM and Trans-FARM are similar, which suggests that Trans-FARM can leverage auxiliary datasets to improve inferential accuracy and can be more reliable.
5.3 Real Data Analysis
In this Section, we apply our method to the FRED-MD dataset (McCracken and Ng, 2016), which contains $134$ monthly U.S. macroeconomic time series capturing various aspects of economic activity. These variables are highly correlated and are commonly modeled through a low-dimensional latent factor structure. In our analysis, we separately consider ‘HOUSTNE’ (housing starts in the northeast region) and ‘GS5’ ($5$-year treasury rate) as response variables, while treating all remaining variables as covariates. Due to pronounced structural changes associated with the $2008$ financial crisis, many series exhibit nonstationary behavior even after standard transformations. To mitigate this issue, following Fan et al. (2024), we conduct our analysis over two relatively stable subsamples: February 1992-October 2007 and August 2010-February 2020.
For each predictive target subsample, we implement a rolling-window forecasting strategy with a window size of 90 months. Each rolling window constitutes the target dataset for the subsequent observation prediction. The non-overlapping observations in the same subsample are used as the first source dataset, and the remaining subsample is treated as an additional source dataset, enabling information transfer across time periods. Prediction performance is then evaluated by aggregating results across all rolling windows. Specifically, for a given subsample and any $0<t\leq T-90$ (where $T$ is the sample size in the subsample), the target data is $\{(\boldsymbol{x}_{i},y_{i})\}_{i=t}^{t+90}$, the corresponding objective is to predict $y_{t+91}$, denoted by $\hat{y}_{t+91}$. We compare our proposed Trans-FARM method with Only-FARM (Fan et al., 2024), Only-Lasso, and Trans-Lasso (Tian and Feng, 2023, Algorithm 2), where we utilize the two aforementioned source dataset as auxiliary dataset in the transfer learning methods. We measure the prediction accuracy through the out-of-sample $R^{2}$:
$R^{2}=1-\frac{\sum_{i=91}^{T}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=91}^{T}(y_{i}-\bar{y}_{i})^{2}},$
where $\bar{y}_{i}$ is the in-sample average $\sum_{j=t}^{t+90}y_{j}/90$. The out-of-sample $R^{2}$ results regarding two subsamples and two responses ‘HOUSTNE’ and ‘GS5’ are exhibited in Table 2.
Table 2: Comparison of out-of-sample $R^{2}$ for four models in predicting two responses across two time periods. Entries in bold correspond to the method achieving the maximum out-of-sample $R^{2}$ in each setting.
Time period Response Only-Lasso Trans-Lasso Only-FARM Trans-FARM
02.1992-10.2007 HOUSTNE 0.717 0.772 0.759 0.892
GS5 0.687 0.722 0.726 0.740
08.2010-02.2020 HOUSTNE 0.478 0.827 0.759 0.895
GS5 0.754 0.787 0.673 0.805
Table 2 shows that Trans-FARM outperforms the other three methods under all predictive targets. These results demonstrate the effectiveness of Trans-FARM in exploiting auxiliary datasets to improve predictive accuracy. Notably, in contrast to approaches that handle structural breaks via explicit subsample partitioning (Fan et al., 2024), Trans-FARM can still extract and useful information despite the presence of pronounced disruptions associated with the financial crisis. This further illustrates the significance of our method.
6 Conclusion
In this paper, we study the factor-augmented sparse linear model within a transfer learning framework, aiming to address the challenge of strong correlations among covariates and to account for potential latent factor structures, along with leveraging these auxiliary datasets to improve estimation accuracy. This setting is particularly relevant to economic, financial, and medical applications, where the target dataset is often limited in size but accompanied by multiple heterogeneous auxiliary sources that can help enhance the estimation accuracy of regression coefficients. We propose transfer learning procedures that effectively exploit these auxiliary datasets and establish non-asymptotic error bounds for the resulting estimators under both $\ell_{1}$- and $\ell_{2}$- estimation errors. To mitigate the risk of negative transfer, we further develop a transferable source detection method in a data-driven manner and provide a thorough theoretical analysis of its consistency. We also adopt an inferential framework to construct valid simultaneous confidence intervals for regression coefficients of interest. The proposed simultaneous confidence intervals, constructed via a studentized bootstrap procedure, are especially noteworthy in that their lengths vary across coefficients, thereby reflecting the heterogeneous influence of different features on the response. Simulation experiments and a real-data study demonstrate the practical advantages of our methodology, showing that the proposed procedures yield substantial improvements in estimation accuracy, deliver reliable uncertainty quantification, and remain robust across heterogeneous environments. Future work could extend the proposed framework to nonlinear or generalized factor-augmented models, allowing more flexible relationships between the response and covariates. Another direction is to incorporate robust estimation techniques, such as Huber regression or other robust loss functions, into the transfer learning framework to improve stability under heavy-tailed errors and domain shifts induced by latent factor heterogeneity.
Acknowledgments
The authors are supported by NSFC under Grant No. 12571311.
Appendix A Proof of Main Results
This section contains the proofs of the main results. Auxiliary lemmas are deferred to Appendix B.
Denote $\hat{L}(\boldsymbol{w})=\frac{1}{2(n_{\mathcal{A}_{\eta}}+n_{0})}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\|\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}\|_{2}^{2},\,\nabla\hat{L}(\boldsymbol{w})=-\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\\ \hat{\boldsymbol{U}}_{k}^{\top}(\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}),\quad\delta\hat{L}(\boldsymbol{t})=\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}}+\boldsymbol{t})-\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})-{\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})}^{\top}\boldsymbol{t}$, and
$D(\boldsymbol{\Delta})=\boldsymbol{\Delta}^{\top}\left(\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}\right)\boldsymbol{\Delta},\qquad\hat{D}(\boldsymbol{\Delta})=\boldsymbol{\Delta}^{\top}\left(\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}\right)\boldsymbol{\Delta}.$
A.1 Proof of Theorem 3.1
Proof.
Let $\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}-\boldsymbol{w}^{\mathcal{A}_{\eta}}$ and $\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}$ be the bias of the the transferring step and eventual estimator respectively. We firstly consider the error bound for $\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$ from which we are committed to derive the error bound for $\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}$.
Step 1: Bounds for $\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$. By the definition of $\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}$, Lemma B.1, and the convexity of $\hat{L}$, under the event $\mathcal{E}_{\boldsymbol{w}}=\{\lambda_{\boldsymbol{w}}\geq 2\|\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})\|_{\infty}\}$, it follows that
$0$ $\leq\delta\hat{L}(\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\leq\lambda_{\boldsymbol{w}}(\|\boldsymbol{w}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})-\lambda_{\boldsymbol{w}}(\|\hat{\boldsymbol{w}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\|\hat{\boldsymbol{w}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})-{\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})}^{\top}\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$ (A.1)
$\leq\lambda_{\boldsymbol{w}}(\|\boldsymbol{w}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})-\lambda_{\boldsymbol{w}}(\|\hat{\boldsymbol{w}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\|\hat{\boldsymbol{w}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})+\frac{1}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}$
$\leq\frac{3}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\lambda_{\boldsymbol{w}}(\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\|\hat{\boldsymbol{w}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})+\frac{1}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}$
$\leq\frac{3}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{1}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{w}}\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}\leq\frac{3}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{1}{2}\lambda_{\boldsymbol{w}}\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{w}}C_{1}\eta$
It can be implied from (A.1) that $\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}\leq 3\|\hat{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta$ such that $\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\in\mathbb{C}_{\eta}$ where $\mathbb{C}_{\eta}=\{\boldsymbol{t}:\|\boldsymbol{t}_{\mathcal{S}^{c}}\|_{1}\leq 3\|\boldsymbol{t}_{\mathcal{S}}\|_{1}+4C_{1}\eta\}$. Meanwhile, it is straightforward that $t\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\in\mathbb{C}_{\eta}$ for any $t\in(0,1)$. We term to consider the event
$\mathcal{E}_{\boldsymbol{w}}^{\prime}=\left\{\frac{D(\boldsymbol{\Delta})}{\|\boldsymbol{\Delta}\|_{2}^{2}}\geq\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\frac{\|\boldsymbol{\Delta}\|_{1}}{\|\boldsymbol{\Delta}\|_{2}}\right),\qquad\text{for all }\|\boldsymbol{\Delta}\|_{2}\leq 1\right\}.$
We aim to show that conditional on $\mathcal{E}_{\boldsymbol{w}}\cap\mathcal{E}_{\boldsymbol{w}}^{\prime}$,
$\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\leq 8\kappa_{2}C_{1}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\eta+3\kappa_{1}^{-1}\sqrt{s}\lambda_{\boldsymbol{w}}+2\sqrt{\frac{\lambda_{\boldsymbol{w}}C_{1}\eta}{\kappa_{1}}}+\sqrt{\frac{2e_{\eta}}{\kappa_{1}}}:=c_{\mathcal{T}}$
through proof by contradiction, where $e_{\eta}$ is defined as $e_{\eta}=c\eta^{2}((|\mathcal{A}_{\eta}|+1)\log p/(n_{\mathcal{A}_{\eta}}+n_{0})+1/p)$ for some positive constant $c<\infty$. If this claim does not hold, we could find some $t_{0}\in(0,1)$ such that $\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}=t\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$ satisfying $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\geq c_{\mathcal{T}}$ and $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\leq 1$ and as long as $c_{\mathcal{T}}\leq 1$. Let $\tilde{\boldsymbol{w}}^{\mathcal{A}_{\eta}}=\boldsymbol{w}^{\mathcal{A}_{\eta}}+\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$. The optimality of $\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}$ implies that $F(\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\leq 0$ with $F(\boldsymbol{t})=\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}}+\boldsymbol{t})-\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})+\lambda_{\boldsymbol{w}}\|\boldsymbol{w}^{\mathcal{A}_{\eta}}+\boldsymbol{t}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}$. The convexity of function $F(\cdot)$ and $F(\boldsymbol{0}_{p})=0$ yield that $F(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\leq tF(\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\leq 0$. Following a same trick of (A.1), together with the convexity of function $F(\cdot)$ yield that
$F(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})\geq\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})^{\top}\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})+\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}$
$\geq$ $-\frac{1}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$
$\geq$ $-\frac{1}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{3}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+\lambda_{\boldsymbol{w}}(\|\tilde{\boldsymbol{w}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1})+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$
$\geq$ $\frac{1}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{3}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}-2\lambda_{\boldsymbol{w}}\|\boldsymbol{w}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}+D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})-D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$
$\geq$ $\frac{1}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}-\frac{3}{2}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}-2\lambda_{\boldsymbol{w}}C_{1}\eta+\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\frac{\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}}{\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}}\right)\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}^{2}+\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})-D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})$
$\geq$ $\kappa_{1}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}^{2}-\kappa_{1}\kappa_{2}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}-\frac{3}{2}\sqrt{s}\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{2}-2\lambda_{\boldsymbol{w}}C_{1}\eta-|\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})-D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})|$
Note that $\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\in\mathbb{C}_{\eta}$, which deduces that $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}\leq 4\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta\leq 4\sqrt{s}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{2}+4C_{1}\eta,$ and by Lemma B.6, it is straightforward that
$|\hat{D}(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})-D(\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}})|$ $\leq\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}^{2}\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\left\|\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}(\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}-\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k})\right\|_{\max}$
$=O_{P}\left(\left(s\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}^{2}+\eta^{2}\right)\left(\frac{(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}+\frac{1}{p}\right)\right).$
Hence, we have
$\frac{1}{2}\kappa_{1}\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}^{2}-\left(4\kappa_{1}\kappa_{2}C_{1}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\eta+\frac{3}{2}\sqrt{s}\lambda_{\boldsymbol{w}}\right)\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}-2\lambda_{\boldsymbol{w}}C_{1}\eta-e_{\eta}\leq 0$
as long as $16\kappa_{2}\sqrt{s\log p/(n_{\mathcal{A}_{\eta}}+n_{0})}\leq 1$ and $s(|\mathcal{A}_{\eta}|+1)\log p/(n_{\mathcal{A}_{\eta}}+n_{0})+s/p<c$ for a constant $c$ small enough. However, the left-hand side of the inequality would be positive if the claim $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\geq c_{\mathcal{T}}$ holds. This leads to a contradiction, implying that $\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}\leq c_{\mathcal{T}}$, i.e.,
$\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}$ $\lesssim\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\eta+\sqrt{s}\lambda_{\boldsymbol{w}}+\sqrt{\lambda_{\boldsymbol{w}}\eta}+\eta\sqrt{\frac{(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\frac{\eta}{\sqrt{p}}$
$\lesssim\eta\sqrt{\frac{(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\sqrt{s}\lambda_{\boldsymbol{w}}+\sqrt{\lambda_{\boldsymbol{w}}\eta}+\frac{\eta}{\sqrt{p}}$
By choosing $\lambda_{\boldsymbol{w}}\asymp\sqrt{\log p/(n_{\mathcal{A}_{\eta}}+n_{0})}$, then we obatin
$\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{2}$ $\lesssim\eta\sqrt{\frac{(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\sqrt{\frac{s\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\left(\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}},$
$\|\tilde{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}$ $\leq 4\sqrt{s}\|\tilde{\boldsymbol{\mathcal{T}}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{2}+4C_{1}\eta$
$\lesssim\eta\sqrt{\frac{s(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+s\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\left(\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}\right)^{\frac{1}{4}}\sqrt{s\eta}+\frac{\sqrt{s}\eta}{\sqrt{p}}+\eta$
$\lesssim\eta\sqrt{\frac{s(|\mathcal{A}_{\eta}|+1)\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+s\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\eta$
It remains to show that the probability of $\mathcal{E}_{\boldsymbol{w}}\cap\mathcal{E}_{\boldsymbol{w}}^{\prime}$ occurs with high probability. Lemma B.2 ensures that $\mathcal{E}_{\boldsymbol{w}}^{\prime}$ occurs with high probability. Lemma B.4 ensures that $\mathcal{E}_{\boldsymbol{w}}$ occurs with high probability if $\eta=O(1)$ and
$\frac{(|\mathcal{A}_{\eta}|+1)(1\vee\eta\log^{\frac{1}{2}}p\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2}\log^{-\frac{1}{2}}p)}{\sqrt{n_{\mathcal{A}_{\eta}}+n_{0}}}=O(1).$
Step 2: Bounds for $\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}$. Define
$\hat{L}^{(0)}(\boldsymbol{w})$ $=\frac{1}{2n_{0}}\|\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}\boldsymbol{w}\|_{2}^{2},\quad\nabla\hat{L}^{(0)}(\boldsymbol{w})=\frac{1}{n_{0}}\hat{\boldsymbol{U}}_{0}^{\top}(\tilde{\boldsymbol{Y}}^{(0)}-\hat{\boldsymbol{U}}_{0}\boldsymbol{w}),$
$\boldsymbol{\delta}^{\mathcal{A}_{\eta}}$ $=\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}},\quad\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}},\quad\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}=\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta},$
$\delta\hat{L}^{(0)}(\boldsymbol{\Delta})$ $=\hat{L}^{(0)}(\boldsymbol{\beta}+\boldsymbol{\Delta})-\hat{L}^{(0)}(\boldsymbol{\beta})-{\nabla\hat{L}^{(0)}(\boldsymbol{\beta})}^{\top}\boldsymbol{\Delta}.$
Similar to (A.1), under the event $\mathcal{E}_{\boldsymbol{\delta}}=\{\lambda_{\boldsymbol{\delta}}\geq 2\|\nabla\hat{L}^{(0)}(\boldsymbol{\beta})\|_{\infty}\}$, we have
$0\leq\delta\hat{L}^{(0)}(\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})$ $\leq\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}\|_{1}+\frac{1}{2}\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}$ (A.2)
$\leq\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}\|_{1}+\frac{1}{2}\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}$
$\leq\frac{3}{2}\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\Delta}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{\delta}}\|(\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})_{\mathcal{S}^{c}}\|_{1}-\frac{1}{2}\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\Delta}}_{\mathcal{S}^{c}}^{\mathcal{A}_{\eta}}\|_{1}$
Since $\hat{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}=\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}=(\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}})-\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}$ with $\|\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}\leq C_{1}\eta$ by Lemma B.1, we conclude that
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}\leq 4\|\hat{\boldsymbol{\Delta}}_{\mathcal{S}}^{\mathcal{A}_{\eta}}\|_{1}+4(\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+C_{1}\eta)\leq 4\sqrt{s}\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}+4\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta.$
We term to consider the event
$\mathcal{E}_{\boldsymbol{\delta}}^{\prime}=\left\{\frac{D^{(0)}(\boldsymbol{\Delta})}{\|\boldsymbol{\Delta}\|_{2}^{2}}\geq\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\frac{\|\boldsymbol{\Delta}\|_{1}}{\|\boldsymbol{\Delta}\|_{2}}\right),\quad\text{for all }\|\boldsymbol{\Delta}\|_{2}\leq 1\right\},$
where $D^{(0)}(\boldsymbol{\Delta})=\boldsymbol{\Delta}^{\top}\left(\boldsymbol{U}_{0}^{\top}\boldsymbol{U}_{0}/n_{k}\right)\boldsymbol{\Delta}$. Now we aim to show that conditional on $\mathcal{E}_{\boldsymbol{\delta}}\cap\mathcal{E}_{\boldsymbol{\delta}}^{\prime}$,
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\leq 8\kappa_{2}C_{1}\sqrt{\frac{\log p}{n_{0}}}\eta+8\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+\sqrt{\frac{2e_{\eta}^{\prime}}{\kappa_{1}}}:=c_{\boldsymbol{\Delta}}$
through proof by contradiction, where $e_{\eta}^{\prime}$ is defined as $e_{\eta}^{\prime}=c(\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}^{2}+\eta^{2})\left(\log p/n_{0}+1/p\right)+2\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{\delta}}C_{1}\eta$ for some positive constant $c<\infty$. If this claim does not hold, we could find some $t_{0}\in(0,1)$ such that $\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}=t\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}$ satisfying $\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\geq c_{\boldsymbol{\Delta}}$ and $\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\leq 1$ and as long as $c_{\boldsymbol{\Delta}}\leq 1$. Also, $\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}\leq 4\sqrt{s}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}+4\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta.$
Similar to Step 1, define $\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})=\boldsymbol{\Delta}^{\top}\left(\hat{\boldsymbol{U}}_{0}^{\top}\hat{\boldsymbol{U}}_{0}/n_{k}\right)\boldsymbol{\Delta}$, then we have
$|\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})-D^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})|$ $\leq\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}^{2}\frac{1}{n_{0}}\left\|\boldsymbol{U}_{0}^{\top}\boldsymbol{U}_{0}-\hat{\boldsymbol{U}}_{0}^{\top}\hat{\boldsymbol{U}}_{0}\right\|_{\max}$
$=O_{P}\left(\left(s\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}^{2}+\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}^{2}+\eta^{2}\right)\left(\frac{\log p}{n_{0}}+\frac{1}{p}\right)\right).$
Write $\tilde{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}=\boldsymbol{\beta}+\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}$. The optimality of $\hat{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}$ implies that $G(\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})\leq 0$ with $G(\boldsymbol{t})=\hat{L}^{(0)}(\boldsymbol{\beta}+\boldsymbol{t})-\hat{L}^{(0)}(\boldsymbol{\beta})+\lambda_{\boldsymbol{w}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}+\boldsymbol{t}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}$. The convexity of function $G(\cdot)$ and $G(\boldsymbol{0}_{p})=0$ yield that $G(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})\leq tG(\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})\leq 0$. Following a same trick of Step 1, conditional on $\mathcal{E}_{\boldsymbol{\delta}}\cap\mathcal{E}_{\boldsymbol{\delta}}^{\prime}$, one attains
$G(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})\geq\nabla\hat{L}^{(0)}(\boldsymbol{\beta})^{\top}\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}+\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})+\lambda_{\boldsymbol{w}}\|\tilde{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{w}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}$
$\geq$ $-\frac{1}{2}\lambda_{\boldsymbol{\delta}}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}+\lambda_{\boldsymbol{\delta}}\|\tilde{\boldsymbol{\beta}}^{\mathcal{A}_{\eta}}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}-\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}+\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})$
$\geq$ $-2\lambda_{\boldsymbol{\delta}}\|\boldsymbol{\beta}-\hat{\boldsymbol{w}}^{\mathcal{A}_{\eta}}\|_{1}+D^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})-|\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})-D^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})|$
$\geq$ $-2\lambda_{\boldsymbol{\delta}}(\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+C_{1}\eta)+\kappa_{1}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}^{2}-\kappa_{1}\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}-|\hat{D}^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})-D^{(0)}(\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}})|.$
It follows that
$\frac{1}{2}\kappa_{1}\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}^{2}-\left(4\kappa_{1}\kappa_{2}C_{1}\sqrt{\frac{\log p}{n_{0}}}\eta+4\kappa_{1}\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}\right)\|\tilde{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}-e_{\eta}^{\prime}\leq 0$
as long as $s\log p/n_{0}+s/p\lesssim c$ for a constant $c$ small enough, where $e_{\eta}^{\prime}=c(\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}^{2}+\eta^{2})\left(\log p/n_{0}+1/p\right)+2\lambda_{\boldsymbol{\delta}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+2\lambda_{\boldsymbol{\delta}}C_{1}\eta$ for some positive constant $c<\infty$. However, the left-hand side of the inequality would be positive if the claim $\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\geq c_{\boldsymbol{\Delta}}$ holds. This leads to a contradiction, implying that $\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\leq c_{\boldsymbol{\Delta}}$, i.e.,
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\lesssim$ $\sqrt{\frac{\log p}{n_{0}}}\eta+\sqrt{\frac{\log p}{n_{0}}}\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+\sqrt{e_{\eta}^{\prime}}$
$\lesssim$ $\sqrt{\frac{\log p}{n_{0}}}\eta+\frac{s\log p}{\sqrt{n_{0}(n_{\mathcal{A}_{\eta}}+n_{0})}}+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\frac{s}{\sqrt{p}}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}$
as long as $s(|\mathcal{A}_{\eta}|+1)\log p/(n_{\mathcal{A}_{\eta}}+n_{0})=O(1)$. Furthermore, assume that $(s\log p/n_{0})^{\frac{1}{2}}=O(1)$, we have
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}\lesssim\sqrt{\frac{\log p}{n_{0}}}\eta+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}.$
Meanwhile,
$\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{1}$ $\leq 4\sqrt{s}\|\hat{\boldsymbol{\Delta}}^{\mathcal{A}_{\eta}}\|_{2}+4\|\hat{\boldsymbol{\mathcal{T}}}^{\mathcal{A}_{\eta}}\|_{1}+4C_{1}\eta\lesssim s\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}+\eta+\sqrt{s\eta}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}$
It remains to show that the probability of $\mathcal{E}_{\boldsymbol{\delta}}\cap\mathcal{E}_{\boldsymbol{\delta}}^{\prime}$ occurs with high probability. Lemma B.3 ensures that $\mathcal{E}_{\boldsymbol{w}}^{\prime}$ occurs with high probability. Lemma B.5 ensures that $\mathcal{E}_{\boldsymbol{w}}$ occurs with high probability under conditions (3.5) and (3.6). ∎
A.2 Proof of Theorem 3.2
Proof.
We write $\hat{\boldsymbol{U}}_{0}^{[r]}=\boldsymbol{e}^{(0)[r]\top}\hat{\boldsymbol{U}}_{0}$, and $L_{0}^{[r]}(\boldsymbol{w})=\mathbb{E}\left[\frac{1}{n_{0}/3}\|\boldsymbol{e}^{(0)[r]\top}(\boldsymbol{Y}^{(0)}-\boldsymbol{U}_{0}\boldsymbol{w}-\boldsymbol{F}_{0}\boldsymbol{\gamma}^{(k)})\|_{2}^{2}\right]$ for convenience. First, we consider $k\in\mathcal{A}_{\eta}$. For $r\in\{1,2,3\}$, $\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})$ can be decomposed as
$\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})=$ $\{\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})+L_{0}^{[r]}(\boldsymbol{\beta})\}+\{L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-L_{0}^{[r]}(\boldsymbol{\beta})\}$
$+\{\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})-\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-L_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+L_{0}^{[r]}(\boldsymbol{\beta})\}+\{L_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})-L_{0}^{[r]}(\boldsymbol{\beta})\}$
$=$ $I_{1}+I_{2}+I_{3}+I_{4}.$
For $I_{2}$, we have
$|I_{2}|=|L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-L_{0}^{[r]}(\boldsymbol{\beta})|=|\frac{1}{2}(\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta})^{\top}\mathbb{E}\left[\frac{\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}}{n_{0}/3}\right](\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta})|\lesssim\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{2}^{2}$
by the definition of $\boldsymbol{\beta}$. Similarly, we have $|I_{4}|\lesssim\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}^{2}$. By Lemma B.7, we have
$\max\{|I_{1}|,|I_{3}|\}\lesssim$ $\sqrt{\frac{\log p}{n_{0}}}(\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}\vee\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{1}\vee\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}^{2}\vee\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{2}^{2})$
$+\frac{\log p}{n_{0}}(\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}^{2}\vee\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{1}^{2}).$
By a similar proof of Theorem 3.1, we have $\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}\lesssim\sqrt{\frac{s\log p}{n_{0}}},\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}\lesssim s\sqrt{\frac{\log p}{n_{0}}},$ and
$\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{2}$ $\lesssim\sqrt{\frac{\log p}{n_{0}}}\eta+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{0}+n_{k}}},$
$\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{\beta}\|_{1}$ $\lesssim s\sqrt{\frac{\log p}{n_{0}+n_{k}}}+\eta+\sqrt{s\eta}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}$
under Assumptions 14. Hence, as long as $\eta=O(1)$, we conclude that
$|I_{1}|+|I_{2}|+|I_{3}|+|I_{4}|\lesssim$ $\frac{s\log p}{n_{0}}+\eta\sqrt{\frac{\log p}{n_{0}}}+\frac{\eta^{2}}{p}+\left(s\sqrt{\frac{\log p}{n_{0}}}+\eta\right)\sqrt{\frac{\log p}{n_{0}}}+\frac{\log p}{n_{0}}\left(s\sqrt{\frac{\log p}{n_{0}}}+\eta\right)^{2}$
$\lesssim$ $\frac{s\log p}{n_{0}}+\eta\sqrt{\frac{\log p}{n_{0}}}+\frac{\eta^{2}}{p}+s^{2}\left(\frac{\log p}{n_{0}}\right)^{\frac{3}{2}}.$
Therefore, under Assumption 4 (d), we have
$\mathbb{P}$ $\left(\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})\geq\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+\epsilon_{0}\hat{\sigma}^{2}\right)\leq\mathbb{P}\left(\epsilon_{0}\hat{\sigma}^{2}\leq|I_{1}|+|I_{2}|+|I_{3}|+|I_{4}|\right)$
$\leq\mathbb{P}\left(\epsilon_{0}\hat{\sigma}^{2}\lesssim\frac{s\log p}{n_{0}}+\eta\sqrt{\frac{\log p}{n_{0}}}+\frac{\eta^{2}}{p}+s^{2}\left(\frac{\log p}{n_{0}}\right)^{\frac{3}{2}}\right)\to 0.$
Now we term to $k\in\mathcal{A}_{\eta}^{c}$. Let
$\boldsymbol{w}^{(0,k)}=\mathop{\arg\min}_{\boldsymbol{w}\in\mathbb{R}^{p}}\mathbb{E}\Bigg[\frac{1}{2n_{0}/3+n_{k}}\Big($ $\|(\boldsymbol{1}_{n_{0}}-\boldsymbol{e}^{(0)[r]})^{\top}(\boldsymbol{Y}^{(0)}-\boldsymbol{U}_{0}\boldsymbol{w}-\boldsymbol{F}_{0}\boldsymbol{\gamma}^{(k)})\|_{2}^{2}+\|\boldsymbol{Y}^{(k)}-\boldsymbol{U}_{k}\boldsymbol{w}-\boldsymbol{F}_{k}\boldsymbol{\gamma}^{(k)}\|_{2}^{2}\Big)\Bigg].$
For $r\in\{1,2,3\}$, $\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})$ can be decomposed as
$\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})$
$=$ $\{\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-\hat{L}_{0}^{[r]}(\boldsymbol{w}^{(0,k)})-L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})+L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})\}+\{L_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})-L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})\}$
$+\{\hat{L}_{0}^{[r]}(\boldsymbol{w}^{(0,k)})-\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})+L_{0}^{[r]}(\boldsymbol{\beta})\}+\{L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})-L_{0}^{[r]}(\boldsymbol{\beta})\}$
$+\{\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})-\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-L_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+L_{0}^{[r]}(\boldsymbol{\beta})\}+\{L_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})-L_{0}^{[r]}(\boldsymbol{\beta})\}$
$=$ $E_{1}+E_{2}+E_{3}+E_{4}+E_{5}+E_{6}.$
Similar to the proof for $k\in\mathcal{A}_{\eta}$ case, we have
$\max\{|E_{1}|,|E_{3}|,|E_{5}|\}$
$\lesssim$ $\left(\sqrt{\frac{\log p}{n_{0}}}(\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{1}\vee\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{1}\vee 1\right)\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{1}\sqrt{\frac{\log p}{n_{0}}}$
$+\left(\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{1}\vee 1\right)\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{1}\sqrt{\frac{\log p}{n_{0}}}+\left(\sqrt{\frac{\log p}{n_{0}}}\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}\vee 1\right)\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{1}\sqrt{\frac{\log p}{n_{0}}}$
$+\sqrt{\frac{\log p}{n_{0}}}(\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}^{2}\vee\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}^{2}\vee\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}^{2})$
and $|E_{6}|\lesssim\|\hat{\boldsymbol{\beta}}^{[r]}-\boldsymbol{\beta}\|_{2}^{2}$. For $E_{2}$, we have
$|E_{2}|$ $\leq|\mathbb{E}[\nabla L_{0}^{[r]}(\boldsymbol{w}^{(0,k)})]^{\top}(\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)})+\frac{1}{2}v_{0}\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}^{2}$
$\leq v_{0}\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}+\frac{1}{2}v_{0}\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}^{2}.$
For $E_{4}$, we obtain $E_{4}\geq\lambda_{\min}(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)})\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}^{2}:=\underline{\kappa}\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}^{2}$. Assumption 4 ensures that we have
$\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{2}$ $\lesssim\sqrt{\frac{\log p}{n_{0}}}\tilde{\eta}+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\tilde{\eta}}+\frac{\tilde{\eta}}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{0}+n_{k}}},$
$\|\hat{\boldsymbol{w}}^{(k)[r]}-\boldsymbol{w}^{(0,k)}\|_{1}$ $\lesssim s\sqrt{\frac{\log p}{n_{0}+n_{k}}}+\tilde{\eta}+\sqrt{s\tilde{\eta}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}$
with probability at least $1-C\exp(-Cn_{0})$. Moreover, by a similar proof of Lemma B.1, we have $\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{2}\leq\|\boldsymbol{w}^{(0,k)}-\boldsymbol{\beta}\|_{1}\lesssim\tilde{\eta}$. Therefore, we have
$|E_{1}|+|E_{2}|+|E_{3}|+|E_{5}|+|E_{6}|\lesssim\frac{s\log p}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}(\tilde{\eta}\vee\tilde{\eta}^{2})+\tilde{\eta}^{\frac{3}{2}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}+\tilde{\eta}\sqrt{\frac{s\log p}{n_{0}+n_{k}}}.$
Hence, we conclude that for $k\in\mathcal{A}_{\eta}^{c}$,
$\mathbb{P}$ $\left(\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})\leq\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+\epsilon_{0}\hat{\sigma}^{2}\right)\leq\mathbb{P}\left(E_{4}\leq\epsilon_{0}\hat{\sigma}^{2}+|E_{1}|+|E_{2}|+|E_{3}|+|E_{5}|+|E_{6}|\right).$
As long as
$\inf_{k\in\mathcal{A}_{\eta}^{c}}\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{2}^{2}\gtrsim\epsilon_{0}\vee\left(\frac{s\log p}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}(\tilde{\eta}\vee\tilde{\eta}^{2})+\tilde{\eta}^{\frac{3}{2}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}+\tilde{\eta}\sqrt{\frac{s\log p}{n_{0}+n_{k}}}\right),$
the probability approaches $1$. Note that by union bounds, we have
$\mathbb{P}(\widehat{\mathcal{A}}\neq\mathcal{A}_{\eta})\leq$ $\sum_{r=1}^{3}\sum_{k\in\mathcal{A}_{\eta}}\mathbb{P}\left(\hat{L}_{0}(\hat{\boldsymbol{w}}^{(k)[r]})\geq\hat{L}_{0}(\hat{\boldsymbol{\beta}}^{[r]})+\epsilon_{0}\hat{\sigma}^{2}\right)$
$+\sum_{r=1}^{3}\sum_{k\in\mathcal{A}_{\eta}^{c}}\mathbb{P}\left(\hat{L}_{0}^{[r]}(\hat{\boldsymbol{w}}^{(k)[r]})\leq\hat{L}_{0}^{[r]}(\hat{\boldsymbol{\beta}}^{[r]})+\epsilon_{0}\hat{\sigma}^{2}\right).$
Here, the probability above tends to 0 since we can change all the error bounds previously analyzed that contained $\log p$ into $\log(Kp)\lesssim(\log K)\vee(\log p)$ and we assume that $K\lesssim p$ without loss of generality. ∎
A.3 Proof of Theorem 4.1
Proof.
Before the proof, the following example illustrates the specific forms of $\Delta_{1}$, $\Delta_{\max}$, and $\Delta_{\infty}$ on Assumption 5.
Example 1 ((Fan et al., 2024, Proposition E.1)).
We have
$\|\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{\Sigma}}_{\boldsymbol{u}}^{(0)}\|_{\max}$ $=O_{P}\left(\sqrt{\frac{\log p}{n_{0}}}+\frac{1}{\sqrt{p}}+\max_{j\in[p]}\frac{\mathcal{V}_{n_{0},p}}{n_{0}}\|\boldsymbol{B}_{0}^{\top}\boldsymbol{\omega}_{j}^{(0)}\|_{2}\right),$
$\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\max}$ $=O_{P}\left(\max_{j\in[p]}\sqrt{|\mathcal{S}_{j}|\left(\frac{\log p}{n_{0}}+\frac{1}{p}\right)}+\max_{j\in[p]}\frac{\mathcal{V}_{n_{0},p}}{n_{0}}\sqrt{|\mathcal{S}_{j}|}\|\boldsymbol{B}_{0}^{\top}\boldsymbol{\omega}_{j}^{(0)}\|_{2}\right),$
$\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\infty}$ $=O_{P}\left(\max_{j\in[p]}|\mathcal{S}_{j}|\sqrt{\frac{\log p}{n_{0}}+\frac{1}{p}}+\max_{j\in[p]}\frac{\mathcal{V}_{n_{0},p}}{n_{0}}|\mathcal{S}_{j}|\|\boldsymbol{B}_{0}^{\top}\boldsymbol{\omega}_{j}^{(0)}\|_{2}\right).$
First, we have the following decomposition:
[…]
The last term in (A.3) can be bounded as
$\|(\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{\Sigma}}_{\boldsymbol{u}})(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\max}\leq\|\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{\Sigma}}_{\boldsymbol{u}}\|_{\max}\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{1}=\Delta_{1}\mathcal{R}_{1},$
where
$\mathcal{R}_{1}=\sqrt{\frac{\log p}{n_{0}}}\eta+\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\eta}+\frac{\eta}{\sqrt{p}}+\sqrt{\frac{s\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}$
is the $\ell_{1}$ error bound established in Theorem 3.1. Additionally, we have
$\|\frac{1}{n_{0}}\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{U}}_{0}^{\top}\boldsymbol{F}_{0}\boldsymbol{\mathcal{\varphi}}^{(0)}\|_{\infty}=O_{P}(\mathcal{V}_{n_{0},p}\|\boldsymbol{\Theta}\|_{\infty}\|\boldsymbol{\mathcal{\varphi}}^{(0)}\|_{2}/n_{0}),$
by Lemma B.6 and
[…]
$=O_{P}\left(\frac{\sqrt{\log p}}{n_{0}}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}\sqrt{\frac{\log p}{n_{0}}}\right).$
[…]
$\Delta_{n_{0},p}$ $=\sqrt{n_{0}}\Delta_{1}\mathcal{R}_{1}+\frac{\mathcal{V}_{n_{0},p}\|\boldsymbol{\Theta}\|_{\infty}\|\boldsymbol{\mathcal{\varphi}}^{(0)}\|_{2}}{\sqrt{n_{0}}}+\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}\sqrt{\log p}=o\left(1\right).$
Additionally, since $\lambda_{\max}(\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)})\leq\|\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\|_{1}\leq 1/\iota$, we have $\min_{j\in[p]}\boldsymbol{\Theta}_{j,j}\geq\lambda_{\min}(\boldsymbol{\Theta})\geq\iota$. Then, $\{\boldsymbol{\Theta}_{j}\boldsymbol{u}_{i}^{(0)}\mathcal{E}_{i}^{(0)}\}_{i=1}^{n_{0}}$ are i.i.d. zero-mean sub-exponential random variables with $\sigma^{2}\boldsymbol{\Theta}_{j,j}\geq\sigma^{2}\iota$ and
$\max_{j\in[p]}\|\boldsymbol{\Theta}_{j}\boldsymbol{u}_{i}^{(0)}\mathcal{E}_{i}^{(0)}\|_{\psi_{1}}\leq\frac{\|\boldsymbol{u}_{i}^{(0)}\|_{\psi_{1}}\|\mathcal{E}_{i}^{(0)}\|_{\psi_{1}}}{\iota}.$
By Lemma B.8, we have
[…]
By Nazarov’s inequality (Chernozhukov et al., 2023a, Proposition 1), since the variance of each element in $\boldsymbol{Q}$ is bounded away from 0 and $\infty$ by assumptions, we can infer that
$\sup_{t\in\mathbb{R}}\mathbb{P}\left(|\|\boldsymbol{Q}\|_{\infty}-t|<\delta\log^{-1/2}(p)\right)\leq C\delta\log^{-1/2}(p)\{\sqrt{2\log p}+2\}.$
By choosing a $\delta$ small enough, the right hand side of the inequality can be arbitrarily small. From Theorem 4.2, the remainder is $o_{P}(1)$, and thus applying Lemma B.9,
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n_{0}}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\infty}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}\|_{\infty}\leq t)\right|=o(1).$
A.4 Proof of Theorem 4.2
Proof.
[…]
$\|\boldsymbol{\Upsilon}-\hat{\boldsymbol{\Upsilon}}\|_{\max}$ $\leq\hat{\sigma}^{2}\|\hat{\boldsymbol{\Theta}}\|_{\infty}\|\boldsymbol{I}_{p}-\hat{\boldsymbol{\Theta}}\hat{\boldsymbol{\Sigma}}_{\boldsymbol{u}}^{(0)}\|_{\max}+\hat{\sigma}^{2}\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}\|_{\infty}\|\hat{\boldsymbol{\Theta}}\|_{\max}+|\sigma^{2}-\hat{\sigma}^{2}|\|\boldsymbol{\Theta}\|_{\max}$
$=O_{P}(\Delta_{1}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}+\Delta_{\sigma})$
Thus, $\log p\|\sigma^{2}\boldsymbol{\Theta}-\hat{\boldsymbol{\Upsilon}}\|_{\max}=o_{P}(1)$ by assumptions. Since the smallest eigenvalue of $\sigma^{2}\boldsymbol{\Theta}$ is bounded away from 0, by the Gaussian comparision result (Chernozhukov et al., 2023b, Lemma 2.1), it can be derived that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}^{e}\|_{\infty}\leq t|\mathcal{D}\right)-\mathbb{P}\left(\|\boldsymbol{Q}\|_{\infty}\leq t\right)\right|=o(1).$
A.5 Proof of Corollary 4.1
Proof.
The proof follows from the proof of Theorems 4.14.2 by considering the subset $\mathcal{G}$. ∎
A.6 Proof of Theorem 4.3
Proof.
[…]
$\|\hat{\boldsymbol{S}}-\boldsymbol{S}\|_{\infty}\leq\|\boldsymbol{\Theta}-\hat{\boldsymbol{\Theta}}\|_{\infty}=O_{P}(\Delta_{\infty}).$
Since $\iota\leq\lambda_{\min}(\boldsymbol{\Theta})\leq\min_{j\in[p]}\boldsymbol{\Theta}_{j,j}\leq\max_{j\in[p]}\boldsymbol{\Theta}_{j,j}\leq\lambda_{\max}(\boldsymbol{\Theta})\leq\iota^{-1}$, which implies that the diagonal elements of $\boldsymbol{\Theta}$ are bounded away from 0, it can be inferred that $\|\hat{\boldsymbol{S}}^{-1/2}-\boldsymbol{S}^{-1/2}\|_{\infty}=O_{P}(\Delta_{\infty}).$ Then it can be concluded that
[…]
$=$ $O_{P}\left(\Delta_{\infty}\|\boldsymbol{\Theta}\|_{\infty}\sqrt{\frac{\log p}{n_{0}}}+\Delta_{n_{0},p}\right),$
where $\Delta_{n_{0},p}$ is defined in the proof of Theorem 4.1. The error bound above is $o_{P}(1)$ by assumptions. Let $\boldsymbol{Q}^{stu}$ be a $p$-dimensional Gaussian random vector with mean $\mathbf{0}$ and variance $\sigma^{2}\boldsymbol{S}^{-1/2}\boldsymbol{\Theta}\boldsymbol{S}^{-1/2}$. Similar to the proof of Theorem 4.1, it can be verified that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\sqrt{n}\hat{\boldsymbol{S}}^{-1/2}(\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta})\|_{\mathcal{G}}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}^{stu}\|_{\mathcal{G}}\leq t)\right|=o(1).$
Following the proof and notation of Theorem 4.2, we have $\hat{\boldsymbol{Q}}^{e}_{stu}\sim\mathbb{N}(\boldsymbol{0}_{p},\hat{\boldsymbol{S}}^{-1/2}\hat{\boldsymbol{\Upsilon}}\hat{\boldsymbol{S}}^{-1/2})$. Since $\hat{\boldsymbol{S}}^{-1/2}$ and $\boldsymbol{S}^{-1/2}$ are bounded diagonal matrices, it can be deduced that
$\|\hat{\boldsymbol{S}}^{-1/2}$ $\hat{\boldsymbol{\Upsilon}}\hat{\boldsymbol{S}}^{-1/2}-\boldsymbol{S}^{-1/2}\boldsymbol{\Upsilon}\boldsymbol{S}^{-1/2}\|_{\max}\leq\|\hat{\boldsymbol{S}}^{-1/2}-\boldsymbol{S}^{-1/2}\|_{\max}\|\hat{\boldsymbol{\Upsilon}}\|_{\max}\|\hat{\boldsymbol{S}}^{-1/2}\|_{\max}$
$+\|\boldsymbol{S}^{-1/2}\|_{\max}\|\hat{\boldsymbol{\Upsilon}}-\hat{\boldsymbol{\Upsilon}}\|_{\max}\|\hat{\boldsymbol{S}}^{-1/2}\|_{\max}+\|\boldsymbol{S}^{-1/2}\|_{\max}\|\boldsymbol{\Upsilon}\|_{\max}\|\hat{\boldsymbol{S}}^{-1/2}-\boldsymbol{S}^{-1/2}\|_{\max}$
$=$ $O_{P}(\Delta_{1}\|\boldsymbol{\Theta}\|_{\infty}+\Delta_{\infty}+\Delta_{\sigma})=o_{P}\left((\log p)^{-1}\right)$
by assumptions. Similar to the proof of Theorem 4.2, we obtain that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\|\hat{\boldsymbol{Q}}_{stu}^{e}\|_{\mathcal{G}}\leq t\right)-\mathbb{P}(\|\boldsymbol{Q}^{stu}\|_{\mathcal{G}}\leq t)\right|=o(1).$
Appendix B Proof of Auxiliary Lemmas
Lemma B.1.
Under Assumption 3, we have $\|\boldsymbol{\boldsymbol{\delta}}^{\mathcal{A}_{\eta}}\|_{1}=\|\boldsymbol{w}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}\|_{1}\leq C_{1}\eta$.
Proof.
By definition, we have $\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})=\boldsymbol{0}_{p},$ which indicates that
$\sum_{k\in\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}}(\boldsymbol{w}^{(k)}-\boldsymbol{\beta})=\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\alpha_{k}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}}(\boldsymbol{w}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}).$
It follows that $\|\boldsymbol{w}^{\mathcal{A}_{\eta}}-\boldsymbol{\beta}\|_{1}\leq\sum_{k\in\mathcal{A}_{\eta}}\alpha_{k}\|(\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}})^{-1}\boldsymbol{\Sigma}^{(k)}_{\boldsymbol{u}}\|_{1}\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{1}\leq C_{1}\eta$. ∎
Lemma B.2 (Lemma 2 of Tian and Feng (2023)).
Suppose that Assumption 1 holds for $k\in\{0\}\cup\mathcal{A}_{\eta}$, then there exists some positive constant $\kappa_{1},\kappa_{2},C_{3}$ and $C_{4}$ such that
$\frac{D(\boldsymbol{\Delta})}{\|\boldsymbol{\Delta}\|_{2}^{2}}\geq\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\frac{\|\boldsymbol{\Delta}\|_{1}}{\|\boldsymbol{\Delta}\|_{2}}\right),\quad\text{for all }\|\boldsymbol{\Delta}\|_{2}\leq 1$
with probability at least $1-C_{3}\exp(-C_{4}(n_{\mathcal{A}_{\eta}}+n_{0}))$.
Proof.
The proof simply follows from Lemma 2 of Tian and Feng (2023). ∎
Lemma B.3 (Proposition 2 of Negahban et al. (2009)).
Suppose that Assumption 1 holds for $k\in\{0\}\cup\mathcal{A}_{\eta}$, then there exists some positive constant $\kappa_{1},\kappa_{2},C_{3}$ and $C_{4}$ such that
$\frac{D^{(0)}(\boldsymbol{\Delta})}{\|\boldsymbol{\Delta}\|_{2}^{2}}\geq\kappa_{1}\left(1-\kappa_{2}\sqrt{\frac{\log p}{n_{0}}}\frac{\|\boldsymbol{\Delta}\|_{1}}{\|\boldsymbol{\Delta}\|_{2}}\right),\quad\text{for all }\|\boldsymbol{\Delta}\|_{2}\leq 1$
with probability at least $1-C_{3}\exp(-C_{4}n_{0})$, where $D^{(0)}(\boldsymbol{\Delta})=\boldsymbol{\Delta}^{\top}\left(\boldsymbol{U}_{0}^{\top}\boldsymbol{U}_{0}/n_{k}\right)\boldsymbol{\Delta}$.
Proof.
The proof simply follows from Proposition 2 of Negahban et al. (2009). ∎
Lemma B.4.
Suppose that Assumptions 1-2 hold for $k\in\{0\}\cup\mathcal{A}_{\eta}$, if $\eta=O(1)$, we have
$\|\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})\|_{\infty}=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)(\sqrt{\log p}\vee\eta\log p\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2})}{n_{\mathcal{A}_{\eta}}+n_{0}}+\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right).$
Proof.
Note that
$-\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})$ $=\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\hat{\boldsymbol{U}}_{k}^{\top}(\tilde{\boldsymbol{Y}}^{(k)}-\hat{\boldsymbol{U}}_{k}\boldsymbol{w}^{(k)})+\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})$
$:$ $=E_{1}+E_{2}.$
[…]
$=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)(\sqrt{\log p}\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2})}{n_{\mathcal{A}_{\eta}}+n_{0}}+\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right).$
For $E_{2}$, according to Lemma B.1 and the definition of $\mathcal{A}_{\eta}$, we have $\|\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}\leq\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta}\|_{1}+\|\boldsymbol{\beta}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}\leq(C_{1}+1)\eta$. This equation, together with Lemma B.6 implies that
$\left\|\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}(\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}-\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k})(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})\right\|_{\infty}$
$\leq$ $\frac{|\mathcal{A}_{\eta}|+1}{(n_{\mathcal{A}_{\eta}}+n_{0})}\|\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}-\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}\|_{\max}\|\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)\eta\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}\right)$
Moreover, by the definition of $\boldsymbol{w}^{\mathcal{A}_{\eta}}$, we have $\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})/(n_{\mathcal{A}_{\eta}}+n_{0})$ is a mean of the sample $\boldsymbol{u}_{i}^{(k)}\boldsymbol{u}_{i}^{(k)\top}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})$ with expectation $\boldsymbol{0}_{p}$ and
$\|\boldsymbol{u}_{i}^{(k)}\boldsymbol{u}_{i}^{(k)\top}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})\|_{\psi_{1}}\leq\|\boldsymbol{u}_{i}^{(k)}\|_{\psi_{2}}^{2}\|\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}}\|_{1}\lesssim\eta.$
It follows from Bernstein’s inequality that
$\frac{1}{n_{\mathcal{A}_{\eta}}+n_{0}}\sum_{k\in\{0\}\cup\mathcal{A}_{\eta}}\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}(\boldsymbol{w}^{(k)}-\boldsymbol{w}^{\mathcal{A}_{\eta}})=O_{P}\left(\eta\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right).$
Hence, we have
$\|E_{2}\|_{\infty}=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)\eta\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}+\eta\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right),$
which indicates that, if $\eta=O(1)$, we have
$\|\nabla\hat{L}(\boldsymbol{w}^{\mathcal{A}_{\eta}})\|_{\infty}=O_{P}\left(\frac{(|\mathcal{A}_{\eta}|+1)(\sqrt{\log p}\vee\eta\log p\vee\max_{k\in\{0\}\cup\mathcal{A}_{\eta}}\mathcal{V}_{n_{k},p}\|\varphi^{(k)}\|_{2})}{n_{\mathcal{A}_{\eta}}+n_{0}}+\sqrt{\frac{\log p}{n_{\mathcal{A}_{\eta}}+n_{0}}}\right).$
Lemma B.5.
Suppose that Assumptions 1-2 hold for $k=0$, we have
$\|\nabla\hat{L}^{(0)}(\boldsymbol{\beta})\|_{\infty}=O_{P}\left(\frac{\sqrt{\log p}\vee\mathcal{V}_{n_{0},p}\|\varphi^{(0)}\|_{2}}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}\right).$
Proof.
Similar to the proof Lemma B.4, we have
[…]
$=$ $O_{P}\left(\frac{\sqrt{\log p}\vee\mathcal{V}_{n_{0},p}\|\varphi^{(0)}\|_{2}}{n_{0}}+\sqrt{\frac{\log p}{n_{0}}}\right).$
Lemma B.6 (Lemma C.3, Lemma C.4, and Lemma C.6 of Fan et al. (2024)).
Suppose that Assumptions 1-2 hold for $k$, and assume that $n_{k}=O(p)$, define
$\mathcal{V}_{n_{k},p}=\frac{n_{k}}{p}+\sqrt{\frac{\log p}{n_{k}}}+\sqrt{\frac{n_{k}\log p}{p}},$
then for any $\boldsymbol{\phi}\in\mathbb{R}^{r_{k}}$ with $\|\boldsymbol{\phi}\|_{2}=1$, we have
[…]
and
$\|\hat{\boldsymbol{U}}_{k}^{\top}\hat{\boldsymbol{U}}_{k}-\boldsymbol{U}_{k}^{\top}\boldsymbol{U}_{k}\|_{\max}=O_{P}\left(\frac{n_{k}}{p}+\log p\right).$
Lemma B.7.
Suppose that Assumptions 1-2 hold for $k\in\{1,\ldots,K\}$, for any $1\leq r\leq 3$ and any $\boldsymbol{w}_{1},\boldsymbol{w}_{2}\in\mathbb{R}^{p}$, we have
$|\hat{L}_{0}^{[r]}(\boldsymbol{w}_{1})-\hat{L}_{0}^{[r]}(\boldsymbol{w}_{2})-(L_{0}^{[r]}(\boldsymbol{w}_{1})-L_{0}^{[r]}(\boldsymbol{w}_{2}))|$
$\lesssim$ $\left(\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}\vee\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{1}\vee 1\right)\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}\sqrt{\frac{\log p}{n_{0}}}$
$+\sqrt{\frac{\log p}{n_{0}}}(\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}^{2}+\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{2}),$
where
$L_{0}^{[r]}(\boldsymbol{w})=\mathbb{E}\left[\frac{1}{n_{0}/3}\|\boldsymbol{e}^{(0)[r]\top}(\boldsymbol{Y}^{(0)}-\boldsymbol{U}_{0}\boldsymbol{w}-\boldsymbol{F}_{0}\boldsymbol{\gamma}^{(k)})\|_{2}^{2}\right]$
Proof.
Firstly, denote $\hat{\boldsymbol{U}}_{0}^{[r]}=\boldsymbol{e}^{(0)[r]\top}\hat{\boldsymbol{U}}_{0}$ and $\boldsymbol{U}_{0}^{[r]}=\boldsymbol{e}^{(0)[r]\top}\boldsymbol{U}_{0}$, we have
$\hat{L}_{0}^{[r]}(\boldsymbol{w}_{1})-\hat{L}_{0}^{[r]}(\boldsymbol{w}_{2})=\nabla\hat{L}_{0}^{[r]}(\boldsymbol{w}_{2})^{\top}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})+\frac{1}{2}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})^{\top}\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})$
$=$ $\nabla\hat{L}_{0}^{[r]}(\boldsymbol{\beta})^{\top}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})+(\boldsymbol{w}_{2}-\boldsymbol{\beta})^{\top}\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})+\frac{1}{2}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})^{\top}\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})$
Similarly, apply this equation to $L_{0}^{[r]}(\boldsymbol{w}_{1})-L_{0}^{[r]}(\boldsymbol{w}_{2})$, together with the equation above yields that
$\hat{L}_{0}^{[r]}(\boldsymbol{w}_{1})-\hat{L}_{0}^{[r]}(\boldsymbol{w}_{2})-(L_{0}^{[r]}(\boldsymbol{w}_{1})-L_{0}^{[r]}(\boldsymbol{w}_{2}))$
$=$ $(\nabla\hat{L}_{0}^{[r]}(\boldsymbol{\beta})-\nabla L_{0}^{[r]}(\boldsymbol{\beta}))^{\top}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})+(\boldsymbol{w}_{2}-\boldsymbol{\beta})^{\top}\left(\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\right)(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})$
$+\frac{1}{2}(\boldsymbol{w}_{1}-\boldsymbol{w}_{2})^{\top}\left(\frac{1}{n_{0}/3}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\right)(\boldsymbol{w}_{1}-\boldsymbol{w}_{2}):=E_{1}+E_{2}+E_{3}.$
By Lemma B.5 and the assumptions, we have $|E_{1}|\leq\|\nabla\hat{L}_{0}^{[r]}(\boldsymbol{\beta})\|_{\infty}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}\lesssim\sqrt{\log p/n_{0}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}$. Now we consider $(n_{0}/3)^{-1}\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}$. By following a similar proof of the line of Lemma C.3 and Lemma C.6 in Fan et al. (2024), we have
$\frac{1}{n_{0}/3}\|\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{U}_{0}^{[r]\top}\boldsymbol{U}_{0}^{[r]}\|_{\max}=O_{P}\left(\frac{\log p}{n_{0}}\right).$
Furthermore, since $\|\boldsymbol{a}^{\top}\boldsymbol{u}^{(0)}\boldsymbol{u}^{(0)\top}\boldsymbol{b}\|_{\psi_{1}}\leq\|\boldsymbol{u}^{(0)}\|_{\psi_{2}}^{2}\|a\|_{2}\|b\|_{2}$ for any vectors $\boldsymbol{a}$ and $\boldsymbol{b}$, by the Bernstein’s inequality, one attains that
$\left\|\boldsymbol{a}^{\top}\left(\frac{1}{n_{0}/3}\boldsymbol{U}_{0}^{[r]\top}\boldsymbol{U}_{0}^{[r]}-\boldsymbol{\Sigma}_{\boldsymbol{u}}^{(0)}\right)\boldsymbol{b}\right\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{a}\|_{2}\|\boldsymbol{b}\|_{2}\right).$
Hence, we conclude that
$|E_{2}|+|E_{3}|\lesssim$ $\left\|\frac{1}{n_{0}/3}(\hat{\boldsymbol{U}}_{0}^{[r]\top}\hat{\boldsymbol{U}}_{0}^{[r]}-\boldsymbol{U}_{0}^{[r]\top}\boldsymbol{U}_{0}^{[r]})\right\|_{\max}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}(\frac{1}{2}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}+\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{1})$
$+\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}(\frac{1}{2}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}+\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{2})$
$\lesssim$ $\frac{\log p}{n_{0}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}(\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{1}\vee\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{1})$
$+\sqrt{\frac{\log p}{n_{0}}}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}(\frac{1}{2}\|\boldsymbol{w}_{1}-\boldsymbol{w}_{2}\|_{2}+\|\boldsymbol{w}_{2}-\boldsymbol{\beta}\|_{2})$
The proof is complete by leveraging the bounds for $|E_{1}|$ and $|E_{2}|+|E_{3}|$. ∎
Lemma B.8 ((Chernozhukov et al., 2013, Corollary 2.1)).
Let $\boldsymbol{X}_{1},\cdots,\boldsymbol{X}_{n}=(X_{i1},\\ \cdots,X_{id})^{\top}$ be n i.i.d. $d$-dimensional random vectors with mean $\mathbf{0}_{d}$. Suppose that there are some constants $0<C_{1}<C_{2}$ such that $\mathbb{E}[X_{ij}|^{2}\geq C_{1}$ and $X_{ij}$ is sub-exponential with $\|X_{ij}\|_{\psi_{1}}\leq C_{2}$ for all $1\leq j\leq d$. If $\log^{7}(dn)/n=o(1)$, then
$\lim_{n\to\infty}\sup_{t\in\mathbb{R}}\left|\mathbb{P}\left(\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\boldsymbol{X}_{i}\right\|_{\infty}\leq t\right)-\mathbb{P}(\|\boldsymbol{N}\|_{\infty}\leq t)\right|=0,$
where $\boldsymbol{N}\sim\mathbb{N}(\mathbf{0}_{d},\mathbb{E}[\boldsymbol{X}_{i}\boldsymbol{X}_{i}^{\top}])$.
Lemma B.9 ((Cai et al., 2025, Lemma 7)).
For any $\delta>0$ and sequences of random vectors $\{X_{n}\}_{n\in\mathbb{N}},\{Y_{n}\}_{n\in\mathbb{N}},\{Z_{n}\}_{n\in\mathbb{N}}$ satisfying
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}(\|X_{n}\|_{\infty}\leq t)-\mathbb{P}(\|Z_{n}\|_{\infty}\leq t)\right|=\Delta_{n},$
we have that
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}(\|X_{n}+Y_{n}\|_{\infty}\leq t)-\mathbb{P}(\|Z_{n}\|_{\infty}\leq t)\right|\leq\mathbb{P}(\|Y_{n}\|_{\infty}>\delta)+\sup_{t\in\mathbb{R}}\mathbb{P}(|\|Z_{n}\|_{\infty}-t|<\delta)+\Delta_{n}.$